### v1

Abstract
This paper investigates the nuanced interplay of gender, birth order, and innate ability in shaping educational disparities among children within households, employing both a reduced-form analysis and a structural model of household resource allocation. By decomposing overall disparities, I identify the contributions of gender, birth order, and innate ability differences. In the context of Benin, for households with non-educated heads and mixed-gender children, total inequality comprises 50% gender disparity, 20% birth order effect, and 30% ability gap. Conversely, in households with only sons or only daughters and non-educated heads, total inequality is predominantly driven by ability disparities—70% for daughters and 83% for sons, with lesser contributions from birth order effects. Furthermore, my analysis reveals that in households with non-educated heads, firstborn daughters surpass their younger brothers in educational attainment on the intensive margin if their innate ability exceeds their brothers’ by at least 13%, a figure reduced to 8% in households with college-educated parents. Additionally, the study unveils parental preferences favoring sons’ education over daughters’ among non-educated parents, with a perceived 22% higher average benefit. Targeted policies aimed at reducing composite education costs prove effective in mitigating gender and birth order gaps, albeit with a modest 5% reduction in total inequality. Overall, this research underscores the complex dynamics influencing intra-household educational inequalities and suggests policy avenues to address them.
JEL classification: I220, I240, H520
Keywords: Education policy, Inequality, Equality of education opportunities, gender gap.
1 Introduction
The trade-off between the number of children one chooses to have and the quality of those children has been a subject of extensive research. This exploration began with the seminal contributions of De Tray (1970) and Becker and Lewis (1973). The quantity-quality trade-off model delves into how changes in the quantity of children can impact their average quality. Numerous empirical studies have provided support for this model in various contexts, consistently revealing a negative relationship between the number of children, often used as an empirical measure of quantity, and the educational attainment of those children, typically used as an empirical measure of quality. These findings align with the predictions put forth by the theoretical framework presented by Becker and Lewis (1973) and have been substantiated by research such as Montgomery (1995) and Li et al. (2008). The current state of the literature on fertility choices (Conley and Glauber (2006), Maralani (2008), Li et al. (2008), Weng et al. (2019)) states that rich and educated families tend to have fewer children while allocating greater investments in the education of those children, in comparison to less affluent and less educated families. In other words, the average level of education among children varies across different families, influenced by endogenous factors such as family size and exogenous factors like the socio-economic background of the parents.
When we extend the analysis beyond the choice of the aggregate quality of children to how it is distributed among children, we observe that, not only does the average change across households, but so does the variance of quality. In addition, the variance is most likely non-zero for a majority of households when there are budget constraints and no compulsory education laws. This presumption of a non-zero variance is justified by evidence of the influence of factors such as gender, birth order and innate ability on education attainment of children. The educational outcomes of children are strongly influenced by their individual characteristics, leading to disparities in the amount of education they receive. Several studies have shed light on these disparities, with findings indicating that girls tend to receive less education due to factors such as gender bias or gender preference (as documented by Biswas (2000) and Ota and Moffatt (2007)). Additionally, birth order can play a role in the educational opportunities afforded to children, with elder siblings benefiting or facing disadvantages, as observed in studies like Ota and Moffatt (2007), Weng et al. (2019), Fergusson et al. (2006), De Haan (2010), Moshoeshoe et al. (2016), and Esposito et al. (2020). Furthermore, children with higher abilities are more likely to receive increased educational opportunities and education attainment, as suggested by the research of Becker and Tomes (1976), Dizon-Ross (2019), and Giannola (2023). However, there exists a notable gap in the research landscape, as there is a scarcity of studies that comprehensively analyze all these various sources of disparities within the same analytical framework. Such a framework, capable of simultaneously examining gender-based differences, birth order effects, and the impact of individual abilities on educational inequality within households, holds the potential for interesting counterfactual and education policies analyses. Specifically, it enables us to explore how these factors interact with one another to shape inequality. Furthermore, this comprehensive approach is crucial for estimating the proportion of inequality attributable to disadvantages and the portion caused by ability-based resource allocation within households. The former allows us to design education policies which target inequality due gender or birth order disadvantages.
In this paper, I use reduced form approach and suggest a structural model to break down the overall intra-household differences in children’s education into parts related to gender, birth order, and ability. Firstly, the model helps us find out if the total disparities observed is solely due to biases from parents and society or because there is high variation in children’s abilities. Secondly, it allows us to measure how much more intelligent a daughter (or firstborn) needs to be in order to have the same educational opportunities as a son (or second born). Lastly, using this model, we can examine how differences in education of children within a family change with a cost reduction policy compared to a policy which aim to change parents’ preference.
I estimate the model using data from Benin in West Africa. Benin is a good place to apply this model for three main reasons. First, even though there is a law stipulating that everybody must finish at least primary school, it’s not always followed by parents or enforced by the government. Second, there are big differences in how much education individuals in the same household get. Also, these differences can be very different from one family to another. Lastly, there is evidence that girls and firstborns often don’t get the same education opportunities as others. For this exercise, I focused on households where there were only two adult children still residing with their parents. Among those households there is difference in how much education resources they have to distribute and the education attainment of the head of household. Taking that into account, I applied my method to households with different characteristics separately.
My analysis uncovered three important findings. First, when it comes to educating daughters compared to their brothers within a household with a non-educated head of household, large part (70%) of the differences in the education of brothers and sisters is due to disadvantages rather than variations in abilities. That number is 33% among college educated parents. However, when examining the education of firstborns compared to their siblings of the same gender, more than half (70% for female and 80% for male) of the disparities are attributed to differences in abilities. These results suggests that in some households gender disadvantage reduces disparities in children’s education through the reduction of the extent to which the talent of a child affect their education. This implies that removing/reducing gender and birth order disadvantages will not systematically reduce inequality for all households. It will increase the amount of disparities in some households.
Second, When we reduce education costs for firstborn children and daughters, we can decrease the differences in education within certain families. This is especially true in families where the firstborn children and daughters have lower abilities compared to their siblings or where parents don’t really take the children’s abilities into consideration when making decisions about education. However, if the firstborn children and daughters are actually more capable than their siblings and parents do consider their abilities, a policy that eases the financial burden of schooling for them might unintentionally increase educational differences within the family. This brings up a dilemma about finding the right balance between giving everyone equal educational opportunities, regardless of factors like gender, birth order, and ability, and ensuring that children with similar abilities get the same opportunities, regardless of gender and birth order. In the latter case we are allowing for rewarding high abilities but not for penalizing gender and birth order.
In terms of policy changes, it seems more implementable to aim for the latter option. That means focusing on policies that increase the net return from education for firstborn children and daughters either through schooling cost reduction or reduction of salary gap between male and female. These policies will help decrease the impact of gender and birth order on educational differences. On the other hand, it’s trickier to implement a policy that reduces how much parents consider a child’s abilities in distributing the few resources that they have, when facing financial challenges. Especially given that it is more costly to give high level formal education to a low ability child, and that there are other more suitable alternatives such as apprenticeship.
Third, When all kids in a family are considered equally capable, the differences in education between them created by gender and birth order is at its highest point. This happens because parents see the benefit from their investment in education and the child’s ability as complementary. In other words, parents think that spending more on the education of a child with higher ability will bring greater benefits compared to investing in a child with lower ability. So, if the firstborn or a daughter has lower abilities than their siblings, they will receive less investment in education even without considering gender and birth order disadvantages. This means that the disadvantages they face because of gender and/or birth order have a smaller impact on increasing the overall disparities in education in term of magnitude. On the flip side, if the firstborn or a daughter is more capable than their siblings, they would get more investment in their education compared to their siblings, in the absence of gender and birth order disadvantages. As a result, the disadvantage they face due to gender and birth order leads to a decrease in the educational differences within the family. This decrease becomes less pronounced as their intelligence level increases.
The subsequent sections of this paper are structured as follows. In Section 2, I present the literature review along with the specific contributions of this paper. Section 3 offers an overview of the data utilized for the estimation of the model, while Section 4 presents key empirical observations and facts derived from the analysis. Section 5 is dedicated for the model’s setup, outlining the identification strategy for key parameters, and describing the inference and estimation procedures employed in this study. Lastly, in Section 6, I present counterfactual analysis to further explore the implications of my findings. Section 7 presents a robustness analysis of the key parameters of interest, and Section 8 concludes the paper.
2 Literature Review
This paper contributes to the extensive literature on fertility choice models and the quantity-quality trade-off faced by households within a country. Early developed theoretical models (DeTray (1970), Becker and Lewis (1973)) predicts that quantity and quality of children are negatively related due to the observation that the shadow price of quantity is influenced by quality, and vice versa. More specially, they have shown that a reduction in the quantity of children results in a decrease in the shadow price of quality. Consequently, this decline in the shadow price of quality leads to an increase in the overall quality of the children. The current state of the literature on fertility choices (Montgomery (1995), Conley and Glauber (2006), Maralani (2008), Li et al. (2008), Weng et al. (2019)) states that rich and educated families tend to have fewer children while allocating greater investments in the education of those children, in comparison to less affluent and less educated families. This trend reflects a significant pattern where socio-economic factors play a crucial role in shaping fertility decisions and educational investments within households. However, some studies still found a positive or no correlation between fertility and education of children either for rural or old households (Montgomery (1995), Black et al. (2005), Maralani (2008)). This is often explained by the lack of contraception methods and poor public education systems. This paper contributes to that strand of the literature by offering two key contributions. Firstly, it empirically demonstrates the existence of a quantity-quality trade-off within the context of Benin. Second, it builds on this result to relax the assumption of equal education for children in the same household.
This paper also contributes to the literature on the within-household schooling decision, particularly factors influencing parents’ distribution decision of education resources among siblings. One key determinant of these distribution decisions is the gender or gender composition of households. Previous research has shown that daughters are less likely to receive education or have lower educational attainment on average. Studies have shown that, while the presence of elder sisters tends to increase the likelihood of schooling, the presence of younger brothers may decrease it (Biswas (2000), Ota and Moffatt (2007), Ombati and Ombati (2012), Osadan and Burrage (2014), Psaki et al. (2018)). Another influential factor is the birth order of children, with mixed findings in previous studies. Some papers suggest a positive effect of birth order on children’s education (Ota and Moffatt (2007), Weng et al. (2019)), while others have shown that later-born children have lower educational attainment (Fergusson et al. (2006), De Haan (2010), Moshoeshoe et al. (2016), Esposito et al. (2020)). Finally, a child’s innate ability or talent plays a role in parental distribution decisions. Studies have demonstrated that parents invest more in the human capital of endowed children and allocate more nonhuman capital to less endowed children (Becker and Tomes (1976), Dizon-Ross (2019), Giannola (2023)). When parents are compelled to invest in the nonhuman capital of low-ability children, this leads to an inefficient equilibrium, where the investment in the human capital of high-ability children is not optimized (Nerlove et al. (1984)). This paper adds to this existing literature in two significant ways. Firstly, it examines a context where parents are not constrained to compensate lower ability children by investing in their nonhuman capital but, instead, rely on family taxes (Wantchekon et al. (2015)). Secondly, this paper develops a household resources distribution structural model that allows for a more flexible analysis of the distribution of education resources within the household. In this model, the assumption of equal distribution is relaxed, enabling a detailed exploration of the interactions between gender and, birth order disadvantages, and the innate abilities of individual children in influencing parental distribution decisions.
Finally, this paper contributes to the literature on within household inequality in children’s human capital (Giannola (2023)). Giannola (2023) has shown in the context of India that observed inequality within households is partly explained by parents investing more in the human capital of high-achieving children, especially when they are financially constrained. This behavior stems from the fact that parents are not particularly averse to inequality and tend to reinforce the gap in learning created by innate ability rather than correct it. This paper contributes to that literature by first building upon the result that parents unequally invest in the human capital of high-achieving children in contexts where the education system is better tailored to serve high-achieving students. Second, this paper interacts with that result and examines how it relates to other sources of unequal distribution, such as gender and birth order.
3 Data Description and Definition of Key Variables
In this section, I present the data used to analyze the question of interest in this paper. The context of Benin is particularly compelling for analyzing the interaction between gender disadvantage, birth order advantage, and innate ability in shaping within household inequality for three reasons. First, there are a lot of disparities in education attainment within households due to the lack of compulsory education law coupled with financial constraints and the availability of outside options other than formal education such as child labor within the households and apprenticeship. Second, there is factual evidence of gender and birth order disadvantages. Lastly, there is observational evidence that despite the prevalence of gender disadvantages in the population, daughters in some households get higher education than their brothers.
3.1 Sample and Data
The 2013 Population and Habitation Census data of Benin; used in this paper; offers a comprehensive insights into households and their members living in the country during that year. Conducted by the National Bureau of Statistics of Benin, this census provides data at both the household and individual levels. For the purpose of this paper, the focus is directed towards individuals who identify themselves as the children of household heads, enabling to get information on parental  [^3] Parents here refers to one of the parents, either the mother or the father. This because it is not possible to have both for household with single parents and to identify the biological mother for polygamous households. and sibling characteristics. To be specific, the variable ”Number of children” represents the observed number of children within each household  [^4] It does not include children who moved out of the family house before the census.. For the primary analysis, only households with children aged between 25 and 40 years are included. This age range is chosen to ensure that the children have either completed their education or nearly achieved complete educational attainment.
The inequality analysis focuses on households with at least 2 such children falling within the specified age range and at least one child with some educational attainment. This specific sample selection is motivated by the desire to examine the consequences of providing equal education to all children, as opposed to the alternative of not educating any children. The resulting sample comprises approximately $90,000$ individuals and $32,000$ households, serving as the basis for further investigation.
3.2 Key variables and measurement
The data set contains several key variables used in this paper, including gender, age, religion, area of residence, family size, household wealth index, and educational attainment of individuals, as well as their parents’ and a subset of their siblings’ variables. Apart from these variables, I also created measures for within household inequality, within household average years of education, and gender composition of children within a household. A description of each variable and their measurement is as follows:
Inequality: It refers to inequality in the education attainment of children within a given household. It is measured by the within household standard deviation and range of children’s education attainment. These two measures are used interchangeably throughout this paper.
High school graduate head of households/parents: It refers to head of households with at least 13 years of education. The education of the head of household is also used to describe the household. In other words, households with head of household with at least 13 years of education are referred to as high school graduate households.
Number of children: It is the total number of people who identify as children of a the head of household. This variable is denoted by $N_{c}$.
Within household average years of education: It is the average education of children between 25 and 40 years for a given household. It serves as a metric for accessing the average quality of children within the household. A related variable is the Within household total years of education, which is the simple sum of children’s years of education. It is used as a proxy for the household’s total investment in education. The within household total and average years of education of children are denoted by $q_{T}$ and $\bar{q}$ respectively.
3.3 Descriptive statistics
Table 1 displays the descriptive statistics for the key variables. Among the offspring, 38% are female, and their average education level is 8 years, with 80% having completed at least one year of education. Additionally, 40% of the heads of households have at least one year of education. About 80% of children without any schooling have parents who also lack formal education, whereas this percentage decreases to 50% for children with schooling. Conversely, approximately 29% of parents without schooling have children who likewise lack schooling, compared to only 8% for parents with schooling (See Figure 1). These statistics provide suggestive evidence of both inter-generational educational mobility  [^5] Children are more educated than parents and inter-generational educational persistence [^6] Children’s education is correlated with parents’ education..
4 Empirical Evidence
4.1 Within and Between Household Inequalities
In this section, I provide empirical evidence using a variance decomposition analysis to calculate the proportion of within-household variation in education and compared that to the overall variation in education. Through this analysis, I seek to understand how much of the educational disparities can be attributed to differences within households. Furthermore, I examine how within and between household variances in education are related. Let $q_{h}=(q_{h,1},q_{h,2},\dots,q_{h,N_{c}})$ be the vector of children’s education attainment in household $h$, and let $q=(q_{1},\dots,q_{n})$ be the education attainment of children in the sample.
$Var(q)=Var_{h}E[q|h]+E_{h}Var(q|h).$
The variance of $q$ is the sum of within and between household variation in $q$.
$\hat{E}_{h}\hat{Var}(q|h)=\hat{E}_{h}\hat{Var}(q_{h})=22.63\text{ and }\hat{Var}(q)=33.66$
This indicates that 67% of the variation in $q$ arises from variation within households. Furthermore, in the absence of within-household inequality, the between-household variance in children’s education attainment is 20.9. However, in the presence of within-household inequality, the between-household variance in children’s education attainment decreases to 11.2. These statistics suggest that, on average, households with some degree of within-household inequality exhibit lower between-household inequality compared to households with no within-household inequality. In conclusion, the analysis highlights on one hand the substantial contribution of within-household inequality to the overall inequality in education attainment. On the other hand, no within household variation in education of children is associated with higher between households variance. This suggests that households with greater disparities in education attainment among their members tend to exhibit lower disparities in education attainment when compared to other households.
4.2 Magnitude of within household inequality in children’s education
This section focuses on the extent of variation in within household inequality across households. I present evidence of differences in the magnitude of within household inequality across households. Understanding these differences can provide valuable insights into the factors that contribute to within household inequality and the potential mechanisms that can be employed to reduce it.
Figure 2 depicts the empirical distribution of the within household range and standard deviation of the education attainment of children. This figure provides compelling evidence of the variation in within household inequality in the educational attainment of children. Specifically, the data reveals that the magnitude of inequality varies across households, with some household having all of their children with the same education attainment while some have at least a child with some college education and at least a child with no education. These findings highlight the importance of considering household-level dynamics when addressing educational inequality and suggest that interventions aimed at reducing disparities in education must be tailored to the unique circumstances of each household. The inequality is present even within gender, although in lower magnitude. About $40\%$ (resp. $60\%$) of household has some level of inequality among daughters (resp. among sons).
The within household variance of children’s education is non-zero on average for all level of parents’ education and wealth index (see Figure 3). However, it appears that within household variance of children’s education decreases with parents’ education level and wealth index.
At the household level, a negative association emerges between the maximum education attainment within a household and the proportion of children within that household who have achieved this maximum education level. We can have of this through an OLS regression of the within household maximum years of education of children on the proportion of children with education attainment equal to that maximum.
$q^{\max}_{h}=\beta_{0}+\beta_{1}\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}1\{q_{i}=q^{\max}_{h}\}+\gamma^{\prime}X_{h}+\varepsilon_{h},$ (1)
where $X_{h}$ include number of children, HWI, area of residence, religion, gender composition of children, and head of household’ s education. The estimation results in column (3)- (4) of Table 2 indicates that, on average, households with a 0.01 higher proportion of children attaining the maximum years of education within the household tend to exhibit around 2.6 years lower overall maximum education levels for children within the household. Additionally, an OLS regression of the within household maximum years of education of children on the within household standard deviation of children’s education indicates that households characterized by higher levels of educational inequality demonstrate, on average, higher within-household maximum education attainment (see column (1)- (2) of Table 2). These findings suggest a trade-off involved in households’ education decision. The same argument as Becker and Lewis (1973) applies here, i.e. an increase in quality  [^7] Here quality refer to the within household maximum years of education of children is more expensive if there are more children with that quality. An increase in quantity [^8] Quantity refers to the number of children with the within household maximum years of education of children is more expensive if children are of high quality. This trade-off is a direct effect of the limited education resources available to households. In conclusion, due to financial constraints within the household, parents are facing a trade-off between reducing inequality within the household or reducing inequality between them and other households.
4.3 Parents’ education, average within household education of children, and inequality
The level of education attained by the head of a household has been found to be a significant factor in determining the level of inequality in children’s education attainment within that household. In particular, an increase in the head of household’s education level is associated with a decrease in inequality. However, it remains unclear whether this is a direct result of more educated parents’ aversion for inequality or an indirect result of their preference for education. To shed light on this issue, this section will investigate the factors that contribute to the observed negative correlation between parents’ educational attainment and within household inequality.
In addition to having lower level of inequality, households with more educated head of household also tend to have higher average years of education for their children (See Figure 5). This observation is particularly interesting given the hump-shaped relationship between inequality and average education of children (See Panel (b) of Figure 6). To investigate this relationship further, I estimation a OLS regression model of within inequality on average education of children, and parents’ level of education, with a quadratic interaction between between this two variables.
$Inequality_{h}=\alpha+\beta_{1}\bar{q}_{h}+\beta_{2}\bar{q}^{2}_{h}\beta_{3}hh\_Educ_{h}+\beta_{4}\bar{q}_{h}hh\_Educ_{h}+\beta_{5}\bar{q}^{2}_{h}hh\_Educ_{h}+\gamma^{\prime}X_{h}+\varepsilon_{h},$ (2)
where $X_{h}$ include number of children, HWI, area of residence, religion and gender composition. The estimation results in column (5)- (7) of Table 2 suggest that the negative dependence between parents’ education and within inequality is a result of both variable being correlated with the within household average education of children. In particular, the positive parents education between parents’ education and the within household average education of children combined with the hump shaped relation between within household inequality and the within household average education of children is translated into the observed spurious negative relationship between parents’ education and within household inequality.
4.4 Observed and unobserved sources of inequality
In the preceding sections, I have presented evidence at the household level, revealing that various factors contribute to the heterogeneity observed in the level of educational inequality across households. Notably, factors such as budget constraints, total investment in education, and the number of children play significant roles. By identifying and understanding these sources of heterogeneity, we gain valuable insights into the underlying mechanisms driving educational inequalities at the household level. In this section, the focus is on exploring the observed characteristics of children who received less education compared to their siblings. The examination of these characteristics is essential for developing effective strategies to address inequality and promote equality of opportunity for all children. Figure 7 graphs the average years of education based on the gender of children and the gender composition of households. To ensure accurate comparisons, the graph holds the within household average education of children constant. In the first panel, the analysis centers around households that are only able to finance primary school education for all their children. In the second panel, households that can only afford to provide education up to junior high school level are considered. The figure reveals some interesting trends in educational attainment among different gender composition households. Specifically, girls from only-daughter households, on average, have the same level of education as the household average, while boys from only-son households have similar education levels as well. However, in both-gender households, girls’ average education is lower than the household average, whereas boys’ average education is higher. These findings suggest that there is discrimination against daughters when it comes to the allocation of education quotas, when the alternative of giving more to a son is available.
Figure 8 allows similar analysis in terms of children’s birth order after controlling for number of children, and within household average years of education of children. The figure provides insight into the average years of education of first and second children from two children households. In panel (a) of Figure 8, the plot is for households that can afford to educate all their children up to primary education, and for households that can afford to educate all their children up to junior HS education is in panel (b). The figure demonstrates that the average years of education for the firstborn children is below the household average for both type of households, whereas the average years of education of the second-born children is above the household average. This monotonic increase in education by birth order applies to any family size (See Appendix). The findings of Figure 8 suggest that there is disadvantage in birth order regarding the allocation of education quotas.
In summary, a child’s gender, the gender and their siblings and their birth order are key determinants of the years of education they will receive. Despite taking into account observable household and children characteristics, a significant amount of variation in inequality across households remains unexplained, as evidenced by the $R^{2}$ value obtained from the regression of within household standard deviation of children’s education on those observable household and children characteristics (See the last two columns of Table 2). In addition, despite the presence of gender disadvantage against daughters, it appears that in some households, daughters receive higher education than their brothers (see Figure 9). I hypothesize that the unexplained difference in inequality can be attributed to the variance in children’s innate abilities, which differs across households. In other words, the fact that some daughters receive higher education compared to their brothers despite gender disadvantages can be attributed to high ability draws by these girls. This is a significant aspect of the household’s education distribution model, which I present in the next section.
4.5 Decomposition of Within Household Inequality
In the previous section, I have presented some empirical evidence about the observed characteristics of children which explain the within household inequality in their education. Additionally, it was demonstrated that a portion of this inequality can be attributed to the children’s unobserved abilities. In this section, I will provide a decomposition of the average within households inequality, categorizing it into components associated with gender disparity, birth order effects, and variations in children’s abilities. Such decomposition is conducted across various average educational levels within households on one hand and parents’ education level on the other hand. I used a household fixed-effect regression approach to achieve this breakdown.
Regression with Household Fixed Effects
To decompose the overall within household inequality into components categorized as disadvantage source and ability-based difference, I consider the following regressions:
$Educ_{i,h}=\beta_{1}Female_{i,h}+\beta_{2}Firstborn_{i,h}+\beta_{3}Female_{i,h}\times Firstborn_{i,h}+\nu_{h}+\varepsilon_{i,h}$ (3)
$Educ_{i,h}=\beta_{1}Firstborn_{i,h}+\nu_{h}+\varepsilon_{i,h}$ (4)
where $Educ_{i,h}$ is the years of education of child $i$ in household $h$, $Female_{i,h}$ is a gender indicator variable equals to 1 if child $i$ in household $h$ is a daughter, $Firstborn_{i,h}$ is a birth order indicator variable equals to 1 if child $i$ in household $h$ is a firstborn, and $\nu_{h}$ is the household fixed effect. Equation 3 is for households with both sons and daughters, while equation 4 is for households with either only sons or only daughters. The estimates from equation 3 and 4 are presented in Table 3 by average education of children and in Table 4 by parents’ education. The results suggest, on one hand, that about $63\%$ of the observed within household inequality in children’s education is due to gender and birth order disadvantages for households with both son and daughter. On the other hand, for households with only daughters or only sons, about 33% of the observed inequality is due to birth order disadvantages. This change is due to the fact that part of the ability-based inequality is muted by gender disadvantage. In other words, gender disadvantage not only exacerbates overall inequality but also diminishes the inequality driven by differences in ability.
For the primary analysis, which focuses on households with just two adult children living at home, the reliability of the estimates shown in Tables 3 and 4 may be compromised. This unreliability stems from the incidental parameter problem, a consequence of having only two data points per household for the fixed effect regressions. To validate the initial findings, I use the following alternative regression for a more robust examination.
$\Delta_{daughter-son}Educ_{h}=\beta_{0}+\beta_{1}Firstborn\_daughter_{h}+\varepsilon_{h},$ (5)
where $\Delta_{daughter-son}Educ_{h}$ is the average difference in the education of sons and daughters in household $h$, $Firstborn\_daughter_{h}$ is an indicator variable equals to 1 if the firstborn in household $h$ is a daughter. The estimates are summarized in Figure 11. Figure 11 illustrates the mean disparity in educational attainment between daughters and sons, with households having a firstborn male and female shown in red and blue, respectively. These measurements are provided across various average educational levels of the children in the panel a) and across education of the head of household in panel b), facilitating a decomposition for each average education level as follows:
[…]
and
[…]
Overall inequality
By combining equations 6 and 7, we can derive the average impact of gender and birth order disadvantages on the inequality within households at each level of the children’s average educational attainment and parents’ education. This integration enables us to break down the average inequality found within households into the three factors illustrated in Figure 12. This figure showcases how the average inequality is divided among gender disadvantage, birth order disadvantage, and differences in ability. It reveals that gender disadvantage is the predominant factor contributing to inequality. As the average educational level of children or the education of the head of household increases, the influence of ability differences becomes more significant, while the impact of birth order diminishes. Similarly, as parents’ education level increases, total inequality is smaller on average, and the share of birth order disadvantage reduces; but the share of gender disadvantage does not change significantly. This indicates not only the presence of variability in the degree of inequality across different levels of children’s average education and parents’ education but also in the way it is broken down.
Extensive versus Intensive margin inequality
It is relevant to analysis how within household inequality in education is decomposed for the extensive margin compared to the intensive margin. To analyze that, I run the previous fixed effect regression in equations 3 and 4 for households with only educated children— for the intensive margin analysis—, and the following regression for households with at least one non-educated child— for the extensive margin analysis.
$1\{Educ_{i,h}>0\}=\beta_{1}Female_{i,h}+\beta_{2}Firstborn_{i,h}+\beta_{3}Female_{i,h}\times Firstborn_{i,h}+\nu_{h}+\varepsilon_{i,h}$ (8)
The estimates are presented in Table 5, and the decomposition of inequality at extensive and intensive margins is presented in Figure 13. The numbers indicate that parents’ education is negatively related to inequality in children’s education mostly at the extensive margin. In particular, panel a) of Figure 13 shows the proportion of households with a non-educated child by parents’ education. That number is the highest among non-educated parents ($\approx 50\%$) and close to 00 ($\approx 3\%$) among college educated parents. There is also a substantial heterogeneity in the decomposition of inequality at the extensive margin. Specifically, for most of non-educated households with a non-educated child, the non-educated child is either a daughter or a firstborn. This is not true among college educated parents.
5 The Model
5.1 Setup
The model I have developed adopts a unique approach, considering children as investment goods rather than mere consumption goods. In other words, the number of children does not enter parents’ utility function directly like in Becker and Lewis (1973). The model consists of 2 distinct stages. In the first stage, households make decisions regarding the number of children, denoted as $N_{c}$, and observe their abilities, represented by the vector $\omega=\big{(}\omega_{1},\dots,\omega_{N_{c}}\big{)}$. They then choose the aggregate total years of education attainment, denoted as $q_{T}$, for these $N_{c}$ children. This leads to a within-household average years of education of children, represented as $\bar{q}=\frac{1}{N_{c}}q_{T}$. This initial stage can be viewed as choices derived from solving a fertility choice model, resembling the one described in Becker and Tomes (1976), with the distinction that each child is not assumed to receive $\bar{q}$ years of education. In other words, the decisions made in the first stage are based on the quantity-quality trade-off theory. This leads different choices for parents with different level of education. In addition to allowing an influence of parents’ education on the quantity of children and resources devoted to their schooling, I account for an unobserved heterogeneity that reflects parents’ aversion to having a child without education. This unobserved aversion parameter dictates the percentage of uneducated children within the family.
In the second stage of the model, households shift their focus to deciding on the distribution of $q_{T}$. This decision is dictated by the aversion parameter combines with children’s observed and unobserved characteristics. Specifically, each household is characterized by a type $\nu_{h}$ (their aversion to having an uneducated child). Given $\nu_{h}$, a household chooses the proportion of $N_{c}$ that receives some education, and distributes $q_{T}$ among those children taking into consideration their gender, birth order and innate ability. The decision of parents is to choose the distribution $(q_{1},\dots q_{N_{c}})$ of $q_{T}$, which maximizes the household’s utility function.
$\max_{q_{i}}U(q,\theta)$ (9)
$\text{subject to }\sum q_{i}\leq q_{T},\text{ }q_{i}>0,\text{ }q_{i}\leq q_{\max}$
$U(.)$ is increasing and concave, and $q_{\max}$ is the maximum years of education a child can receive. $\theta$ is described in the next section.
5.2 Functional form of households’ utility
Drawing upon the empirical evidence presented in Section 3, I have adopted the following generalized utilitarian social welfare function with concave utility function for each child to represent parents’ utility function. This form reflects the preferences and decision-making processes observed in the data and serves as a crucial component in the model. Let $q_{h}=(q_{1,h},\dots,q_{N_{c},h})$ be the distribution of $q_{T,h}$ in household $h$. The utility function for households with 2 children has the following expression [^9] The utility function for households with more than 2 children is presented in Appendix B.:
$U(q_{h},\theta)=\nu_{h}\Big{[}\sum_{i=1}^{N_{c}}a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i}\Big{]}+(1-\nu_{h})\Big{\{}\sum_{i=1}^{N_{c}}s_{h}\Big{[}\nu_{i}^{s}.a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i}\Big{]}+$ (10)
$\sum_{i=1}^{N_{c}}b_{h}^{d,s}\Big{[}\nu_{i}^{b_{d,s}}.a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i}\Big{]}+\sum_{i=1}^{N_{c}}b_{h}^{s,d}\Big{[}\nu_{i}^{b_{s,d}}.a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i}\Big{]}\Big{\}}$
where $a_{i}$ captures parents’ preference for child $i$ relative to other children based on their ability draws, and $\alpha_{i}$ is the costs (financial and opportunity costs) of giving a year of education to the $i^{th}$ child. Note that $(\alpha_{(1)},\dots,\alpha_{(N_{c})})$ is the ordering of cost of education by children’s birth order.
$a_{i}=\frac{\omega_{i}}{\sum_{j=1}^{N_{c}}\omega_{j}},\text{ }\nu_{i}^{gender\_comp}=1\{a_{i}.(q_{T})^{\delta_{i,h}}-\alpha_{i}q_{T}>a_{j}.(q_{T})^{\delta_{j,h}}-\alpha_{j}q_{T}\},\text{ }\nu_{j}^{gender\_comp}=1-\nu_{i}^{s},$
$gender\_comp\in\{s,b_{d,s},b_{s,d}\}.$
$s_{h}=Female_{i}.Female_{j}=(1-Female_{i}).(1-Female_{j})$
$b_{h}^{d,s}=Female_{1}.(1-Female_{2}),\text{ }b_{h}^{s,d}=(1-Female_{1}).Female_{2}$
For households with 2 children $\nu_{h}$ takes two possible values (High: H/Low:L). Households with high aversion for having an uneducated child choose to educate both children, whereas households with low aversion choose to educate only child out of the two children. For those who choose to educate only one child, choose the educated child based on a draw $\nu^{gender\_comp}_{i}$.
$\delta_{i,h}=\delta(gender_{i},gender\_comp_{h})=\gamma-\theta_{1}Female_{i}\frac{1}{N_{c}-1}\sum_{\{i,j\in h\},j\neq i}(1-Female_{j})$
$\nu_{h}\sim Bernoulli(p_{h}),\text{ }\nu_{1}^{s}\sim Bernoulli(p_{1}),\nu_{1}^{b_{d,s}}\sim Bernoulli(p_{fb,d}),\text{ }\nu_{2}^{b_{s,d}}\sim Bernoulli(p_{sb,d})$
The vector of parameters of interest is $\theta=(\theta_{1},\alpha_{1}-\alpha_{2},p_{1},p_{fb,d},p_{sb,d})$.
Assumption 1: The relative ability $a_{i}$ of a child $i$, is drawn from a distribution $G(.)$, with the constraint that $\sum_{i=1}^{n}a_{i,h}=1$. I assume that $G(.)$ is not dependent of gender or birth order.
The incorporation of differences in $\delta$ across children and the household’s gender composition within the model allows for the consideration of disadvantage that females face at the intensive margin in terms of human capital investment when they have a brother. This feature accounts for the varying investment in education based on the child’s gender, reflecting potential gender disadvantage that may exist within the household— as evidenced in Figure 7. The assumed functional form is designed to capture the idea that girls with brothers receive a penalty in the distribution decision of the education quota made by parents. Additionally that penalty is an increasing function of the proportion of boys among the siblings.
Similarly, the model’s incorporation of differences in $\alpha$ across children’s birth order enables the examination of the monotonic increase in education attainment as birth order advances. This acknowledges the tendency for education levels to rise sequentially with the child’s position within the family birth order— as observed in Figure 8. The integration of these elements into the model, allows for a more comprehensive understanding of the factors influencing educational outcomes, and facilitates the identification of potential areas for intervention to address gender-related disparities and birth order effects on educational attainment.
The parameters $p_{1}$, $p_{fb,d}$, and $p_{sb,d}$ allow to take into account disadvantages faced by firstborn children and daughters in their education attainment at the extensive margin. Finally, by incorporating the ability-based preference parameter into the model, I enable the consideration of unexplained inequality that may arise from factors beyond the observable household and individual characteristics. This parameter captures the influence of individual abilities, which is not captured by the observed variables.
To estimate the vector of parameter $\theta$, I used an indirect inference approach. In particular a simulated method of moments (SMM) approach. The procedure is outlined in the next section.
5.3 Inference Strategy
In this section, I provide an overview of the data moments utilized to identify the key parameters in the model. These data moments serve as essential empirical inputs that align the model’s predictions with the observed real-world outcomes.
For the estimation of the parameters, I use two set of moments. The first set of moments includes two moments. Firstly, I consider the difference in average education between girls from households composed of only daughters and households with a mix of genders, while controlling for parents’ education and the number of children ($N_{c}$). This difference in educational outcomes can be attributed solely to gender disadvantage within households, making it a critical determinant in identifying the parameters involved in $\delta$. By isolating the impact of gender disadvantage on educational disparities, this moment provides a valuable means of disentangling the specific effects of gender-related preferences within the model. Secondly, I consider the average education attainment by birth order in households with children of the same gender only (either only daughters or only sons), holding fix the years of education of the head of household, and the number of children in the household. These average educational levels for different birth orders are used to calculate the differences in average education attainment between successive birth orders. These differences play a fundamental role in uniquely identifying the parameter associated with the opportunity costs ($\alpha_{(i)}$).
The second set of moments used involves the proportion of educated daughters and the proportion of educated firstborn children by gender composition of households. This allows me to uniquely identify $p_{1}$, $p_{df,d}$, and $p_{sb,d}$. In particular the proportion of educated firstborn children in households with only children of the same gender, is used to estimate $p_{1}$. Similarly the proportion of educated firstborn daughters and second born daughters are used to estimate $p_{fb,d}$, and $p_{sb,d}$ respectively.
For the rest of the analysis, let’s define the variables $Y_{h}^{d}$ as daughters’ education in household $h$ and $Y_{h}^{s}$ as sons’ education in the same household $h$. And let’s $Y_{h}^{1}$, $Y_{h}^{2}$, be the education of firstborn and second children respectively. Additionally, let $Z$ represent a vector comprising the observables, specifically the education of the head of the household and the number of children ($N_{c}$). Note that households with identical observable characteristics share the same type, and any differences observed in the variables $Y_{h}^{d}$ and $Y_{h}^{s}$, or in $Y_{h}^{1}$ and $Y_{h}^{2}$ between these households stem from disparities in the unobservable difference in children’s ability. This unobservable component captures the variability in educational outcomes that cannot be explained by the observable characteristics, such as parental ability-based preferences that influence educational investment decisions within households.
[…]
  1. 1.
    $S_{d}^{d}=s\times H_{d}^{d}$ predictions of the education of daughters in households with only daughters,
  2. 2.
    $S_{d}^{b}=s\times H_{d}^{b}$ predictions of the education of daughters in households with both genders,
  3. 3.
    $S_{d}^{i}=s\times H_{d}$ predictions of the education of the $i^{th}$ born daughter in households with only daughters,
  4. 4.
    $S_{s}^{i}=s\times H_{d}$ predictions of the education of the $i^{th}$ born son in households with only sons,
These predicted education attainments represent the educational outcomes based on the given parameter values. I then take the average of the $S_{l}^{m}$ predictions for each moment, where $l,m\in\{s,d,b\}$ = {only sons, only daughters, both gender}.
To do inference on the parameters in $\delta$, the model and data moments are matched across various gender compositions. This process involves normalizing the parameter $a$ and estimating $\theta_{1}$ by comparing the model’s predictions to the observed data in terms of the difference in educational attainment for girls from households with only daughters and households with both genders.
To do inference on the parameters $(\alpha_{(t)})_{t=1}^{N_{c}}$ associated with birth order, the model and data moments are matched across different birth orders. This process entails normalizing $\alpha_{(N_{c})}$ and estimating $(\alpha_{(t)})_{t=1}^{N_{c}-1}$ by comparing the model’s predictions to the observed data regarding the difference in educational attainment between $t^{th}$ and $(t+1)^{th}$ born children.
To do inference on the parameters $p_{1}$, $p_{fb,d}$, and $p_{sb,d}$, the data moments and the model moments on the proportion of educated firstborn children, firstborn daughter, and second born daughters are matched with the model moments.
5.4 Estimation
In this section, I outline the estimation procedure used to estimate the parameters in the model. Let $\theta=(\theta_{1},\theta_{2},\theta_{3})$, where $\theta_{2}=(\alpha_{(1)},\dots,\alpha_{(N_{c}-1)})$, and $\theta_{3}=(p_{1},p_{fb,d},p_{sb,d})$. To estimate these parameters, I conduct $S$ simulations and compute the average education attainment of daughters in households with only daughters, only sons, and both genders as explained in the previous Section. Let $\hat{\mu}_{l}^{d}(\theta,Z)$ represent the average education attainment of daughters in these different household types, where $l\in{d,s,b}$ denotes households with only daughters, only sons, and both genders, respectively. Similarly, I compute the vector $\hat{\mu}(\theta,Z)=(\hat{\mu}_{1}(\theta,Z),\dots,\hat{\mu}_{N_{c}}(\theta,Z))$, which comprises the average education attainment by birth order. I can also compute $\hat{\pi}=(\hat{\pi}_{1},\hat{\pi}_{fb,d},\hat{\pi}_{sb,d})$, the model prediction of the proportions of firstborn children, of firstborn daughters and second born daughters. These simulations provide estimates of the model’s predictions for various household compositions, gender and birth orders, allowing for the comparison of the model’s outcomes with the observed data. The data moments are defined as follows:
5.4.1 Moments Matched
$m_{1}(Z)=E[Y^{d}|Z,Gender\_Comp=d]-E[Y^{d}|Z,Gender\_Comp=b],$
$m_{2}(Z)=E[1\{Y^{1}>0\}|Z,Gender\_Comp=\{d,s\}],$
$m_{3}(Z)=E[1\{Y^{d}>0\}|Z,Gender\_Comp=b,firstborn=d],$
$m_{4}(Z)=E[1\{Y^{d}>0\}|Z,Gender\_Comp=b,firstborn=s],$
$m_{t+4}(Z)=E[Y|Z,birth\_order=t+1]-E[Y|Z,birth\_order=t],\text{ }t\in\{1,\dots,N_{c}\}.$
I matched the following data and model moments to estimate $\theta$.
$m_{1}(Z)=\hat{\mu}^{d}_{d}(Z)-\hat{\mu}^{d}_{b}(Z),\text{ }m_{2}(Z)=\hat{\pi}_{1},\text{ }m_{3}(Z)=\hat{\pi}_{fb,d},\text{ }m_{4}(Z)=\hat{\pi}_{sb,d}\text{ and }$
$m_{t+4}(Z)=\hat{\mu}_{t+1}(Z)-\hat{\mu}_{t}(Z);\text{ }t\in\{1,\dots N_{c}-1\}.$
The corresponding sample objective function is the following expression:
$\hat{Q}(\theta|Z=z)=(\bar{Y}_{d,z}^{d}-\bar{Y}_{b,z}^{d}-(\hat{\mu}_{d,z}^{d}-\hat{\mu}_{b,z}^{d}))^{2}+(\hat{m}_{2}(Z)-\hat{\pi}_{1})^{2}+(\hat{m}_{3}(Z)-\hat{\pi}_{fb,d})^{2}+(\hat{m}_{4}(Z)-\hat{\pi}_{sb,d})^{2}+$ (11)
$\sum_{l\in\{d,s\}}\sum_{t=1}^{N_{c}-1}(\bar{Y}_{t+1,z}^{l}-\bar{Y}_{t,z}^{l}-(\hat{\mu}_{t+1,z}^{l}-\hat{\mu}_{t,z}^{l}))^{2}$
$\hat{\theta}=\text{ argmin}_{\theta\in\Theta}\hat{Q}(\theta).$
The sample objective function to possess a unique optimizer (See Figure 20).
5.5 Estimation of G(.)
I utilized auxiliary data to estimate the parameters of $G(.)$ outside of the model. In particular, I assume that the ability-based parents’ preference for children are i.i.d from a Beta distribution.
$a_{i}\sim^{i.i.d}Beta(\beta_{1},\beta_{2})\text{ with }\sum_{i=1}^{N_{c}}a_{i}=1$
$\beta_{1}$ and $\beta_{2}$ are estimated using auxiliary data. Specifically, I used data on average GPA in junior high school for a sample of student in Benin in 2018 to estimate $\beta_{1}$ and $\beta_{2}$ using maximum likelihood method. The Beta distribution seems to be a good fit for the distribution of relative ability (See Figure 21).
5.6 Standard Errors
The variance-covariance matrix for parameter estimates is given by the following expression:
$\Omega=\Big{[}\frac{\partial m(\theta)^{\prime}}{\partial\theta}\Big{|}_{\hat{\theta}}V^{-1}\frac{\partial m(\theta)}{\partial\theta}\Big{|}_{\hat{\theta}}\Big{]}^{-1}$
where, the partial derivative of the model function $\frac{\partial m(\theta)}{\partial\theta}$ is obtained numerically as,
$\frac{\partial m(\theta)}{\partial\theta}\Big{|}_{\hat{\theta}}=\frac{m(\hat{\theta}+h)-m(\hat{\theta})}{h}$
and, the variance-covariance matrix of the data moments, denoted as $V$, is computed through bootstrapping the data N times. This involves repeatedly calculating the moments N times and then using these N observations of the moments to determine the covariance between them. The standard errors are derived as the square roots of the diagonal elements of the matrix $\Omega$.
5.7 The Model’s Performance
In this section, I assess the effectiveness of stage 3 in the model through a straightforward exercise. This exercise consists of deriving the ability-based parameters $a_{i}$ from the model using $\hat{\theta}$. The empirical distribution of $\hat{a}$ should satisfy assumption 2 if the model has good performance. For a fix $\theta$ define the function
$g:a_{1},a_{2}\in(0,1)^{2}\longrightarrow[0,21]^{2}$
$(q_{1},q_{2})=g(\theta;a_{1},a_{2})$
Proposition 1: $g(\theta;.)$ is injective, i.e. $g(\theta;a_{1},a_{2})=g(\theta,a_{1}^{\prime},a_{2}^{\prime})$ implies that $(a_{1},a_{2})=(a_{1}^{\prime},a_{2}^{\prime})$.
This allows me to get the $\hat{a}_{i}$ for each children in each household. The distribution of $\hat{a}$ is plotted in Figure 15 for non-educated parents and Figure 16 for college educated parents. It shows that the distribution of $\hat{a}$ exhibits similarity across gender and birth order, albeit with minor differences. These disparities primarily stem from the heterogeneity in the disadvantages effects by different level of education, a feature not replicated in the predicted educational attainment by the model. To insure that this disparities do not affect the policy counterfactual analysis, I will use the distribution of $\hat{a}$ instead of $G(.)$ to do them.
6 Model predictions and Counterfactual Analysis
6.1 Results and Model Predictions
The estimates of $\theta$ are provided in Table 6 for households with $N_{c}=2$ and non-educated and college educated parents. The estimation results reveal that both at the extensive and intensive margins households with non-educated head of households exhibit higher values for the gender disadvantage parameter when compared to households with college educated parents. This is in line with the reduced form results, which suggest that college-educated parents are less biased against daughters in terms of education based on the comparison of educational outcomes between daughters and sons in households with college-educated parents versus non-educated parents. In particular, $\hat{p}_{fb,d}=0.1124$, $\hat{p}_{1}=0.3663$ and $\hat{p}_{sb,d}=0.3217$ for non-educated parents. Recall that:
$\nu_{1}^{fb,d}=1\{a_{1}.(q_{T})^{\delta_{1,h}}-\alpha_{1}q_{T}>a_{2}.(q_{T})^{\delta_{2,h}}-\alpha_{2}q_{T}\}\text{ and }\nu_{1}^{fb,d}\sim Bernoulli(p_{fb,d})$ (12)
Equation 12 implies that:
$\nu_{1}^{fb,d}=1\Big{\{}a_{1}>\frac{(\alpha_{1}-\alpha_{2})q_{T}+q_{T}^{\delta_{2,h}}}{q_{T}^{\delta_{1,h}}+q_{T}^{\delta_{2,h}}}\Big{\}}$
This implies that
$P[\nu_{1}^{fb,d}=1]=p_{fb,d}=P\Big{[}a_{1}>\frac{(\alpha_{1}-\alpha_{2})q_{T}+q_{T}^{\delta_{2,h}}}{q_{T}^{\delta_{1,h}}+q_{T}^{\delta_{2,h}}}\Big{]}$
Given the distribution of $a_{1}$, and $\hat{p}_{fb,d}$, we get:
$\frac{(\alpha_{1}-\alpha_{2})q_{T}+q_{T}^{\delta_{2,h}}}{q_{T}^{\delta_{1,h}}+q_{T}^{\delta_{2,h}}}=0.579$ (13)
For households with both gender and second born daughter, we have:
$\frac{(\alpha_{2}-\alpha_{1})q_{T}+q_{T}^{\delta_{1,h}}}{q_{T}^{\delta_{1,h}}+q_{T}^{\delta_{2,h}}}=0.53$ (14)
For households with only daughters or only sons, we have:
$\alpha_{1}-\alpha_{2}=\frac{1}{q_{T}}\Big{(}0.046\times q_{T}^{(1/2)}\Big{)}$ (15)
This suggests that, on average, for parents without formal education and low aversion for having an uneducated child, the likelihood of the firstborn child being educated compared to a second-born child of the same gender is approximately 0.3663, which correspond to an average cost difference of $0.019$. To illustrate, if it costs $2500$ USD  [^10] This is the estimated cost of completing high school in Benin in 2003 (Foko et al. (2012)) to provide a high school education for a second-born child, it would require an additional $47$ USD to provide the same level of education for the firstborn child in families without formal education. This difference represents around 13% of Benin’s GDP per capita in the 1990s.
Moreover, they perceive that investing in the education of the second-born son yields on average a 22% higher return compared to investing the same resources in the education of the firstborn daughter. Similarly, the return is estimated to be 27.8% higher for the firstborn son compared to the second-born daughter. Note that estimates for educational probabilities ($p_{1}$, $p_{fb,d}$, and $p_{sb,d}$) are not provided for college-educated parents, as nearly all of them, approximately 98%, have educated children.
At the intensive margin, $\hat{\theta}_{1}=0.0218$, and $\hat{\alpha}_{1}=0.0018$ for non-educated parents. For college educated parents, $\hat{\theta}_{1}=0.0115$, and $\hat{\alpha}_{1}=0.0016$ for college educated parents. This indicates that parents without formal education perceive a 5.75% higher return on graduating high school for sons compared to daughters, while for college-educated parents, the difference is approximately 3%. The cost difference between providing up to high school education for firstborn compared to second born conditional on both being educated represents only 1.3% of Benin’s GDP per capita in the 1990s for non-educated parents.
This disparity can be explained by a couple of key factors. First, college-educated parents often highly value education— they have an extreme aversion to having an uneducated child— and are more likely to prioritize educational opportunities for all their children, regardless of gender. They might view education as a means of empowerment and advancement, and thus, they are more likely to invest equally in the education of both daughters and sons. In addition, college-educated parents often have higher incomes and socioeconomic status, which can provide more resources for their children’s education. This increased financial stability can mitigate concerns about the perceived need to prioritize sons’ education over daughters’. Second, college-educated parents might be more aware of gender biases and stereotypes in society, including those related to education. They may actively work to counter these biases by ensuring equal opportunities for their children. Finally, there could be generational and cultural shifts at play, where younger, college-educated parents are more likely to challenge traditional gender roles and expectations regarding education. Overall, while these are potential reasons, it’s important to note that biases can still exist in any household, regardless of educational background. However, the data presented in the passage suggests a trend where college-educated parents appear to be less biased against daughters in terms of educational opportunities within their households.
6.2 Counterfactual Analysis
In this section, I provide contextual interpretations of the parameters within the model, shedding light on their real-world implications and the underlying mechanisms influencing distribution of education resources within households. I did this through two key counterfactual analyses, which allow me to explore the potential effects of policy interventions and alternative scenarios. These counterfactual analyses allow for a broader exploration of the model’s predictions under different conditions, helping to identify potential strategies to address educational inequalities and promote equitable educational opportunities for all children. The contextual interpretations and counterfactual analyses together contribute to a comprehensive evaluation of the model’s insights and provide valuable guidance for policymakers and researchers in devising effective measures to improve educational outcomes and reduce disparities within diverse household contexts.
6.2.1 Counterfactual 1: Gender, birth order, and ability
In this first counterfactual analysis, my main objective is to interpret the parameters $\theta$ by quantifying the additional level of ability required to counterbalance educational inequality resulting from gender and birth order disadvantages, respectively. To achieve this, I compute the extra ability needed by daughters and elder siblings to offset the effect of gender and birth order disadvantages on their education attainment. In order to do that I solve the household maximization problem in equation 9 with (using $\hat{\theta}$) and without disadvantages for a grid of relative ability of children for two-child families with a firstborn daughter and a second-born son, and compute the following quantity:
  1. 1.
    Ability of the firstborn daughter relative to the second born son at which the average difference between daughter’s and son’s education is equal to zero in presence of disadvantages.
  2. 2.
    The change in inequality due to gender and birth order disadvantages, by level of relative ability of the firstborn daughter.
Figure 14 presents the first counterfactual analysis for non-educated and college educated parents. It suggests three main conclusions. First, for the same ability draws, gender and birth order disadvantages reduce the education attainment of the first born daughter by $\approx 4.6$ and 2.2 for non-educated parents and college educated parents respectively. Second, the average difference between firstborn daughters and second born sons is equal to 0 in the presence of disadvantages when the firstborn daughter’s ability draw is $\approx 78\%$ (resp. 8%) higher than the ability draw of the second born son for households with non-educated head of household, (resp. households with college educated head of household).
6.2.2 Counterfactual 2: Education Policies
In situations where there is no inherent disadvantage against daughters or firstborn children, the primary factor leading to differences in education within a household between daughters and sons lies in variations in their individual abilities. Assuming that the initial abilities of children are distributed independently of their gender and birth order, the average difference in education between daughters and sons, in the absence of any disadvantages, forms a symmetric distribution centered around 0. However, this distribution shifts towards the negative side in instances where disadvantages are present. Simply put, in the absence of disadvantages, the distribution of the average education difference between daughters and sons exhibits first-order stochastic dominance over its counterpart in scenarios where disadvantages exist (See Figure 17).
The objective of this section is to identify an education cost reduction strategy that would result in an education distribution between daughters and sons resembling scenarios where no disadvantages exist, even in the presence of such biases. Additionally, I examine the efficacy of a policy focusing on the extensive margin, wherein all parents strongly prioritize avoiding uneducated children. The cost reduction policy encompasses the following measures:
  1. 1.
    At the extensive (resp. intensive) margin remove cost difference between firstborn and second born. $\approx 1.9\%$ (resp. $\approx 0.18\%$) reduction in schooling cost for firstborn children.
  2. 2.
    At the extensive (resp. intensive) margin reduced schooling cost for firstborn daughters. $\approx 3\%$ (resp. $\approx 1.3\%$) reduction in schooling cost for daughters compared to sons.
  3. 3.
    At the extensive (resp. intensive) margin reduced schooling cost for second daughters. $\approx 3.8\%$ (resp. $\approx 1.3\%$) reduction in schooling cost for daughters compared to sons.
Figure 18 displays the distribution of the average education difference between daughters and sons in panel (a), and between firstborn and second-born children of the same gender in panel (b), across four distinct scenarios. The blue curve represents the empirical distribution when disadvantages against daughters exist. Conversely, the red curve depicts the distribution in the absence of such biases. In the purple curve, we observe the distribution when disadvantages persist, but a cost reduction policy is implemented. Lastly, the orange curve illustrates the distribution when the disadvantage at the extensive margin is eradicated.
By construction, the cost reduction policy effectively mirrors the distribution in scenarios devoid of disadvantages. The elimination of the extensive margin disadvantage proves effective only when the ability draws of sons are smaller than the ability draws of daughters. This explained by the fact that high draws of ability by daughters can counterbalance the effect of gender disadvantage in the intensive margin. However, this balancing effect is not observed at the extensive margin. Consequently, in households where daughters demonstrate higher ability draws than their brothers, the elimination of disadvantages at the extensive margin alone is sufficient to align the distribution of average education differences between daughters and sons with scenarios where there is no gender disadvantage.
6.3 Counterfactual 3: Comparative Statics of Increase in Education Resources
This section delves into the analysis of how a non-targeted increase in total education resources among non-educated parents affects the impact of gender and birth order disadvantages on within-household inequality. Initial empirical findings reveal that the average education level of children from college-educated parents with two educated adult children is 14.5, while it stands at 9.2 for children from non-educated parents.
The exercise conducted in this section involves an exogenous increase in the average education level of children from non-educated parents, raising it from approximately 9.2 to 14.5. Subsequently, the household’s maximization problem is solved with this new average and $\hat{\theta}$. Finally, a comparison is drawn between the education differences of daughters and sons in this new scenario against the original disparity and the disparity observed for college-educated parents. The results are depicted in Figure 19. The Figure suggests that this non-targeted increase in education resources leads to increase in inequality against daughters both compared to the original disparities and compared to college educated parents. In summary, the analysis suggests that the non-targeted increase in education resources, while aiming to improve overall educational outcomes, had unintended consequences of widening the gap in educational attainment between daughters and sons within households where parents are not educated.
7 Robustness of Estimates to Missing Siblings
7.1 Gender Disadvantage
The main dataset used for this analysis comprises adult children who were living in the same household as their parents during the census period. This sample represents a specific subgroup within the larger population of adult children. Importantly, the decision for children to leave the parental home is often influenced by factors such as their occupation and educational accomplishments, making it an endogenous process. Moreover, the motives for leaving home frequently differ between daughters, commonly associated with marriage, and sons. Given these dynamics, there’s a potential for bias in our estimates. This would be particularly concerning if, firstly, the children who remained at home are more similar to each other, and secondly, if they significantly differ from those who moved out.
The wide range in both educational attainment and gender among children residing in the same household as their parents suggests that the first concern may not be significant. However, the second concern could lead to either overestimation – if women who moved out are more educated and men who moved out are less educated – or underestimation – if women who moved out are less educated and men who moved out are more educated, compared to those who remained at home.
In this section, I delve into the potential bias in estimating the effect of gender disadvantage on within-household inequality. To investigate this, I compare the educational attainment of adult women and men living in the same households as their parents to those who have moved out. The mean comparison between these two groups is presented in Table 9. This comparison suggests that the difference in average education between men and women is more pronounced in the sub-sample that is not included in my analysis. As a result, it implies that, if anything, I may be underestimating the effect of gender disadvantage. Consequently, my estimate of gender disadvantage can be interpreted as an estimate of the lower bound of the true parameter.
7.2 Birth Order Disadvantage
A similar argument to the one presented in the previous section also applies to the birth order disadvantage parameter. The decision for children to move out is closely linked to their age, with older children being more inclined to leave their parents’ household. Consequently, we may have a selected sample of younger children in some households. In specific cases, children referred to as firstborns in certain households might actually be of a higher birth order. Additionally, more accomplished younger siblings may have already moved out. It’s important to note that both of these situations would potentially bias our estimate of the birth order disadvantage parameter downward.
8 Conclusion
In this paper, I examine the interaction between the three empirically known sources of disparities in children’s educational attainment within households. I constructed a structural model of households’ distribution of education resources among children, allowing for the influence of factors such gender, birth order and ability of children. The model not only allows me to decompose, for each relative ability draw, the total observed inequality into parts due gender, birth order, and ability differences; it also gives a platform for analysing how education cost reduction policies affects within household inequalities.
The construction of the model is motivated by contexts similar to the one of Benin; a setting marked by notable disparities in children’s education within households, coupled with evidence of gender and birth order disadvantages. To ensure tractability, certain aspects of the parental decision-making process regarding education resources distribution are omitted. Notably, the model adopts a static approach, although the education decision of children is inherently dynamic. The primary objective of the paper being to rationalize the observed differences in children’s education, attributing them to gender disadvantages, birth order disadvantages, or variations in innate ability draws; despite its static nature, the model proves relevant, as it effectively incorporates and analyzes the interactions among these three factors. Additionally, the paper attributes any unexplained differences in children’s education, not accounted for by gender and birth order, to differential draws of innate ability. However, it acknowledges the potential influence of other unobserved factors, such as varying preferences of mothers in polygamous households, which could lead to increased parental investment in the education of specific children. In recognizing this, the interpretation of unexplained inequality within households is acknowledged as an upper bound of the effect of differential ability.
In light of the findings in this paper, we can expect a reduction in the opportunity cost of girls education such as education support in the form of cash transfers, scholarships, and school kits for girls; to reduce within household inequality in children’s education that is due to gender disadvantage. Additionally, a reduction in the opportunity cost of education for firstborn, such as cash transfers and school kits, to young parents (first-time parents) or scholarships for firstborn children; is expected to reduce within household inequality in children’s education that is due to birth order disadvantage. However, these two policies need to be combined for an effective reduction in disadvantaged-based inequality. This is due to the possibility of displacement of disadvantage from one group to another. In particular, if the policy only targets firstborn children, the disadvantage against daughters might increase, and vice versa.
9 Tables and Figures
(a) Children’s education as function of parents’ education
(b) Parents’ education as function of children’s education
Figure 1: Parents and children’s education.
(a) Range
(b) Standard Deviation
Figure 2: Empirical cdf of within household range and standard deviation of education attainment of children.
(a) Parents Education
(b) Household Wealth Index (HWI)
Figure 3: Distribution of inequality by socio-economic groups.
Figure 4: Distribution of inequality by number of children.
Figure 5: Within household average and standard deviation of children’s education as function of parents’ level of education.
Figure 6: Distribution of average education attainment of children.
Figure 7: Average years of education by gender and households gender composition ($N_{c}=2$)
Figure 8: Average years of education by birth order ($N_{c}=2$)
(a) Between daughters and sons
(b) Between first and second born
Figure 9: Histogram of within household difference in average education (Benin, 2013)
(a) By gender
(b) By birth order
Figure 10: Distribution of children’s education for number of $N_{c}=2$ (Benin, 2013)
(a) As function of within household average education of children
(b) As function of head of household’s education
Figure 11: Effect of gender and birth order disadvantages on within household inequality ($N_{c}=2$)
(a) As function of within household average education of children
(b) As function of head of household’s education
Figure 12: Inequality decomposition ($N_{c}=2$)
(a) Extensive Margin
(b) Intensive margin
Figure 13: Inequality decomposition as function of head of household’s education ($N_{c}=2$)
(a) Non-educated head of household
(b) College educated head of households
Figure 14: Effect of gender and birth order disadvantages on inequality ($N_{c}=2$)
(a) Firstborn vs. second born
(b) Daughter vs. son
Figure 15: Distribution of estimated ability-based parameters ($N_{c}$ = 2 and non-educated parents)
(a) Firstborn vs. second born
(b) Daughter vs. son
Figure 16: Distribution of estimated ability-based parameters ($N_{c}$ = 2 and college educated parents)
(a) Daughter v.s son
(b) Firstborn v.s second born
Figure 17: Distribution of the difference in children’s education ($N_{c}=2$)
(a) Daughter v.s son
(b) Firstborn v.s second born
Figure 18: Distribution of the difference in children’s education ($N_{c}=2$)
Figure 19: Distribution of difference in daughter and son’s education by parents’ education level
Figure 20: Plot of $\hat{Q}(\theta)$
Figure 21: Histogram of relative GPA in junior high school and histogram of random draws from Beta (28.82, 28.78).
Tables
Table 1: Descriptive Statistics
Statistic N Mean St. Dev. Min Max
Age 89,594 29.452 3.993 25 40
Female 89,594 0.380 0.485 0 1
Years of education 89,594 7.760 5.802 0 21
At least one years of education 89,594 0.776 0.417 0 1
Range of children’s education 89,594 6.821 4.873 0 21
Standard deviation of children’s education 89,594 3.874 2.744 0.000 14.142
Educated head of household 85,407 0.390 0.488 0 1
Number of children between 25 and 40 89,594 3.055 1.557 2 16
Number of children 89,594 6.340 4.354 2 79
Educated with educated head of household 85,407 0.353 0.478 0 1
Non-educated with non-educated head of household 89,594 0.180 0.384 0 1
$\bar{q}$ 89,594 7.760 4.444 0.143 20.250
$q_{T}$ 89,594 22.354 15.660 1 148
Table 2: Regression of within household maximum years of education on within household inequality and of within household standard deviation of children’s education on households’ characteristics
Maximum years of education Standard deviation
(1) (2) (3) (4) (5) (6) (7)
(Intercept) $9.40^{*}$ $6.16^{*}$ $11.94^{*}$ $10.24^{*}$ $4.24^{*}$ $1.46^{*}$ $0.87^{*}$
Standard deviation $0.50^{*}$ $0.66^{*}$
$\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}1\{q_{i}=q_{\max}\}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-1.35^{*}}$ ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}-2.64^{*}}$
[…]
hh Educ = Junior HS $2.62^{*}$ $2.14^{*}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1.17^{*}}$ $-0.66^{*}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}0.86^{*}}$
hh Educ = Senior HS $4.24^{*}$ $3.56^{*}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1.59^{*}}$ $-0.47^{*}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}2.98^{*}}$
hh Educ = College $5.74^{*}$ $4.93^{*}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}-1.94^{*}}$ $0.24^{*}$ ${\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}4.75^{*}}$
Average years of education ($\bar{q}$) $0.91^{*}$ $1.16^{*}$
$\bar{q}^{2}$ $-0.06^{*}$ $-0.07^{*}$
hh Educ = Primary:$\bar{q}$ $-0.29^{*}$
hh Educ = Junior HS:$\bar{q}$ $-0.56^{*}$
hh Educ = Senior HS:$\bar{q}$ $-0.94^{*}$
hh Educ = College:$\bar{q}$ $-1.14^{*}$
hh Educ = Primary:$\bar{q}^{2}$ $0.02^{*}$
hh Educ = Junior HS:$\bar{q}^{2}$ $0.04^{*}$
hh Educ = Senior HS:$\bar{q}^{2}$ $0.05^{*}$
hh Educ = College:$\bar{q}^{2}$ $0.06^{*}$
Number of children $0.04^{*}$ $0.05^{*}$ $0.04^{*}$ $0.05^{*}$
HWI $0.45^{*}$ $0.38^{*}$ $-0.10^{*}$ $-0.10^{*}$
Urban $0.89^{*}$ $0.71^{*}$ $-0.32^{*}$ $-0.32^{*}$
Christian $0.90^{*}$ $0.74^{*}$ $-0.28^{*}$ $-0.27^{*}$
Both gender $0.31^{*}$ $0.42^{*}$ $0.20^{*}$ $0.22^{*}$
R2 $0.10$ $0.31$ $0.00$ $0.16$ $0.04$ $0.22$ $0.24$
Num. obs. $32729$ $32729$ $32729$ $32729$ $32729$ $32729$ $32729$
Null hypothesis value outside the confidence interval.
Table 3: Regression of children’s education on their gender and birth order with household fixed effect $(N_{c}=2)$
(1) (2) (3) (4) (5) (6)
$q_{T}=12$ $q_{T}=20$ All $q_{T}$
Female $-3.03^{*}$ $-2.75^{*}$ $-2.46^{*}$
First born $-3.24^{*}$ $-1.90^{*}$ $-2.59^{*}$ $-2.61^{*}$ $-0.95^{*}$ $-1.24^{*}$
Firstborn female $1.26^{*}$ $0.38$ $-0.27^{*}$
R2 $0.21$ $0.09$ $0.11$ $0.09$ $0.67$ $0.70$
Adj. R2 $-0.59$ $-0.82$ $-0.79$ $-0.83$ $0.34$ $0.40$
Num. obs. $1632$ $300$ $1558$ $278$ $43970$ $7562$
RMSE $4.39$ $3.91$ $5.71$ $5.76$ $4.52$ $4.23$
Household fixed effects
Average inequality (Both gender: Firstborn female) 6.23 7.22 5.84
Average inequality (Only daughters) 3.81 6.01 4.59
Average inequality (Only sons) 5.07 6.09 5.08
Explained proportion
Gender 50.1% - 38.1% - 33.7% -
Birth order 30.5% 49.9% 35.9% 43.4% 29.3% 32.9%
Unexplained 19.6% 50.1% 26% 56.6% 37% 67.1%
Null hypothesis value outside the confidence interval.
Note: Columns (2), (4), and (6) are for households with only daughters. For households with only sons the decomposition is 19% birth order + 81% ability.
Table 4: Regression of children’s education on their gender and birth order with household fixed effect $(N_{c}=2)$
(1) (2) (3) (4) (5) (6)
Non-educated College educated All
parents parents
Female $-3.16^{*}$ $-0.90^{*}$ $-2.47^{*}$
Firstborn $-1.19^{*}$ $-1.55^{*}$ $-0.41$ $-0.13$ $-0.93^{*}$ $-1.24^{*}$
Firstborn female $-0.39^{*}$ $0.25$ $-0.34^{*}$
R2 $0.58$ $0.60$ $0.68$ $0.69$ $0.67$ $0.70$
Num. obs. $22540$ $3528$ $1884$ $478$ $40884$ $6956$
Household fixed effects
Average inequality (Both gender: Firstborn female) 6.76 3.27 5.84
Average inequality (Only daughters) 5.29 3.16 4.59
Average inequality (Only sons) 5.71 2.78 5.08
Explained proportion
Gender 47.2% - 36% - 33.7% -
Birth order 23% 29.3% 4% 4.1% 29.3% 32.9%
Unexplained 29.8% 70.7% 60% 95.9% 37% 67.1%
Null hypothesis value outside the confidence interval.
Note: Columns (2), (4), and (6) are for households with only daughters. For households with only sons the decomposition is respectively 17% birth order + 83% ability for college educated parents and 21% birth order + 79% ability for non-educated parents. For the whole sample it is 18% birth order + 82% ability.
Table 5: Regression of children’s education on their gender and birth order with household fixed effect (Extensive vs Intensive margin) $(N_{c}=2)$
Non-educated parents College educated parents All
Extensive Intensive Extensive Intensive Extensive Intensive
(1) (2) (3) (4) (5) (6)
Female $-0.54^{*}$ $-1.73^{*}$ $0.05$ $-0.94^{*}$ $-0.52^{*}$ $-1.38^{*}$
Firstborn $-0.23^{*}$ $-0.39^{*}$ $0.11$ $-0.47^{*}$ $-0.22^{*}$ $-0.29^{*}$
Firstborn Female $-0.05$ $0.05$ $-0.35$ $0.45$ $-0.05$ $-0.25^{*}$
R2 $0.21$ $0.71$ $0.03$ $0.71$ $0.20$ $0.75$
Num. obs. $10166$ $12374$ $62$ $1822$ $12846$ $28038$
Household fixed effects
Null hypothesis value outside the confidence interval.
Note:
Table 6: Estimates of $\hat{\theta}$
Non-educated parents College educated parents
$\hat{\theta}_{1}$ $\hat{\theta}_{2}$ $\hat{p}_{1}$ $\hat{p}_{fb,d}$ $\hat{p}_{sb,d}$ $\hat{\theta}_{1}$ $\hat{\theta}_{2}$
Estimates 0.0218 0.0018 0.3663 0.1124 0.3217 0.0115 0.0016
Standard errors 0.0017 0.005 0.0145 0.0119 0.0136
Number of observations 11,270 942
Table 7: Data and Model Moments: Targeted Moments
Data Moments Model moments Difference
$\bar{Y}_{d}^{d}-\bar{Y}_{d}^{b}$ 1.7028 1.7169 -0.0141
Non-educated parents $\bar{Y}_{sb}^{d}-\bar{Y}_{fb}^{d}$ 0.2754 0.4502 -0.1748
$\bar{Y}_{sb}^{s}-\bar{Y}_{fb}^{s}$ 0.3682 0.1616 0.2066
Proportion of educated firstborn 0.3642 0.3678 0.0036
Proportion of educated firstborn daughters 0.1116 0.1156 0.004
Proportion of educated second born daughters 0.3177 0.3177 0
$\bar{Y}_{d}^{d}-\bar{Y}_{d}^{b}$ 1.56 1.55 0.01
College educated parents $\bar{Y}_{sb}^{d}-\bar{Y}_{fb}^{d}$ 2.61 2.76 -0.15
$\bar{Y}_{sb}^{s}-\bar{Y}_{fb}^{s}$ 2.47 2.88 -0.41
Proportion of educated firstborn
Proportion of educated firstborn daughters
Proportion of educated second born daughters
Table 8: Data and Model Moments: Non targeted Moments
Data Moments Model moments Difference
Average inequality 5.87 5.97 -0.1
Average inequality (only daughters households) 5.29 5.46 -0.17
Average inequality (only sons households) 5.71 5.87 -0.16
Average inequality (both gender households) 6.29 6.29 0
Average inequality ($\bar{q}<6$) 5.05 5.04 0.01
Non-educated parents Average inequality ($106\leq\bar{q}<$ 10) 8.46 7.56 0.9
Average inequality ($1310\leq\bar{q}<$ 13) 5.27 5.34 -0.07
Average inequality ($\bar{q}\geq 13$) 2.1 5.62 -3.52
Average inequality (intensive margin) 3.59 3.77 -0.18
Standard deviation of firstborn’s education 5.55 5.5
Standard deviation of second born education 5.19 5.29
Average inequality
Average inequality (only daughters households)
Average inequality (only sons households)
Average inequality (both gender households)
Average inequality ($\bar{q}<6$)
College educated parents Average inequality ($106\leq\bar{q}<$ 10)
Average inequality ($1310\leq\bar{q}<$ 13)
Average inequality ($\bar{q}\geq 13$)
Average inequality (intensive margin)
Standard deviation of firstborn’s education
Standard deviation of second born education
Table 9: Difference in Average education of men and women
Observed Unobserved
Average years of education
Male 5.35 4.66
Female 4.73 2.03
Difference 0.62 2.63
Proportion of married
Male 0.37 0.84
Female 0.38 0.88
Difference 0.01 -0.04
Quantity-quality trade-off in Benin
Table 10: Regression of education attainment dummy on number of children and other covariates
1998-2013 1970-1998 1960-1970
OLS IV OLS IV OLS IV
(Intercept) $0.74^{*}$ $0.71^{***}$ $0.44^{*}$ $0.46^{***}$ $0.16^{*}$ $0.62^{*}$
$[0.74;0.75]$ $(0.01)$ $[0.43;0.44]$ $(0.02)$ $[0.12;0.19]$ $(0.26)$
Family size:Urban $-0.01^{*}$ $-0.01^{***}$ $-0.00^{*}$ $-0.00$ $0.00$ $0.01$
$[-0.01;-0.01]$ $(0.00)$ $[-0.00;-0.00]$ $(0.01)$ $[-0.01;0.01]$ $(0.06)$
Family size:Rural $-0.02^{*}$ $-0.02^{***}$ $-0.00^{*}$ $-0.01^{*}$ $0.00$ $-0.16$
$[-0.03;-0.02]$ $(0.00)$ $[-0.00;-0.00]$ $(0.00)$ $[-0.00;0.01]$ $(0.10)$
Education of hh $0.02^{*}$ $0.02^{***}$ $0.03^{*}$ $0.03^{***}$ $0.05^{*}$ $0.05^{***}$
$[0.02;0.02]$ $(0.00)$ $[0.03;0.03]$ $(0.00)$ $[0.05;0.06]$ $(0.00)$
Female hh $0.09^{*}$ $0.10^{***}$ $0.11^{*}$ $0.11^{***}$ $0.15^{*}$ $0.03$
$[0.09;0.09]$ $(0.00)$ $[0.11;0.12]$ $(0.01)$ $[0.13;0.17]$ $(0.09)$
HWI $0.04^{*}$ $0.04^{***}$ $0.06^{*}$ $0.06^{***}$ $0.07^{*}$ $0.07^{***}$
$[0.04;0.04]$ $(0.00)$ $[0.06;0.06]$ $(0.00)$ $[0.06;0.08]$ $(0.01)$
Urban $0.02^{*}$ $0.04^{***}$ $0.13^{*}$ $0.11^{***}$ $0.16^{*}$ $-0.36$
$[0.02;0.02]$ $(0.01)$ $[0.12;0.13]$ $(0.03)$ $[0.12;0.19]$ $(0.30)$
Birth order $0.02^{*}$ $0.01^{***}$ $0.01^{*}$ $0.02^{***}$ $0.01$ $0.12$
$[0.01;0.02]$ $(0.00)$ $[0.01;0.01]$ $(0.00)$ $[-0.01;0.03]$ $(0.08)$
Female $-0.04^{*}$ $-0.04^{***}$ $-0.08^{*}$ $-0.08^{***}$ $-0.12^{*}$ $-0.14^{***}$
$[-0.04;-0.04]$ $(0.00)$ $[-0.09;-0.08]$ $(0.00)$ $[-0.14;-0.10]$ $(0.02)$
R2 $0.14$ $0.14$ $0.21$ $0.21$ $0.27$ $-0.05$
Adj. R2 $0.14$ $0.14$ $0.21$ $0.21$ $0.27$ $-0.05$
Num. obs. $2069182$ $2069182$ $403850$ $403850$ $6461$ $6461$
RMSE $0.40$ $0.43$ $0.40$
${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$ (or Null hypothesis value outside the confidence interval).
10 Appendix A: Additional Figures
(a) All Children.
(b) Children between 25-40 years old.
Figure 22: Histogram of number of children.
(a) Whole sample
(b) Sample of educated individuals
Figure 23: Histogram of years of education.
Figure 24: Distribution of within household range of children’s education.
Figure 25: Average education by birth order for households with $N_{c}=2$ (Benin, 2013)
Figure 26: Average education by birth order for households with $N_{c}=3$ (Benin, 2013)
Figure 27: Average education by birth order for households with $N_{c}=4$ (Benin, 2013)
Figure 28: Average education by birth order for households with $N_{c}=5$ (Benin, 2013)
Figure 29: Average education by birth order for households with $N_{c}=6$ (Benin, 2013)
(a) 2 daughters, 1 son
(b) 2 sons, 1 daughter
Figure 30: Effect of gender and birth order disadvantages on within household inequality as function of head of household’s education ($N_{c}=3$)
(a) 2 daughters, 1 son
(b) 2 sons, 1 daughter
Figure 31: Effect of gender and birth order disadvantages on within household inequality as function of head of household’s education ($N_{c}=3$)
11 Appendix B: Households’ utility function for Number of Children equals 3
The utility function maximized by households with 3 children has the following expression:
$U(q_{h},\theta)=\nu_{h}\Big{[}\sum_{i=1}^{N_{c}}a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i}\Big{]}+\nu_{m}\Big{\{}\sum_{i=1}^{N_{c}}s_{h}\Big{[}\nu_{m,i}^{s}.(a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i})\Big{]}+$ (16)
$\sum_{gender\_comp}\sum_{i=1}^{N_{c}}b_{h}^{gender\_comp}\Big{[}\nu_{m,i}^{{gender\_comp}}.(a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i})\Big{]}\Big{\}}+\nu_{l}\Big{\{}\sum_{i=1}^{N_{c}}s_{h}\Big{[}\nu_{l,i}^{s}.(a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i})\Big{]}+$
$\sum_{gender\_comp}\sum_{i=1}^{N_{c}}b_{h}^{gender\_comp}\Big{[}\nu_{l,i}^{{gender\_comp}}.(a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i})\Big{]}\Big{\}}$
where
$gender\_comp\in\{(d,s,s);(d,d,s);(d,s,d);(s,d,d);(s,s,d);(s,d,s)\}$
$s_{h}=Female_{i}.Female_{j}=(1-Female_{i}).(1-Female_{j})$
$b_{h}^{d,s}=Female_{1}.(1-Female_{2}),\text{ }b_{h}^{s,d}=(1-Female_{1}).Female_{2}$
For households with 3 children, aversion for having uneducated children takes three possible values (High: H/ Medium:M/ Low:L). Households with high aversion for having an uneducated child choose to educate both children. Households with medium aversion choose to educate 2 of their children, whereas households with low aversion choose to educate only child out of the three children. For those who choose to educate two children and only one child, choose the educated children based on a draw $\nu^{gender\_comp}_{m,i}$ and $\nu^{gender\_comp}_{l,i}$.
$\delta_{i,h}=\delta(gender_{i},gender\_comp_{h})=\gamma-\theta_{1}Female_{i}\frac{1}{N_{c}-1}\sum_{\{i,j\in h\},j\neq i}(1-Female_{j})$
$\nu_{h}\sim Bernoulli(p_{h}),\text{ }\nu_{m}\sim Bernoulli(p_{m}),\text{ }\nu_{l}\sim Bernoulli(p_{l}),\text{ with the constraint that }\nu_{h}+\nu_{m}+\nu_{l}=1$
In practice, for each simulated household, the algorithm is as follows:
  1. 1.
    Take draw for $\nu_{h}\sim Bernoulli(p_{h})$,
  2. 2.
    If $\nu_{h}=1$, $\nu_{m}=\nu_{l}=0$ for that household,
  3. 3.
    If $\nu_{h}=0$, take draw for $\nu_{m}\sim Bernoulli(p_{m})$,
  4. 4.
    If $\nu_{m}=1$, $\nu_{l}=0$ for that household,
  5. 5.
    If not $\nu_{l}=1$. Note that this work because by construction in the data $\hat{p}_{h}+\hat{p}_{m}+\hat{p}_{l}=1$.
$\nu_{m,i}^{s}\sim Bernoulli(p_{m,i}^{s}),\text{ }\nu_{l,i}^{s}\sim Bernoulli(p_{l,i}^{s})\text{ with }i\in\{1,2,3\}\text{ and }\sum_{i=1}^{3}\nu_{m,i}^{s}=2\text{ and }\sum_{i=1}^{3}\nu_{l,i}^{s}=1$
Note that $p_{m,i}$ and $p_{l,i}$ are parameters of interest to be estimated. In practice for households with medium aversion, for each simulated household, the algorithm is as follows:
  1. 1.
    Take a draw for $\nu_{m,1}^{s}\sim Bernoulli(p_{m,1})$,
  2. 2.
    If $\nu_{m,1}^{s}=1$, take a draw for $\nu_{m,2}^{s}\sim Bernoulli(p_{m,2})$,
  3. 3.
    If $\nu_{m,2}=1$, $\nu_{m,3}=0$. If not $\nu_{m,3}=1$.
For households with low aversion, the algorithm is similar to the one for $\nu_{h}$, $\nu_{m}$ and $\nu_{l}$.
$\nu_{m,1}^{b_{gender\_comp}}\sim Bernoulli(p_{m}^{gender\_comp}),\text{ }\nu_{m,2}^{b_{gender\_comp}}\sim Bernoulli(p_{m}^{gender\_comp}),$
$\nu_{m,3}^{b_{gender\_comp}}\sim Bernoulli(p_{m}^{gender\_comp})\nu_{l,1}^{b_{gender\_comp}}\sim Bernoulli(p_{l}^{gender\_comp}),\text{ }$
$\nu_{l,2}^{b_{gender\_comp}}\sim Bernoulli(p_{l}^{gender\_comp}),\nu_{l,3}^{b_{gender\_comp}}\sim Bernoulli(p_{l}^{gender\_comp})$
with the following conditions on $\nu_{m,i}^{gender\_comp}$s, and $\nu_{l,i}^{gender\_comp}$s.
$\sum_{i=1}^{3}\nu_{m,i}^{gender\_comp}=2\text{ and }\sum_{i=1}^{3}\nu_{l,i}^{gender\_comp}=1$
The vector of parameters of interest is
$\theta=(\theta_{1},\alpha_{(1)}-\alpha_{(2)},\alpha_{(2)}-\alpha_{(3)},p_{m,1}^{s},p_{m,2}^{s},p_{l,1}^{s},p_{l,2}^{s},p_{m,1}^{gender\_comp},p_{m,2}^{gender\_comp},p_{l,1}^{gender\_comp},p_{l,2}^{gender\_comp})$
and
$p_{m,3}^{s}=1-(p_{m,1}^{s}+p_{m,2}^{s}),\text{ }p_{l,3}^{s}=1-(p_{l,1}^{s}+p_{l,2}^{s})$
$p_{m,3}^{gender\_comp}=1-(p_{m,1}^{gender\_comp}+p_{m,2}^{gender\_comp}),\text{ }p_{l,3}^{gender\_comp}=1-(p_{l,1}^{gender\_comp}+p_{l,2}^{gender\_comp})$
The dimension of $\theta$ is $1\times 31$.
12 Appendix C: Model Performance
For a fix $\theta$ define the function
$g:a_{1},a_{2}\in(0,1)^{2}\longrightarrow[0,21]^{2}$
$(q_{1},q_{2})=g(\theta;a_{1},a_{2})$
Proposition 1: $g(\theta;.)$ is injective, i.e. $g(\theta;a_{1},a_{2})=g(\theta,a_{1}^{\prime},a_{2}^{\prime})$ implies that $(a_{1},a_{2})=(a_{1}^{\prime},a_{2}^{\prime})$.
Proof of Proposition 1
Let $g(\theta;a_{1},a_{2})=g(\theta;a_{1}^{\prime},a_{2}^{\prime})=(x,q_{T}-x)$. From the household’s maximization we must have $\forall x^{\prime}\neq x$
$a_{1}x^{\delta_{1}}+a_{2}(q_{T}-x)^{\delta_{2}}-\alpha_{1}x-\alpha_{2}(q_{T}-x)\geq a_{1}x^{\prime\delta_{1}}+a_{2}(q_{T}-x^{\prime})^{\delta_{2}}-\alpha_{1}x^{\prime}-\alpha_{2}(q_{T}-x^{\prime})$
and
$a_{1}^{\prime}x^{\delta_{1}}+a_{2}^{\prime}(q_{T}-x)^{\delta_{2}}-\alpha_{1}x-\alpha_{2}(q_{T}-x)\geq a_{1}^{\prime}x^{\prime\delta_{1}}+a_{2}^{\prime}(q_{T}-x^{\prime})^{\delta_{2}}-\alpha_{1}x^{\prime}-\alpha_{2}(q_{T}-x^{\prime})$
This implies that
$(a_{1}+a_{1}^{\prime})x^{\delta_{1}}+(a_{2}+a_{2}^{\prime})(q_{T}-x)^{\delta_{2}}\geq(a_{1}+a_{1}^{\prime})x^{\prime\delta_{1}}+(a_{2}+a_{2}^{\prime})(q_{T}-x^{\prime})^{\delta_{2}}$ (17)
which is not necessarily true if $(a_{1},a_{1})\neq(a_{1}^{\prime},a_{2}^{\prime})$. Hence $g(.)$ is injective.
To see that equation 17 only holds if $(a_{1},a_{1})\neq(a_{1}^{\prime},a_{2}^{\prime})$, consider the following:
  • $N_{c}=2$, $q_{T}=20$, $Female_{1}=0$ and $Female_{2}=1$.
  • Initial ability vector correspond to $(0.57,0.43)$ with optimal distribution of $q_{T}$ equal to $q_{1}=15$, $q_{2}=5$
  • The blue curve correspond to utility with change in ability but no change in education distribution.
  • The red curve correspond to utility with change in ability and the corresponding optimal distribution of education.

### v3

@@
 Abstract
-This paper investigates the nuanced interplay of gender, birth order, and innate ability in shaping educational disparities among children within households, employing both a reduced-form analysis and a structural model of household resource allocation.
-By decomposing overall disparities, I identify the contributions of gender, birth order, and innate ability differences.
-In the context of Benin, for households with non-educated heads and mixed-gender children, total inequality comprises 50% gender disparity, 20% birth order effect, and 30% ability gap.
-Conversely, in households with only sons or only daughters and non-educated heads, total inequality is predominantly driven by ability disparities—70% for daughters and 83% for sons, with lesser contributions from birth order effects.
-Furthermore, my analysis reveals that in households with non-educated heads, firstborn daughters surpass their younger brothers in educational attainment on the intensive margin if their innate ability exceeds their brothers’ by at least 13%, a figure reduced to 8% in households with college-educated parents.
-Additionally, the study unveils parental preferences favoring sons’ education over daughters’ among non-educated parents, with a perceived 22% higher average benefit.
-Targeted policies aimed at reducing composite education costs prove effective in mitigating gender and birth order gaps, albeit with a modest 5% reduction in total inequality.
-Overall, this research underscores the complex dynamics influencing intra-household educational inequalities and suggests policy avenues to address them.
+This paper examines how gender, birth order, and innate ability shape within-household disparities in children’s educational attainment in developing countries.
+Using data from Benin, I find that in households with non-educated parents, gender and birth order drive over two-thirds of the average educational attainment disparities among adult children, while their influence decreases to one-third in households with college-educated parents.
+Furthermore, average inequality, measured by the range of children’s educational attainment is twice as high among non-educated parents compared to college educated parents.
+I propose and estimate a structural model of educational attainment choices within-family.
+Using the model, I show that the absence of gender and birth order effects does not lead to a significant reduction in the average within-family disparities in children’s educational attainment.
+Additionally, in theory, ensuring that every child has at least one year of education lowers average within-family educational inequality.
+Yet, even in this scenario, daughters tend to receive less education than sons, and practical efforts to achieve universal entry are less effective than the theoretical model.
+The latest version of this paper is available here.
 JEL classification: I220, I240, H520
@@
 1 Introduction
-The trade-off between the number of children one chooses to have and the quality of those children has been a subject of extensive research.
-This exploration began with the seminal contributions of De Tray (1970) and Becker and Lewis (1973).
-The quantity-quality trade-off model delves into how changes in the quantity of children can impact their average quality.
-Numerous empirical studies have provided support for this model in various contexts, consistently revealing a negative relationship between the number of children, often used as an empirical measure of quantity, and the educational attainment of those children, typically used as an empirical measure of quality.
-These findings align with the predictions put forth by the theoretical framework presented by Becker and Lewis (1973) and have been substantiated by research such as Montgomery (1995) and Li et al.
-(2008).
+The trade-off between the number of children one chooses to have and the average education of those children has been a subject of extensive research.
+The quantity-quality trade-off model predicts a negative relationship between the quantity and the average quality of children.
+Numerous empirical studies have provided support for this model in various contexts, consistently revealing a negative relationship between the number of children, and the average educational attainment of those children (Becker and Lewis (1973), Montgomery (1995) and Li et al.
+(2008)).
 The current state of the literature on fertility choices (Conley and Glauber (2006), Maralani (2008), Li et al.
@@
 (2019)) states that rich and educated families tend to have fewer children while allocating greater investments in the education of those children, in comparison to less affluent and less educated families.
-In other words, the average level of education among children varies across different families, influenced by endogenous factors such as family size and exogenous factors like the socio-economic background of the parents.
-When we extend the analysis beyond the choice of the aggregate quality of children to how it is distributed among children, we observe that, not only does the average change across households, but so does the variance of quality.
-In addition, the variance is most likely non-zero for a majority of households when there are budget constraints and no compulsory education laws.
-This presumption of a non-zero variance is justified by evidence of the influence of factors such as gender, birth order and innate ability on education attainment of children.
-The educational outcomes of children are strongly influenced by their individual characteristics, leading to disparities in the amount of education they receive.
+However, in contexts where fertility can not be controlled by parents— low access to contraceptive methods— or when the quality of and access to public education system is poor or there is no clear understanding of the economic benefits of schooling, this trade-off seems to vanish (Montgomery (1995), Black et al.
+(2005), Maralani (2008)).
+When we extend the analysis beyond the choice of the aggregate education of children to how it is distributed among children within a family, we evidence that, not only does the average change across households, but so does the variance of education.
+In addition, the variance is most likely non-zero for a majority of households when there are budget constraints and no or poorly enforced compulsory education laws.
+In such circumstances, the educational outcomes of children are strongly influenced by their observed and unobserved individual characteristics, leading to disparities in the amount of education they receive.
 Several studies have shed light on these disparities, with findings indicating that girls tend to receive less education due to factors such as gender bias or gender preference (as documented by Biswas (2000) and Ota and Moffatt (2007)).
@@
 (2020).
-Furthermore, children with higher abilities are more likely to receive increased educational opportunities and education attainment, as suggested by the research of Becker and Tomes (1976), Dizon-Ross (2019), and Giannola (2023).
-However, there exists a notable gap in the research landscape, as there is a scarcity of studies that comprehensively analyze all these various sources of disparities within the same analytical framework.
-Such a framework, capable of simultaneously examining gender-based differences, birth order effects, and the impact of individual abilities on educational inequality within households, holds the potential for interesting counterfactual and education policies analyses.
+Furthermore, children with higher abilities are more likely to receive increased educational opportunities and educational attainment, as suggested by the research of Becker and Tomes (1976), Dizon-Ross (2019), and Giannola (2024).
+However, there exists a notable gap in the research landscape, as there are few studies that comprehensively analyze all these various sources of disparities within the same analytical framework.
+Such a framework, capable of simultaneously examining gender-based differences, birth order effects, and the impact of individual unobserved abilities on educational inequality within-households, holds the potential for the analysis of the effectiveness of educational counterfactual scenarios in reducing inequality.
 Specifically, it enables us to explore how these factors interact with one another to shape inequality.
-Furthermore, this comprehensive approach is crucial for estimating the proportion of inequality attributable to disadvantages and the portion caused by ability-based resource allocation within households.
-The former allows us to design education policies which target inequality due gender or birth order disadvantages.
-In this paper, I use reduced form approach and suggest a structural model to break down the overall intra-household differences in children’s education into parts related to gender, birth order, and ability.
-Firstly, the model helps us find out if the total disparities observed is solely due to biases from parents and society or because there is high variation in children’s abilities.
-Secondly, it allows us to measure how much more intelligent a daughter (or firstborn) needs to be in order to have the same educational opportunities as a son (or second born).
-Lastly, using this model, we can examine how differences in education of children within a family change with a cost reduction policy compared to a policy which aim to change parents’ preference.
-I estimate the model using data from Benin in West Africa.
-Benin is a good place to apply this model for three main reasons.
-First, even though there is a law stipulating that everybody must finish at least primary school, it’s not always followed by parents or enforced by the government.
-Second, there are big differences in how much education individuals in the same household get.
-Also, these differences can be very different from one family to another.
-Lastly, there is evidence that girls and firstborns often don’t get the same education opportunities as others.
-For this exercise, I focused on households where there were only two adult children still residing with their parents.
-Among those households there is difference in how much education resources they have to distribute and the education attainment of the head of household.
-Taking that into account, I applied my method to households with different characteristics separately.
-My analysis uncovered three important findings.
-First, when it comes to educating daughters compared to their brothers within a household with a non-educated head of household, large part (70%) of the differences in the education of brothers and sisters is due to disadvantages rather than variations in abilities.
-That number is 33% among college educated parents.
-However, when examining the education of firstborns compared to their siblings of the same gender, more than half (70% for female and 80% for male) of the disparities are attributed to differences in abilities.
-These results suggests that in some households gender disadvantage reduces disparities in children’s education through the reduction of the extent to which the talent of a child affect their education.
-This implies that removing/reducing gender and birth order disadvantages will not systematically reduce inequality for all households.
-It will increase the amount of disparities in some households.
-Second, When we reduce education costs for firstborn children and daughters, we can decrease the differences in education within certain families.
-This is especially true in families where the firstborn children and daughters have lower abilities compared to their siblings or where parents don’t really take the children’s abilities into consideration when making decisions about education.
-However, if the firstborn children and daughters are actually more capable than their siblings and parents do consider their abilities, a policy that eases the financial burden of schooling for them might unintentionally increase educational differences within the family.
-This brings up a dilemma about finding the right balance between giving everyone equal educational opportunities, regardless of factors like gender, birth order, and ability, and ensuring that children with similar abilities get the same opportunities, regardless of gender and birth order.
-In the latter case we are allowing for rewarding high abilities but not for penalizing gender and birth order.
-In terms of policy changes, it seems more implementable to aim for the latter option.
-That means focusing on policies that increase the net return from education for firstborn children and daughters either through schooling cost reduction or reduction of salary gap between male and female.
-These policies will help decrease the impact of gender and birth order on educational differences.
-On the other hand, it’s trickier to implement a policy that reduces how much parents consider a child’s abilities in distributing the few resources that they have, when facing financial challenges.
-Especially given that it is more costly to give high level formal education to a low ability child, and that there are other more suitable alternatives such as apprenticeship.
-Third, When all kids in a family are considered equally capable, the differences in education between them created by gender and birth order is at its highest point.
-This happens because parents see the benefit from their investment in education and the child’s ability as complementary.
-In other words, parents think that spending more on the education of a child with higher ability will bring greater benefits compared to investing in a child with lower ability.
-So, if the firstborn or a daughter has lower abilities than their siblings, they will receive less investment in education even without considering gender and birth order disadvantages.
-This means that the disadvantages they face because of gender and/or birth order have a smaller impact on increasing the overall disparities in education in term of magnitude.
-On the flip side, if the firstborn or a daughter is more capable than their siblings, they would get more investment in their education compared to their siblings, in the absence of gender and birth order disadvantages.
-As a result, the disadvantage they face due to gender and birth order leads to a decrease in the educational differences within the family.
-This decrease becomes less pronounced as their intelligence level increases.
+Furthermore, this comprehensive approach is crucial for estimating the proportion of inequality attributable to gender and birth order and the portion caused by other unobserved factors such as the cognitive ability of a child compared to their siblings.
+The former allows us to design educational policies which target inequality due to gender or birth order effects.
+This micro-level analysis of educational disparities is particularly relevant to the broader context of educational challenges in Sub-Saharan Africa, where lack of education remains a critical obstacle to development.
+Despite numerous reforms, the region continues to struggle with high illiteracy rates, significant educational inequality, and have a substantial proportion of the world’s out-of-school children (UNESCO (2021)).
+While primary school completion rates are approaching or exceeding 90% in most regions globally, Sub-Saharan Africa lags behind, with only two out of three children completing primary school (UNESCO (2021)).
+Educational disparities are the results of various barriers, which can be categorized into three types: situational (life circumstances), dispositional (personal attitudes), and institutional (structural conditions) (UNESCO (2021)).
+A crucial but often overlooked factor contributing to these disparities is the within-family inequality in children’s educational attainment, which accounts for approximately 40% of the variation in educational attainment in the developing world (Giannola (2024)).
+Using this regional evidence of educational inequality as a foundation, this paper focuses on the specific context of Benin.
+Benin is an ideal location for this study for three main reasons: the non enforcement of compulsory education laws, the significant variability in educational attainment among individuals within the same household, and the observed disparities in educational opportunities based on gender and birth order.
+For this exercise, I focused on households where there are two adult children residing in the same household as their parents [^3] The extension to households with more than 2 children is straightforward.
+All the estimates and analysis in the paper are also done after including households with 3 children and the results are presented in the appendix of the paper..
+Among those households there are differences in how much educational resources parents have to distribute and the education attainment of the head of household.
+Taking that into account, I perform my analysis on households with different observed characteristics separately.
+In the first part of the paper, I establish two key stylized facts about intra-household educational inequality.
+First, there is a non-linear, hump-shaped relationship between household-level mean and variance of educational attainment.
+Second, I decompose the average within-household educational attainment disparities into gender, birth order, and unobserved residual effects, finding that among households with non-educated heads and one child of each gender, over two-thirds of the average disparities in children’s educational attainment is due to gender and birth order, while among college-educated parents, only one-third is due to these factors.
+Furthermore, average inequality, measured by the range of children’s educational attainment is twice as high among non-educated parents compared to college educated parents.
+I then propose a structural model of household educational resources allocation to rationalize the observed inequality with budget constraints and parents’ preferences.
+The model accounts for gender effects by introducing a difference in the marginal utility of education between sons and daughters, with the marginal utility being higher for sons when birth order and other unobserved factors are held constant.
+Additionally, birth order effects are integrated through variations in the marginal cost of education by birth order, holding gender and unobserved factors fixed.
+Each child’s unique unobserved characteristics that influence their educational outcomes are modeled through an unobserved weight on the benefit derived from that child’s education.
+I estimate the model using the Simulated Method of Moments, and analyze diverse counterfactual scenarios in how they affect average inequality and the share of gender and birth order effects in the average inequality among non-educated parents.
+The first counterfactual (1) which theoretically set gender and birth order effects to zero; did not lead to a significant reduction in the average inequality in the sample.
+The second counterfactual (2) which theoretically removed barriers to school entry for all children reduces the average inequality in the sample by about half, but does not reduce the share of gender effect.
+The third counterfactual (3) which is more practical, relaxes the financial burden of non-educated parents to the same level as their college-educated counterparts.
+This only leads to a 10% reduction in the average inequality in the sample, with more than 50% of the remaining inequality due to gender effects.
 The subsequent sections of this paper are structured as follows.
-In Section 2, I present the literature review along with the specific contributions of this paper.
-Section 3 offers an overview of the data utilized for the estimation of the model, while Section 4 presents key empirical observations and facts derived from the analysis.
-Section 5 is dedicated for the model’s setup, outlining the identification strategy for key parameters, and describing the inference and estimation procedures employed in this study.
-Lastly, in Section 6, I present counterfactual analysis to further explore the implications of my findings.
-Section 7 presents a robustness analysis of the key parameters of interest, and Section 8 concludes the paper.
-2 Literature Review
-This paper contributes to the extensive literature on fertility choice models and the quantity-quality trade-off faced by households within a country.
-Early developed theoretical models (DeTray (1970), Becker and Lewis (1973)) predicts that quantity and quality of children are negatively related due to the observation that the shadow price of quantity is influenced by quality, and vice versa.
-More specially, they have shown that a reduction in the quantity of children results in a decrease in the shadow price of quality.
-Consequently, this decline in the shadow price of quality leads to an increase in the overall quality of the children.
-The current state of the literature on fertility choices (Montgomery (1995), Conley and Glauber (2006), Maralani (2008), Li et al.
-(2008), Weng et al.
-(2019)) states that rich and educated families tend to have fewer children while allocating greater investments in the education of those children, in comparison to less affluent and less educated families.
-This trend reflects a significant pattern where socio-economic factors play a crucial role in shaping fertility decisions and educational investments within households.
-However, some studies still found a positive or no correlation between fertility and education of children either for rural or old households (Montgomery (1995), Black et al.
-(2005), Maralani (2008)).
-This is often explained by the lack of contraception methods and poor public education systems.
-This paper contributes to that strand of the literature by offering two key contributions.
-Firstly, it empirically demonstrates the existence of a quantity-quality trade-off within the context of Benin.
-Second, it builds on this result to relax the assumption of equal education for children in the same household.
-This paper also contributes to the literature on the within-household schooling decision, particularly factors influencing parents’ distribution decision of education resources among siblings.
-One key determinant of these distribution decisions is the gender or gender composition of households.
-Previous research has shown that daughters are less likely to receive education or have lower educational attainment on average.
+In Section 2, I present an overview of the data used for this study, while Section 3 presents key empirical evidence and stylized facts derived from the data description.
+Section 4 is dedicated for the model’s setup, outlining the estimation strategy for key parameters, and describing the inference and estimation procedures employed in this study.
+Lastly, in Section 5, I present counterfactual analysis to further explore the implications of my findings, and Section 7 concludes the paper.
+Other relevant literature
+This paper contributes to the literature on within-household schooling decision, particularly factors influencing parents’ distribution decision of education resources among siblings.
+A key determinant of these distribution decisions is the gender of the child and the gender composition of the household.
+Previous research has shown that daughters are less likely to receive some education; or have lower educational attainment on average.
 Studies have shown that, while the presence of elder sisters tends to increase the likelihood of schooling, the presence of younger brothers may decrease it (Biswas (2000), Ota and Moffatt (2007), Ombati and Ombati (2012), Osadan and Burrage (2014), Psaki et al.
@@
 Finally, a child’s innate ability or talent plays a role in parental distribution decisions.
-Studies have demonstrated that parents invest more in the human capital of endowed children and allocate more nonhuman capital to less endowed children (Becker and Tomes (1976), Dizon-Ross (2019), Giannola (2023)).
-When parents are compelled to invest in the nonhuman capital of low-ability children, this leads to an inefficient equilibrium, where the investment in the human capital of high-ability children is not optimized (Nerlove et al.
+Studies have demonstrated that parents invest more in the human capital of high-ability children and allocate more nonhuman capital to low-ability children (Becker and Tomes (1976), Dizon-Ross (2019), Giannola (2024)).
+When parents are compelled to invest in the nonhuman capital— for example inheritance in form of land or financial assets— of low-ability children, this leads to an inefficient equilibrium, where the investment in the human capital of high-ability children is not optimized (Nerlove et al.
 (1984)).
 This paper adds to this existing literature in two significant ways.
-Firstly, it examines a context where parents are not constrained to compensate lower ability children by investing in their nonhuman capital but, instead, rely on family taxes (Wantchekon et al.
+First, it examines a context where parents are not constrained to compensate lower ability children by investing in their nonhuman capital but, instead, rely on family taxes (Wantchekon et al.
 (2015)).
-Secondly, this paper develops a household resources distribution structural model that allows for a more flexible analysis of the distribution of education resources within the household.
-In this model, the assumption of equal distribution is relaxed, enabling a detailed exploration of the interactions between gender and, birth order disadvantages, and the innate abilities of individual children in influencing parental distribution decisions.
-Finally, this paper contributes to the literature on within household inequality in children’s human capital (Giannola (2023)).
-Giannola (2023) has shown in the context of India that observed inequality within households is partly explained by parents investing more in the human capital of high-achieving children, especially when they are financially constrained.
-This behavior stems from the fact that parents are not particularly averse to inequality and tend to reinforce the gap in learning created by innate ability rather than correct it.
+Second, this paper proposes a household educational resources distribution model which allows for a more flexible analysis of the distribution of education resources within the household.
+In this model, the assumption of equal distribution is relaxed, enabling a detailed exploration of the interactions between gender and, birth order effects, and the innate abilities of children in influencing household distribution decisions.
+This paper also contributes to the literature on educational Kuznets curve theory ( Londoño (1990) and Ram (1990), Thomas et al.
+(2003), Morrisson and Murtin (2013)).
+Previous studies have analyzed the relationship between the mean and variance of education using cross country data or within-country time series data.
+This paper contributes to that literature by analyzing the relationship between the mean and variance of education using within-country cross household data.
+Specifically, it shows that in Benin at the household level; the relationship mean-variance of education is inverted U shaped with the peak occurring at approximately 7 years of education..
+Finally, this paper contributes to the literature on within-household inequality in children’s human capital (Giannola (2024)).
+Giannola (2024) has shown in the context of India that observed inequality within-households is partly explained by parents investing more in the human capital of high-achieving children, especially when they are financially constrained.
+This behavior stems from the fact that parents are not particularly averse to inequality and tend to reinforce the gap in learning created by innate ability rather than correcting it.
 This paper contributes to that literature by first building upon the result that parents unequally invest in the human capital of high-achieving children in contexts where the education system is better tailored to serve high-achieving students.
-Second, this paper interacts with that result and examines how it relates to other sources of unequal distribution, such as gender and birth order.
+Second, this paper interacts with that result and examines how it relates to other sources of inequality, such as gender and birth order.
 2 Data Description and Definition of Key Variables
-In this section, I present the data used to analyze the question of interest in this paper.
-The context of Benin is particularly compelling for analyzing the interaction between gender disadvantage, birth order advantage, and innate ability in shaping within household inequality for three reasons.
-First, there are a lot of disparities in education attainment within households due to the lack of compulsory education law coupled with financial constraints and the availability of outside options other than formal education such as child labor within the households and apprenticeship.
-Second, there is factual evidence of gender and birth order disadvantages.
-Lastly, there is observational evidence that despite the prevalence of gender disadvantages in the population, daughters in some households get higher education than their brothers.
 2.1 Sample and Data
-The 2013 Population and Habitation Census data of Benin; used in this paper; offers a comprehensive insights into households and their members living in the country during that year.
-Conducted by the National Bureau of Statistics of Benin, this census provides data at both the household and individual levels.
+In this section, I present the data used.
+I use data from the 2013 Population and Habitation Census of Benin.
+This census data provides information on all households and their members.
+It is conducted by the National Bureau of Statistics of Benin, and has information at both the household and individual levels.
 For the purpose of this paper, the focus is directed towards individuals who identify themselves as the children of household heads, enabling to get information on parental  [^4] Parents here refers to one of the parents, either the mother or the father.
 This because it is not possible to have both for household with single parents and to identify the biological mother for polygamous households.
-and sibling characteristics.
-To be specific, the variable ”Number of children” represents the observed number of children within each household  [^4] It does not include children who moved out of the family house before the census..
+and sibling characteristics for a sub-sample of siblings.
+The variable “Number of children” represents the observed number of children within each household  [^5] It does not include children who moved out of the family house before the census..
 For the primary analysis, only households with children aged between 25 and 40 years are included.
-This age range is chosen to ensure that the children have either completed their education or nearly achieved complete educational attainment.
+This age range is chosen to ensure that the children have either completed their education or nearly achieved maximum educational attainment.
+The summary statistics of all individuals between 25 and 40 years living in the same household as their parents is presented in Table 9 in Appendix A.
 The inequality analysis focuses on households with at least 2 such children falling within the specified age range and at least one child with some educational attainment.
-This specific sample selection is motivated by the desire to examine the consequences of providing equal education to all children, as opposed to the alternative of not educating any children.
-The resulting sample comprises approximately $90,000$ individuals and $32,000$ households, serving as the basis for further investigation.
+This specific condition on the sample is motivated by the goal to examine the reasons for providing equal education to all children, as opposed to the alternative of not educating any children.
+In particular households with only non-educated children do not offer any information about the distribution of education resources which this paper aims to analyze.
+The resulting sample comprises approximately $90,000$ individuals and $\approx 33,000$ households, serving as the basis for further investigation.
+The sample description is as follows:
+  1. 1.
+    Sample 0: All households with children between 25-40 yrs old ($\approx 160,000$ households)
+  2. 2.
+    Sample 1: Households in sample 0 with at least 2 children between 25 and 40 years old. ($\approx$ 51,600 households)
+  3. 3.
+    Sample 2: Households in sample 1 with at least one educated child between 25 and 40 years old. ($\approx$ 32,800 households)
+I use sample 1 for stylized facts, estimation of the model, counterfactual analysis, and comparative statics, and sample 2 only for counterfactual analysis, and comparative statics.
 2.2 Key variables and measurement
 The data set contains several key variables used in this paper, including gender, age, religion, area of residence, family size, household wealth index, and educational attainment of individuals, as well as their parents’ and a subset of their siblings’ variables.
-Apart from these variables, I also created measures for within household inequality, within household average years of education, and gender composition of children within a household.
+Apart from these variables, I also created measures for within-household inequality, within-household average years of educational attainment of children, and gender composition of children within a household.
 A description of each variable and their measurement is as follows:
-Inequality: It refers to inequality in the education attainment of children within a given household.
-It is measured by the within household standard deviation and range of children’s education attainment.
-These two measures are used interchangeably throughout this paper.
-High school graduate head of households/parents: It refers to head of households with at least 13 years of education.
-The education of the head of household is also used to describe the household.
-In other words, households with head of household with at least 13 years of education are referred to as high school graduate households.
+Within household inequality: Is the disparity in the educational attainment of children within a given household.
+It is measured by the within-household range of children’s education attainment for households with 2 children.
+For households with more than 2 children, it is measured by the standard deviation of education.
 Number of children: It is the total number of people who identify as children of a the head of household. This variable is denoted by $N_{c}$.
-Within household average years of education: It is the average education of children between 25 and 40 years for a given household.
-It serves as a metric for accessing the average quality of children within the household.
-A related variable is the Within household total years of education, which is the simple sum of children’s years of education.
+Within-household average years of educational attainment of children: Is the average education of children between 25 and 40 years for a given household.
+It serves as a metric for accessing the average quality of children within the household, and is used as proxy for educational resources available.
+A related variable is the Within-household total years of educational attainment of children, which is the simple sum of children’s years of education.
 It is used as a proxy for the household’s total investment in education.
-The within household total and average years of education of children are denoted by $q_{T}$ and $\bar{q}$ respectively.
+The within-household total and average years of educational attainment of children are denoted by $q_{T}$ and $\bar{q}$ respectively.
 2.3 Descriptive statistics
 Table 1 displays the descriptive statistics for the key variables.
-Among the offspring, 38% are female, and their average education level is 8 years, with 80% having completed at least one year of education.
-Additionally, 40% of the heads of households have at least one year of education.
+First, among the observed offspring [^6] Observed offspring refers to adult children living in the same household as their parents.
+in sample 2, 38% are female, their average education level is 8 years, with 80% having completed at least one year of education.
+Second, 40% of heads of household have at least one year of education.
+Table 1: Descriptive Statistics
+Statistic Mean St. Dev. Min Max
+Age 29.452 3.993 25 40
+Female 0.380 0.485 0 1
+Years of education 7.760 5.802 0 21
+At least one years of education 0.776 0.417 0 1
+Range of children’s education 6.821 4.873 0 21
+Standard deviation of children’s education 3.874 2.744 0.000 14.142
+Educated head of household 0.390 0.488 0 1
+Number of children between 25 and 40 3.055 1.557 2 16
+Number of children 6.340 4.354 2 79
+$\bar{q}$ 7.760 4.444 0.143 20.250
+$q_{T}$ 22.354 15.660 1 148
+Number of observation 89,594
 About 80% of children without any schooling have parents who also lack formal education, whereas this percentage decreases to 50% for children with schooling.
-Conversely, approximately 29% of parents without schooling have children who likewise lack schooling, compared to only 8% for parents with schooling (See Figure 1).
-These statistics provide suggestive evidence of both inter-generational educational mobility  [^5] Children are more educated than parents and inter-generational educational persistence [^6] Children’s education is correlated with parents’ education..
-4 Empirical Evidence
-4.1 Within and Between Household Inequalities
-In this section, I provide empirical evidence using a variance decomposition analysis to calculate the proportion of within-household variation in education and compared that to the overall variation in education.
-Through this analysis, I seek to understand how much of the educational disparities can be attributed to differences within households.
-Furthermore, I examine how within and between household variances in education are related.
-Let $q_{h}=(q_{h,1},q_{h,2},\dots,q_{h,N_{c}})$ be the vector of children’s education attainment in household $h$, and let $q=(q_{1},\dots,q_{n})$ be the education attainment of children in the sample.
-$Var(q)=Var_{h}E[q|h]+E_{h}Var(q|h).$
-The variance of $q$ is the sum of within and between household variation in $q$.
-$\hat{E}_{h}\hat{Var}(q|h)=\hat{E}_{h}\hat{Var}(q_{h})=22.63\text{ and }\hat{Var}(q)=33.66$
-This indicates that 67% of the variation in $q$ arises from variation within households.
-Furthermore, in the absence of within-household inequality, the between-household variance in children’s education attainment is 20.9.
-However, in the presence of within-household inequality, the between-household variance in children’s education attainment decreases to 11.2.
-These statistics suggest that, on average, households with some degree of within-household inequality exhibit lower between-household inequality compared to households with no within-household inequality.
-In conclusion, the analysis highlights on one hand the substantial contribution of within-household inequality to the overall inequality in education attainment.
-On the other hand, no within household variation in education of children is associated with higher between households variance.
-This suggests that households with greater disparities in education attainment among their members tend to exhibit lower disparities in education attainment when compared to other households.
-4.2 Magnitude of within household inequality in children’s education
-This section focuses on the extent of variation in within household inequality across households.
-I present evidence of differences in the magnitude of within household inequality across households.
-Understanding these differences can provide valuable insights into the factors that contribute to within household inequality and the potential mechanisms that can be employed to reduce it.
-Figure 2 depicts the empirical distribution of the within household range and standard deviation of the education attainment of children.
-This figure provides compelling evidence of the variation in within household inequality in the educational attainment of children.
-Specifically, the data reveals that the magnitude of inequality varies across households, with some household having all of their children with the same education attainment while some have at least a child with some college education and at least a child with no education.
-These findings highlight the importance of considering household-level dynamics when addressing educational inequality and suggest that interventions aimed at reducing disparities in education must be tailored to the unique circumstances of each household.
-The inequality is present even within gender, although in lower magnitude.
-About $40\%$ (resp.
-$60\%$) of household has some level of inequality among daughters (resp.
-among sons).
-The within household variance of children’s education is non-zero on average for all level of parents’ education and wealth index (see Figure 3).
-However, it appears that within household variance of children’s education decreases with parents’ education level and wealth index.
-At the household level, a negative association emerges between the maximum education attainment within a household and the proportion of children within that household who have achieved this maximum education level.
-We can have of this through an OLS regression of the within household maximum years of education of children on the proportion of children with education attainment equal to that maximum.
-$q^{\max}_{h}=\beta_{0}+\beta_{1}\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}1\{q_{i}=q^{\max}_{h}\}+\gamma^{\prime}X_{h}+\varepsilon_{h},$ (1)
-where $X_{h}$ include number of children, HWI, area of residence, religion, gender composition of children, and head of household’ s education.
-The estimation results in column (3)- (4) of Table 2 indicates that, on average, households with a 0.01 higher proportion of children attaining the maximum years of education within the household tend to exhibit around 2.6 years lower overall maximum education levels for children within the household.
-Additionally, an OLS regression of the within household maximum years of education of children on the within household standard deviation of children’s education indicates that households characterized by higher levels of educational inequality demonstrate, on average, higher within-household maximum education attainment (see column (1)- (2) of Table 2).
-These findings suggest a trade-off involved in households’ education decision.
-The same argument as Becker and Lewis (1973) applies here, i.e.
-an increase in quality  [^7] Here quality refer to the within household maximum years of education of children is more expensive if there are more children with that quality.
-An increase in quantity [^8] Quantity refers to the number of children with the within household maximum years of education of children is more expensive if children are of high quality.
-This trade-off is a direct effect of the limited education resources available to households.
-In conclusion, due to financial constraints within the household, parents are facing a trade-off between reducing inequality within the household or reducing inequality between them and other households.
-4.3 Parents’ education, average within household education of children, and inequality
-The level of education attained by the head of a household has been found to be a significant factor in determining the level of inequality in children’s education attainment within that household.
-In particular, an increase in the head of household’s education level is associated with a decrease in inequality.
-However, it remains unclear whether this is a direct result of more educated parents’ aversion for inequality or an indirect result of their preference for education.
-To shed light on this issue, this section will investigate the factors that contribute to the observed negative correlation between parents’ educational attainment and within household inequality.
-In addition to having lower level of inequality, households with more educated head of household also tend to have higher average years of education for their children (See Figure 5).
-This observation is particularly interesting given the hump-shaped relationship between inequality and average education of children (See Panel (b) of Figure 6).
-To investigate this relationship further, I estimation a OLS regression model of within inequality on average education of children, and parents’ level of education, with a quadratic interaction between between this two variables.
-$Inequality_{h}=\alpha+\beta_{1}\bar{q}_{h}+\beta_{2}\bar{q}^{2}_{h}\beta_{3}hh\_Educ_{h}+\beta_{4}\bar{q}_{h}hh\_Educ_{h}+\beta_{5}\bar{q}^{2}_{h}hh\_Educ_{h}+\gamma^{\prime}X_{h}+\varepsilon_{h},$ (2)
-where $X_{h}$ include number of children, HWI, area of residence, religion and gender composition.
-The estimation results in column (5)- (7) of Table 2 suggest that the negative dependence between parents’ education and within inequality is a result of both variable being correlated with the within household average education of children.
-In particular, the positive parents education between parents’ education and the within household average education of children combined with the hump shaped relation between within household inequality and the within household average education of children is translated into the observed spurious negative relationship between parents’ education and within household inequality.
-4.4 Observed and unobserved sources of inequality
-In the preceding sections, I have presented evidence at the household level, revealing that various factors contribute to the heterogeneity observed in the level of educational inequality across households.
-Notably, factors such as budget constraints, total investment in education, and the number of children play significant roles.
-By identifying and understanding these sources of heterogeneity, we gain valuable insights into the underlying mechanisms driving educational inequalities at the household level.
-In this section, the focus is on exploring the observed characteristics of children who received less education compared to their siblings.
-The examination of these characteristics is essential for developing effective strategies to address inequality and promote equality of opportunity for all children.
-Figure 7 graphs the average years of education based on the gender of children and the gender composition of households.
-To ensure accurate comparisons, the graph holds the within household average education of children constant.
-In the first panel, the analysis centers around households that are only able to finance primary school education for all their children.
-In the second panel, households that can only afford to provide education up to junior high school level are considered.
-The figure reveals some interesting trends in educational attainment among different gender composition households.
-Specifically, girls from only-daughter households, on average, have the same level of education as the household average, while boys from only-son households have similar education levels as well.
-However, in both-gender households, girls’ average education is lower than the household average, whereas boys’ average education is higher.
-These findings suggest that there is discrimination against daughters when it comes to the allocation of education quotas, when the alternative of giving more to a son is available.
-Figure 8 allows similar analysis in terms of children’s birth order after controlling for number of children, and within household average years of education of children.
-The figure provides insight into the average years of education of first and second children from two children households.
-In panel (a) of Figure 8, the plot is for households that can afford to educate all their children up to primary education, and for households that can afford to educate all their children up to junior HS education is in panel (b).
-The figure demonstrates that the average years of education for the firstborn children is below the household average for both type of households, whereas the average years of education of the second-born children is above the household average.
-This monotonic increase in education by birth order applies to any family size (See Appendix).
-The findings of Figure 8 suggest that there is disadvantage in birth order regarding the allocation of education quotas.
-In summary, a child’s gender, the gender and their siblings and their birth order are key determinants of the years of education they will receive.
-Despite taking into account observable household and children characteristics, a significant amount of variation in inequality across households remains unexplained, as evidenced by the $R^{2}$ value obtained from the regression of within household standard deviation of children’s education on those observable household and children characteristics (See the last two columns of Table 2).
-In addition, despite the presence of gender disadvantage against daughters, it appears that in some households, daughters receive higher education than their brothers (see Figure 9).
-I hypothesize that the unexplained difference in inequality can be attributed to the variance in children’s innate abilities, which differs across households.
-In other words, the fact that some daughters receive higher education compared to their brothers despite gender disadvantages can be attributed to high ability draws by these girls.
-This is a significant aspect of the household’s education distribution model, which I present in the next section.
-4.5 Decomposition of Within Household Inequality
-In the previous section, I have presented some empirical evidence about the observed characteristics of children which explain the within household inequality in their education.
-Additionally, it was demonstrated that a portion of this inequality can be attributed to the children’s unobserved abilities.
-In this section, I will provide a decomposition of the average within households inequality, categorizing it into components associated with gender disparity, birth order effects, and variations in children’s abilities.
-Such decomposition is conducted across various average educational levels within households on one hand and parents’ education level on the other hand.
-I used a household fixed-effect regression approach to achieve this breakdown.
-Regression with Household Fixed Effects
-To decompose the overall within household inequality into components categorized as disadvantage source and ability-based difference, I consider the following regressions:
-$Educ_{i,h}=\beta_{1}Female_{i,h}+\beta_{2}Firstborn_{i,h}+\beta_{3}Female_{i,h}\times Firstborn_{i,h}+\nu_{h}+\varepsilon_{i,h}$ (3)
-$Educ_{i,h}=\beta_{1}Firstborn_{i,h}+\nu_{h}+\varepsilon_{i,h}$ (4)
-where $Educ_{i,h}$ is the years of education of child $i$ in household $h$, $Female_{i,h}$ is a gender indicator variable equals to 1 if child $i$ in household $h$ is a daughter, $Firstborn_{i,h}$ is a birth order indicator variable equals to 1 if child $i$ in household $h$ is a firstborn, and $\nu_{h}$ is the household fixed effect.
-Equation 3 is for households with both sons and daughters, while equation 4 is for households with either only sons or only daughters.
-The estimates from equation 3 and 4 are presented in Table 3 by average education of children and in Table 4 by parents’ education.
-The results suggest, on one hand, that about $63\%$ of the observed within household inequality in children’s education is due to gender and birth order disadvantages for households with both son and daughter.
-On the other hand, for households with only daughters or only sons, about 33% of the observed inequality is due to birth order disadvantages.
-This change is due to the fact that part of the ability-based inequality is muted by gender disadvantage.
-In other words, gender disadvantage not only exacerbates overall inequality but also diminishes the inequality driven by differences in ability.
-For the primary analysis, which focuses on households with just two adult children living at home, the reliability of the estimates shown in Tables 3 and 4 may be compromised.
-This unreliability stems from the incidental parameter problem, a consequence of having only two data points per household for the fixed effect regressions.
-To validate the initial findings, I use the following alternative regression for a more robust examination.
-$\Delta_{daughter-son}Educ_{h}=\beta_{0}+\beta_{1}Firstborn\_daughter_{h}+\varepsilon_{h},$ (5)
-where $\Delta_{daughter-son}Educ_{h}$ is the average difference in the education of sons and daughters in household $h$, $Firstborn\_daughter_{h}$ is an indicator variable equals to 1 if the firstborn in household $h$ is a daughter.
-The estimates are summarized in Figure 11.
-Figure 11 illustrates the mean disparity in educational attainment between daughters and sons, with households having a firstborn male and female shown in red and blue, respectively.
-These measurements are provided across various average educational levels of the children in the panel a) and across education of the head of household in panel b), facilitating a decomposition for each average education level as follows:
-[…]
-and
-[…]
-Overall inequality
-By combining equations 6 and 7, we can derive the average impact of gender and birth order disadvantages on the inequality within households at each level of the children’s average educational attainment and parents’ education.
-This integration enables us to break down the average inequality found within households into the three factors illustrated in Figure 12.
-This figure showcases how the average inequality is divided among gender disadvantage, birth order disadvantage, and differences in ability.
-It reveals that gender disadvantage is the predominant factor contributing to inequality.
-As the average educational level of children or the education of the head of household increases, the influence of ability differences becomes more significant, while the impact of birth order diminishes.
-Similarly, as parents’ education level increases, total inequality is smaller on average, and the share of birth order disadvantage reduces; but the share of gender disadvantage does not change significantly.
-This indicates not only the presence of variability in the degree of inequality across different levels of children’s average education and parents’ education but also in the way it is broken down.
-Extensive versus Intensive margin inequality
-It is relevant to analysis how within household inequality in education is decomposed for the extensive margin compared to the intensive margin.
-To analyze that, I run the previous fixed effect regression in equations 3 and 4 for households with only educated children— for the intensive margin analysis—, and the following regression for households with at least one non-educated child— for the extensive margin analysis.
-$1\{Educ_{i,h}>0\}=\beta_{1}Female_{i,h}+\beta_{2}Firstborn_{i,h}+\beta_{3}Female_{i,h}\times Firstborn_{i,h}+\nu_{h}+\varepsilon_{i,h}$ (8)
-The estimates are presented in Table 5, and the decomposition of inequality at extensive and intensive margins is presented in Figure 13.
-The numbers indicate that parents’ education is negatively related to inequality in children’s education mostly at the extensive margin.
-In particular, panel a) of Figure 13 shows the proportion of households with a non-educated child by parents’ education.
-That number is the highest among non-educated parents ($\approx 50\%$) and close to 00 ($\approx 3\%$) among college educated parents.
-There is also a substantial heterogeneity in the decomposition of inequality at the extensive margin.
-Specifically, for most of non-educated households with a non-educated child, the non-educated child is either a daughter or a firstborn.
-This is not true among college educated parents.
-5 The Model
-5.1 Setup
-The model I have developed adopts a unique approach, considering children as investment goods rather than mere consumption goods.
-In other words, the number of children does not enter parents’ utility function directly like in Becker and Lewis (1973).
-The model consists of 2 distinct stages.
-In the first stage, households make decisions regarding the number of children, denoted as $N_{c}$, and observe their abilities, represented by the vector $\omega=\big{(}\omega_{1},\dots,\omega_{N_{c}}\big{)}$.
-They then choose the aggregate total years of education attainment, denoted as $q_{T}$, for these $N_{c}$ children.
-This leads to a within-household average years of education of children, represented as $\bar{q}=\frac{1}{N_{c}}q_{T}$.
-This initial stage can be viewed as choices derived from solving a fertility choice model, resembling the one described in Becker and Tomes (1976), with the distinction that each child is not assumed to receive $\bar{q}$ years of education.
-In other words, the decisions made in the first stage are based on the quantity-quality trade-off theory.
-This leads different choices for parents with different level of education.
-In addition to allowing an influence of parents’ education on the quantity of children and resources devoted to their schooling, I account for an unobserved heterogeneity that reflects parents’ aversion to having a child without education.
-This unobserved aversion parameter dictates the percentage of uneducated children within the family.
-In the second stage of the model, households shift their focus to deciding on the distribution of $q_{T}$.
-This decision is dictated by the aversion parameter combines with children’s observed and unobserved characteristics.
-Specifically, each household is characterized by a type $\nu_{h}$ (their aversion to having an uneducated child).
-Given $\nu_{h}$, a household chooses the proportion of $N_{c}$ that receives some education, and distributes $q_{T}$ among those children taking into consideration their gender, birth order and innate ability.
-The decision of parents is to choose the distribution $(q_{1},\dots q_{N_{c}})$ of $q_{T}$, which maximizes the household’s utility function.
-$\max_{q_{i}}U(q,\theta)$ (9)
-$\text{subject to }\sum q_{i}\leq q_{T},\text{ }q_{i}>0,\text{ }q_{i}\leq q_{\max}$
-$U(.)$ is increasing and concave, and $q_{\max}$ is the maximum years of education a child can receive. $\theta$ is described in the next section.
-5.2 Functional form of households’ utility
-Drawing upon the empirical evidence presented in Section 3, I have adopted the following generalized utilitarian social welfare function with concave utility function for each child to represent parents’ utility function.
-This form reflects the preferences and decision-making processes observed in the data and serves as a crucial component in the model.
-Let $q_{h}=(q_{1,h},\dots,q_{N_{c},h})$ be the distribution of $q_{T,h}$ in household $h$.
-The utility function for households with 2 children has the following expression [^9] The utility function for households with more than 2 children is presented in Appendix B.:
-$U(q_{h},\theta)=\nu_{h}\Big{[}\sum_{i=1}^{N_{c}}a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i}\Big{]}+(1-\nu_{h})\Big{\{}\sum_{i=1}^{N_{c}}s_{h}\Big{[}\nu_{i}^{s}.a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i}\Big{]}+$ (10)
-$\sum_{i=1}^{N_{c}}b_{h}^{d,s}\Big{[}\nu_{i}^{b_{d,s}}.a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i}\Big{]}+\sum_{i=1}^{N_{c}}b_{h}^{s,d}\Big{[}\nu_{i}^{b_{s,d}}.a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i}\Big{]}\Big{\}}$
-where $a_{i}$ captures parents’ preference for child $i$ relative to other children based on their ability draws, and $\alpha_{i}$ is the costs (financial and opportunity costs) of giving a year of education to the $i^{th}$ child.
-Note that $(\alpha_{(1)},\dots,\alpha_{(N_{c})})$ is the ordering of cost of education by children’s birth order.
-$a_{i}=\frac{\omega_{i}}{\sum_{j=1}^{N_{c}}\omega_{j}},\text{ }\nu_{i}^{gender\_comp}=1\{a_{i}.(q_{T})^{\delta_{i,h}}-\alpha_{i}q_{T}>a_{j}.(q_{T})^{\delta_{j,h}}-\alpha_{j}q_{T}\},\text{ }\nu_{j}^{gender\_comp}=1-\nu_{i}^{s},$
-$gender\_comp\in\{s,b_{d,s},b_{s,d}\}.$
-$s_{h}=Female_{i}.Female_{j}=(1-Female_{i}).(1-Female_{j})$
-$b_{h}^{d,s}=Female_{1}.(1-Female_{2}),\text{ }b_{h}^{s,d}=(1-Female_{1}).Female_{2}$
-For households with 2 children $\nu_{h}$ takes two possible values (High: H/Low:L).
-Households with high aversion for having an uneducated child choose to educate both children, whereas households with low aversion choose to educate only child out of the two children.
-For those who choose to educate only one child, choose the educated child based on a draw $\nu^{gender\_comp}_{i}$.
-$\delta_{i,h}=\delta(gender_{i},gender\_comp_{h})=\gamma-\theta_{1}Female_{i}\frac{1}{N_{c}-1}\sum_{\{i,j\in h\},j\neq i}(1-Female_{j})$
-$\nu_{h}\sim Bernoulli(p_{h}),\text{ }\nu_{1}^{s}\sim Bernoulli(p_{1}),\nu_{1}^{b_{d,s}}\sim Bernoulli(p_{fb,d}),\text{ }\nu_{2}^{b_{s,d}}\sim Bernoulli(p_{sb,d})$
-The vector of parameters of interest is $\theta=(\theta_{1},\alpha_{1}-\alpha_{2},p_{1},p_{fb,d},p_{sb,d})$.
-Assumption 1: The relative ability $a_{i}$ of a child $i$, is drawn from a distribution $G(.)$, with the constraint that $\sum_{i=1}^{n}a_{i,h}=1$.
-I assume that $G(.)$ is not dependent of gender or birth order.
-The incorporation of differences in $\delta$ across children and the household’s gender composition within the model allows for the consideration of disadvantage that females face at the intensive margin in terms of human capital investment when they have a brother.
-This feature accounts for the varying investment in education based on the child’s gender, reflecting potential gender disadvantage that may exist within the household— as evidenced in Figure 7.
-The assumed functional form is designed to capture the idea that girls with brothers receive a penalty in the distribution decision of the education quota made by parents.
-Additionally that penalty is an increasing function of the proportion of boys among the siblings.
-Similarly, the model’s incorporation of differences in $\alpha$ across children’s birth order enables the examination of the monotonic increase in education attainment as birth order advances.
-This acknowledges the tendency for education levels to rise sequentially with the child’s position within the family birth order— as observed in Figure 8.
-The integration of these elements into the model, allows for a more comprehensive understanding of the factors influencing educational outcomes, and facilitates the identification of potential areas for intervention to address gender-related disparities and birth order effects on educational attainment.
-The parameters $p_{1}$, $p_{fb,d}$, and $p_{sb,d}$ allow to take into account disadvantages faced by firstborn children and daughters in their education attainment at the extensive margin.
-Finally, by incorporating the ability-based preference parameter into the model, I enable the consideration of unexplained inequality that may arise from factors beyond the observable household and individual characteristics.
-This parameter captures the influence of individual abilities, which is not captured by the observed variables.
-To estimate the vector of parameter $\theta$, I used an indirect inference approach. In particular a simulated method of moments (SMM) approach. The procedure is outlined in the next section.
-5.3 Inference Strategy
-In this section, I provide an overview of the data moments utilized to identify the key parameters in the model.
-These data moments serve as essential empirical inputs that align the model’s predictions with the observed real-world outcomes.
-For the estimation of the parameters, I use two set of moments.
-The first set of moments includes two moments.
-Firstly, I consider the difference in average education between girls from households composed of only daughters and households with a mix of genders, while controlling for parents’ education and the number of children ($N_{c}$).
-This difference in educational outcomes can be attributed solely to gender disadvantage within households, making it a critical determinant in identifying the parameters involved in $\delta$.
-By isolating the impact of gender disadvantage on educational disparities, this moment provides a valuable means of disentangling the specific effects of gender-related preferences within the model.
-Secondly, I consider the average education attainment by birth order in households with children of the same gender only (either only daughters or only sons), holding fix the years of education of the head of household, and the number of children in the household.
-These average educational levels for different birth orders are used to calculate the differences in average education attainment between successive birth orders.
-These differences play a fundamental role in uniquely identifying the parameter associated with the opportunity costs ($\alpha_{(i)}$).
-The second set of moments used involves the proportion of educated daughters and the proportion of educated firstborn children by gender composition of households.
-This allows me to uniquely identify $p_{1}$, $p_{df,d}$, and $p_{sb,d}$.
-In particular the proportion of educated firstborn children in households with only children of the same gender, is used to estimate $p_{1}$.
-Similarly the proportion of educated firstborn daughters and second born daughters are used to estimate $p_{fb,d}$, and $p_{sb,d}$ respectively.
-For the rest of the analysis, let’s define the variables $Y_{h}^{d}$ as daughters’ education in household $h$ and $Y_{h}^{s}$ as sons’ education in the same household $h$.
-And let’s $Y_{h}^{1}$, $Y_{h}^{2}$, be the education of firstborn and second children respectively.
-Additionally, let $Z$ represent a vector comprising the observables, specifically the education of the head of the household and the number of children ($N_{c}$).
-Note that households with identical observable characteristics share the same type, and any differences observed in the variables $Y_{h}^{d}$ and $Y_{h}^{s}$, or in $Y_{h}^{1}$ and $Y_{h}^{2}$ between these households stem from disparities in the unobservable difference in children’s ability.
-This unobservable component captures the variability in educational outcomes that cannot be explained by the observable characteristics, such as parental ability-based preferences that influence educational investment decisions within households.
-[…]
-  1. 1.
-    $S_{d}^{d}=s\times H_{d}^{d}$ predictions of the education of daughters in households with only daughters,
-  2. 2.
-    $S_{d}^{b}=s\times H_{d}^{b}$ predictions of the education of daughters in households with both genders,
-  3. 3.
-    $S_{d}^{i}=s\times H_{d}$ predictions of the education of the $i^{th}$ born daughter in households with only daughters,
-  4. 4.
-    $S_{s}^{i}=s\times H_{d}$ predictions of the education of the $i^{th}$ born son in households with only sons,
-These predicted education attainments represent the educational outcomes based on the given parameter values.
-I then take the average of the $S_{l}^{m}$ predictions for each moment, where $l,m\in\{s,d,b\}$ = {only sons, only daughters, both gender}.
-To do inference on the parameters in $\delta$, the model and data moments are matched across various gender compositions.
-This process involves normalizing the parameter $a$ and estimating $\theta_{1}$ by comparing the model’s predictions to the observed data in terms of the difference in educational attainment for girls from households with only daughters and households with both genders.
-To do inference on the parameters $(\alpha_{(t)})_{t=1}^{N_{c}}$ associated with birth order, the model and data moments are matched across different birth orders.
-This process entails normalizing $\alpha_{(N_{c})}$ and estimating $(\alpha_{(t)})_{t=1}^{N_{c}-1}$ by comparing the model’s predictions to the observed data regarding the difference in educational attainment between $t^{th}$ and $(t+1)^{th}$ born children.
-To do inference on the parameters $p_{1}$, $p_{fb,d}$, and $p_{sb,d}$, the data moments and the model moments on the proportion of educated firstborn children, firstborn daughter, and second born daughters are matched with the model moments.
-5.4 Estimation
-In this section, I outline the estimation procedure used to estimate the parameters in the model.
-Let $\theta=(\theta_{1},\theta_{2},\theta_{3})$, where $\theta_{2}=(\alpha_{(1)},\dots,\alpha_{(N_{c}-1)})$, and $\theta_{3}=(p_{1},p_{fb,d},p_{sb,d})$.
-To estimate these parameters, I conduct $S$ simulations and compute the average education attainment of daughters in households with only daughters, only sons, and both genders as explained in the previous Section.
-Let $\hat{\mu}_{l}^{d}(\theta,Z)$ represent the average education attainment of daughters in these different household types, where $l\in{d,s,b}$ denotes households with only daughters, only sons, and both genders, respectively.
-Similarly, I compute the vector $\hat{\mu}(\theta,Z)=(\hat{\mu}_{1}(\theta,Z),\dots,\hat{\mu}_{N_{c}}(\theta,Z))$, which comprises the average education attainment by birth order.
-I can also compute $\hat{\pi}=(\hat{\pi}_{1},\hat{\pi}_{fb,d},\hat{\pi}_{sb,d})$, the model prediction of the proportions of firstborn children, of firstborn daughters and second born daughters.
-These simulations provide estimates of the model’s predictions for various household compositions, gender and birth orders, allowing for the comparison of the model’s outcomes with the observed data.
-The data moments are defined as follows:
-5.4.1 Moments Matched
-$m_{1}(Z)=E[Y^{d}|Z,Gender\_Comp=d]-E[Y^{d}|Z,Gender\_Comp=b],$
-$m_{2}(Z)=E[1\{Y^{1}>0\}|Z,Gender\_Comp=\{d,s\}],$
-$m_{3}(Z)=E[1\{Y^{d}>0\}|Z,Gender\_Comp=b,firstborn=d],$
-$m_{4}(Z)=E[1\{Y^{d}>0\}|Z,Gender\_Comp=b,firstborn=s],$
-$m_{t+4}(Z)=E[Y|Z,birth\_order=t+1]-E[Y|Z,birth\_order=t],\text{ }t\in\{1,\dots,N_{c}\}.$
-I matched the following data and model moments to estimate $\theta$.
-$m_{1}(Z)=\hat{\mu}^{d}_{d}(Z)-\hat{\mu}^{d}_{b}(Z),\text{ }m_{2}(Z)=\hat{\pi}_{1},\text{ }m_{3}(Z)=\hat{\pi}_{fb,d},\text{ }m_{4}(Z)=\hat{\pi}_{sb,d}\text{ and }$
-$m_{t+4}(Z)=\hat{\mu}_{t+1}(Z)-\hat{\mu}_{t}(Z);\text{ }t\in\{1,\dots N_{c}-1\}.$
-The corresponding sample objective function is the following expression:
-$\hat{Q}(\theta|Z=z)=(\bar{Y}_{d,z}^{d}-\bar{Y}_{b,z}^{d}-(\hat{\mu}_{d,z}^{d}-\hat{\mu}_{b,z}^{d}))^{2}+(\hat{m}_{2}(Z)-\hat{\pi}_{1})^{2}+(\hat{m}_{3}(Z)-\hat{\pi}_{fb,d})^{2}+(\hat{m}_{4}(Z)-\hat{\pi}_{sb,d})^{2}+$ (11)
-$\sum_{l\in\{d,s\}}\sum_{t=1}^{N_{c}-1}(\bar{Y}_{t+1,z}^{l}-\bar{Y}_{t,z}^{l}-(\hat{\mu}_{t+1,z}^{l}-\hat{\mu}_{t,z}^{l}))^{2}$
-$\hat{\theta}=\text{ argmin}_{\theta\in\Theta}\hat{Q}(\theta).$
-The sample objective function to possess a unique optimizer (See Figure 20).
-5.5 Estimation of G(.)
-I utilized auxiliary data to estimate the parameters of $G(.)$ outside of the model. In particular, I assume that the ability-based parents’ preference for children are i.i.d from a Beta distribution.
-$a_{i}\sim^{i.i.d}Beta(\beta_{1},\beta_{2})\text{ with }\sum_{i=1}^{N_{c}}a_{i}=1$
-$\beta_{1}$ and $\beta_{2}$ are estimated using auxiliary data.
-Specifically, I used data on average GPA in junior high school for a sample of student in Benin in 2018 to estimate $\beta_{1}$ and $\beta_{2}$ using maximum likelihood method.
-The Beta distribution seems to be a good fit for the distribution of relative ability (See Figure 21).
-5.6 Standard Errors
-The variance-covariance matrix for parameter estimates is given by the following expression:
-$\Omega=\Big{[}\frac{\partial m(\theta)^{\prime}}{\partial\theta}\Big{|}_{\hat{\theta}}V^{-1}\frac{\partial m(\theta)}{\partial\theta}\Big{|}_{\hat{\theta}}\Big{]}^{-1}$
-where, the partial derivative of the model function $\frac{\partial m(\theta)}{\partial\theta}$ is obtained numerically as,
-$\frac{\partial m(\theta)}{\partial\theta}\Big{|}_{\hat{\theta}}=\frac{m(\hat{\theta}+h)-m(\hat{\theta})}{h}$
-and, the variance-covariance matrix of the data moments, denoted as $V$, is computed through bootstrapping the data N times.
-This involves repeatedly calculating the moments N times and then using these N observations of the moments to determine the covariance between them.
-The standard errors are derived as the square roots of the diagonal elements of the matrix $\Omega$.
-5.7 The Model’s Performance
-In this section, I assess the effectiveness of stage 3 in the model through a straightforward exercise.
-This exercise consists of deriving the ability-based parameters $a_{i}$ from the model using $\hat{\theta}$.
-The empirical distribution of $\hat{a}$ should satisfy assumption 2 if the model has good performance.
-For a fix $\theta$ define the function
-$g:a_{1},a_{2}\in(0,1)^{2}\longrightarrow[0,21]^{2}$
-$(q_{1},q_{2})=g(\theta;a_{1},a_{2})$
-Proposition 1: $g(\theta;.)$ is injective, i.e. $g(\theta;a_{1},a_{2})=g(\theta,a_{1}^{\prime},a_{2}^{\prime})$ implies that $(a_{1},a_{2})=(a_{1}^{\prime},a_{2}^{\prime})$.
-This allows me to get the $\hat{a}_{i}$ for each children in each household.
-The distribution of $\hat{a}$ is plotted in Figure 15 for non-educated parents and Figure 16 for college educated parents.
-It shows that the distribution of $\hat{a}$ exhibits similarity across gender and birth order, albeit with minor differences.
-These disparities primarily stem from the heterogeneity in the disadvantages effects by different level of education, a feature not replicated in the predicted educational attainment by the model.
-To insure that this disparities do not affect the policy counterfactual analysis, I will use the distribution of $\hat{a}$ instead of $G(.)$ to do them.
-6 Model predictions and Counterfactual Analysis
-6.1 Results and Model Predictions
-The estimates of $\theta$ are provided in Table 6 for households with $N_{c}=2$ and non-educated and college educated parents.
-The estimation results reveal that both at the extensive and intensive margins households with non-educated head of households exhibit higher values for the gender disadvantage parameter when compared to households with college educated parents.
-This is in line with the reduced form results, which suggest that college-educated parents are less biased against daughters in terms of education based on the comparison of educational outcomes between daughters and sons in households with college-educated parents versus non-educated parents.
-In particular, $\hat{p}_{fb,d}=0.1124$, $\hat{p}_{1}=0.3663$ and $\hat{p}_{sb,d}=0.3217$ for non-educated parents.
-Recall that:
-$\nu_{1}^{fb,d}=1\{a_{1}.(q_{T})^{\delta_{1,h}}-\alpha_{1}q_{T}>a_{2}.(q_{T})^{\delta_{2,h}}-\alpha_{2}q_{T}\}\text{ and }\nu_{1}^{fb,d}\sim Bernoulli(p_{fb,d})$ (12)
-Equation 12 implies that:
-$\nu_{1}^{fb,d}=1\Big{\{}a_{1}>\frac{(\alpha_{1}-\alpha_{2})q_{T}+q_{T}^{\delta_{2,h}}}{q_{T}^{\delta_{1,h}}+q_{T}^{\delta_{2,h}}}\Big{\}}$
-This implies that
-$P[\nu_{1}^{fb,d}=1]=p_{fb,d}=P\Big{[}a_{1}>\frac{(\alpha_{1}-\alpha_{2})q_{T}+q_{T}^{\delta_{2,h}}}{q_{T}^{\delta_{1,h}}+q_{T}^{\delta_{2,h}}}\Big{]}$
-Given the distribution of $a_{1}$, and $\hat{p}_{fb,d}$, we get:
-$\frac{(\alpha_{1}-\alpha_{2})q_{T}+q_{T}^{\delta_{2,h}}}{q_{T}^{\delta_{1,h}}+q_{T}^{\delta_{2,h}}}=0.579$ (13)
-For households with both gender and second born daughter, we have:
-$\frac{(\alpha_{2}-\alpha_{1})q_{T}+q_{T}^{\delta_{1,h}}}{q_{T}^{\delta_{1,h}}+q_{T}^{\delta_{2,h}}}=0.53$ (14)
-For households with only daughters or only sons, we have:
-$\alpha_{1}-\alpha_{2}=\frac{1}{q_{T}}\Big{(}0.046\times q_{T}^{(1/2)}\Big{)}$ (15)
-This suggests that, on average, for parents without formal education and low aversion for having an uneducated child, the likelihood of the firstborn child being educated compared to a second-born child of the same gender is approximately 0.3663, which correspond to an average cost difference of $0.019$.
-To illustrate, if it costs $2500$ USD  [^10] This is the estimated cost of completing high school in Benin in 2003 (Foko et al.
-(2012)) to provide a high school education for a second-born child, it would require an additional $47$ USD to provide the same level of education for the firstborn child in families without formal education.
-This difference represents around 13% of Benin’s GDP per capita in the 1990s.
-Moreover, they perceive that investing in the education of the second-born son yields on average a 22% higher return compared to investing the same resources in the education of the firstborn daughter.
-Similarly, the return is estimated to be 27.8% higher for the firstborn son compared to the second-born daughter.
-Note that estimates for educational probabilities ($p_{1}$, $p_{fb,d}$, and $p_{sb,d}$) are not provided for college-educated parents, as nearly all of them, approximately 98%, have educated children.
-At the intensive margin, $\hat{\theta}_{1}=0.0218$, and $\hat{\alpha}_{1}=0.0018$ for non-educated parents.
-For college educated parents, $\hat{\theta}_{1}=0.0115$, and $\hat{\alpha}_{1}=0.0016$ for college educated parents.
-This indicates that parents without formal education perceive a 5.75% higher return on graduating high school for sons compared to daughters, while for college-educated parents, the difference is approximately 3%.
-The cost difference between providing up to high school education for firstborn compared to second born conditional on both being educated represents only 1.3% of Benin’s GDP per capita in the 1990s for non-educated parents.
-This disparity can be explained by a couple of key factors.
-First, college-educated parents often highly value education— they have an extreme aversion to having an uneducated child— and are more likely to prioritize educational opportunities for all their children, regardless of gender.
-They might view education as a means of empowerment and advancement, and thus, they are more likely to invest equally in the education of both daughters and sons.
-In addition, college-educated parents often have higher incomes and socioeconomic status, which can provide more resources for their children’s education.
-This increased financial stability can mitigate concerns about the perceived need to prioritize sons’ education over daughters’.
-Second, college-educated parents might be more aware of gender biases and stereotypes in society, including those related to education.
-They may actively work to counter these biases by ensuring equal opportunities for their children.
-Finally, there could be generational and cultural shifts at play, where younger, college-educated parents are more likely to challenge traditional gender roles and expectations regarding education.
-Overall, while these are potential reasons, it’s important to note that biases can still exist in any household, regardless of educational background.
-However, the data presented in the passage suggests a trend where college-educated parents appear to be less biased against daughters in terms of educational opportunities within their households.
-6.2 Counterfactual Analysis
-In this section, I provide contextual interpretations of the parameters within the model, shedding light on their real-world implications and the underlying mechanisms influencing distribution of education resources within households.
-I did this through two key counterfactual analyses, which allow me to explore the potential effects of policy interventions and alternative scenarios.
-These counterfactual analyses allow for a broader exploration of the model’s predictions under different conditions, helping to identify potential strategies to address educational inequalities and promote equitable educational opportunities for all children.
-The contextual interpretations and counterfactual analyses together contribute to a comprehensive evaluation of the model’s insights and provide valuable guidance for policymakers and researchers in devising effective measures to improve educational outcomes and reduce disparities within diverse household contexts.
-6.2.1 Counterfactual 1: Gender, birth order, and ability
-In this first counterfactual analysis, my main objective is to interpret the parameters $\theta$ by quantifying the additional level of ability required to counterbalance educational inequality resulting from gender and birth order disadvantages, respectively.
-To achieve this, I compute the extra ability needed by daughters and elder siblings to offset the effect of gender and birth order disadvantages on their education attainment.
-In order to do that I solve the household maximization problem in equation 9 with (using $\hat{\theta}$) and without disadvantages for a grid of relative ability of children for two-child families with a firstborn daughter and a second-born son, and compute the following quantity:
-  1. 1.
-    Ability of the firstborn daughter relative to the second born son at which the average difference between daughter’s and son’s education is equal to zero in presence of disadvantages.
-  2. 2.
-    The change in inequality due to gender and birth order disadvantages, by level of relative ability of the firstborn daughter.
-Figure 14 presents the first counterfactual analysis for non-educated and college educated parents.
-It suggests three main conclusions.
-First, for the same ability draws, gender and birth order disadvantages reduce the education attainment of the first born daughter by $\approx 4.6$ and 2.2 for non-educated parents and college educated parents respectively.
-Second, the average difference between firstborn daughters and second born sons is equal to 0 in the presence of disadvantages when the firstborn daughter’s ability draw is $\approx 78\%$ (resp.
-8%) higher than the ability draw of the second born son for households with non-educated head of household, (resp.
-households with college educated head of household).
-6.2.2 Counterfactual 2: Education Policies
-In situations where there is no inherent disadvantage against daughters or firstborn children, the primary factor leading to differences in education within a household between daughters and sons lies in variations in their individual abilities.
-Assuming that the initial abilities of children are distributed independently of their gender and birth order, the average difference in education between daughters and sons, in the absence of any disadvantages, forms a symmetric distribution centered around 0.
-However, this distribution shifts towards the negative side in instances where disadvantages are present.
-Simply put, in the absence of disadvantages, the distribution of the average education difference between daughters and sons exhibits first-order stochastic dominance over its counterpart in scenarios where disadvantages exist (See Figure 17).
-The objective of this section is to identify an education cost reduction strategy that would result in an education distribution between daughters and sons resembling scenarios where no disadvantages exist, even in the presence of such biases.
-Additionally, I examine the efficacy of a policy focusing on the extensive margin, wherein all parents strongly prioritize avoiding uneducated children.
-The cost reduction policy encompasses the following measures:
-  1. 1.
-    At the extensive (resp. intensive) margin remove cost difference between firstborn and second born. $\approx 1.9\%$ (resp. $\approx 0.18\%$) reduction in schooling cost for firstborn children.
-  2. 2.
-    At the extensive (resp. intensive) margin reduced schooling cost for firstborn daughters. $\approx 3\%$ (resp. $\approx 1.3\%$) reduction in schooling cost for daughters compared to sons.
-  3. 3.
-    At the extensive (resp. intensive) margin reduced schooling cost for second daughters. $\approx 3.8\%$ (resp. $\approx 1.3\%$) reduction in schooling cost for daughters compared to sons.
-Figure 18 displays the distribution of the average education difference between daughters and sons in panel (a), and between firstborn and second-born children of the same gender in panel (b), across four distinct scenarios.
-The blue curve represents the empirical distribution when disadvantages against daughters exist.
-Conversely, the red curve depicts the distribution in the absence of such biases.
-In the purple curve, we observe the distribution when disadvantages persist, but a cost reduction policy is implemented.
-Lastly, the orange curve illustrates the distribution when the disadvantage at the extensive margin is eradicated.
-By construction, the cost reduction policy effectively mirrors the distribution in scenarios devoid of disadvantages.
-The elimination of the extensive margin disadvantage proves effective only when the ability draws of sons are smaller than the ability draws of daughters.
-This explained by the fact that high draws of ability by daughters can counterbalance the effect of gender disadvantage in the intensive margin.
-However, this balancing effect is not observed at the extensive margin.
-Consequently, in households where daughters demonstrate higher ability draws than their brothers, the elimination of disadvantages at the extensive margin alone is sufficient to align the distribution of average education differences between daughters and sons with scenarios where there is no gender disadvantage.
-6.3 Counterfactual 3: Comparative Statics of Increase in Education Resources
-This section delves into the analysis of how a non-targeted increase in total education resources among non-educated parents affects the impact of gender and birth order disadvantages on within-household inequality.
-Initial empirical findings reveal that the average education level of children from college-educated parents with two educated adult children is 14.5, while it stands at 9.2 for children from non-educated parents.
-The exercise conducted in this section involves an exogenous increase in the average education level of children from non-educated parents, raising it from approximately 9.2 to 14.5.
-Subsequently, the household’s maximization problem is solved with this new average and $\hat{\theta}$.
-Finally, a comparison is drawn between the education differences of daughters and sons in this new scenario against the original disparity and the disparity observed for college-educated parents.
-The results are depicted in Figure 19.
-The Figure suggests that this non-targeted increase in education resources leads to increase in inequality against daughters both compared to the original disparities and compared to college educated parents.
-In summary, the analysis suggests that the non-targeted increase in education resources, while aiming to improve overall educational outcomes, had unintended consequences of widening the gap in educational attainment between daughters and sons within households where parents are not educated.
-7 Robustness of Estimates to Missing Siblings
-7.1 Gender Disadvantage
-The main dataset used for this analysis comprises adult children who were living in the same household as their parents during the census period.
-This sample represents a specific subgroup within the larger population of adult children.
-Importantly, the decision for children to leave the parental home is often influenced by factors such as their occupation and educational accomplishments, making it an endogenous process.
-Moreover, the motives for leaving home frequently differ between daughters, commonly associated with marriage, and sons.
-Given these dynamics, there’s a potential for bias in our estimates.
-This would be particularly concerning if, firstly, the children who remained at home are more similar to each other, and secondly, if they significantly differ from those who moved out.
-The wide range in both educational attainment and gender among children residing in the same household as their parents suggests that the first concern may not be significant.
-However, the second concern could lead to either overestimation – if women who moved out are more educated and men who moved out are less educated – or underestimation – if women who moved out are less educated and men who moved out are more educated, compared to those who remained at home.
-In this section, I delve into the potential bias in estimating the effect of gender disadvantage on within-household inequality.
-To investigate this, I compare the educational attainment of adult women and men living in the same households as their parents to those who have moved out.
-The mean comparison between these two groups is presented in Table 9.
-This comparison suggests that the difference in average education between men and women is more pronounced in the sub-sample that is not included in my analysis.
-As a result, it implies that, if anything, I may be underestimating the effect of gender disadvantage.
-Consequently, my estimate of gender disadvantage can be interpreted as an estimate of the lower bound of the true parameter.
-7.2 Birth Order Disadvantage
-A similar argument to the one presented in the previous section also applies to the birth order disadvantage parameter.
-The decision for children to move out is closely linked to their age, with older children being more inclined to leave their parents’ household.
-Consequently, we may have a selected sample of younger children in some households.
-In specific cases, children referred to as firstborns in certain households might actually be of a higher birth order.
-Additionally, more accomplished younger siblings may have already moved out.
-It’s important to note that both of these situations would potentially bias our estimate of the birth order disadvantage parameter downward.
-8 Conclusion
-In this paper, I examine the interaction between the three empirically known sources of disparities in children’s educational attainment within households.
-I constructed a structural model of households’ distribution of education resources among children, allowing for the influence of factors such gender, birth order and ability of children.
-The model not only allows me to decompose, for each relative ability draw, the total observed inequality into parts due gender, birth order, and ability differences; it also gives a platform for analysing how education cost reduction policies affects within household inequalities.
-The construction of the model is motivated by contexts similar to the one of Benin; a setting marked by notable disparities in children’s education within households, coupled with evidence of gender and birth order disadvantages.
-To ensure tractability, certain aspects of the parental decision-making process regarding education resources distribution are omitted.
-Notably, the model adopts a static approach, although the education decision of children is inherently dynamic.
-The primary objective of the paper being to rationalize the observed differences in children’s education, attributing them to gender disadvantages, birth order disadvantages, or variations in innate ability draws; despite its static nature, the model proves relevant, as it effectively incorporates and analyzes the interactions among these three factors.
-Additionally, the paper attributes any unexplained differences in children’s education, not accounted for by gender and birth order, to differential draws of innate ability.
-However, it acknowledges the potential influence of other unobserved factors, such as varying preferences of mothers in polygamous households, which could lead to increased parental investment in the education of specific children.
-In recognizing this, the interpretation of unexplained inequality within households is acknowledged as an upper bound of the effect of differential ability.
-In light of the findings in this paper, we can expect a reduction in the opportunity cost of girls education such as education support in the form of cash transfers, scholarships, and school kits for girls; to reduce within household inequality in children’s education that is due to gender disadvantage.
-Additionally, a reduction in the opportunity cost of education for firstborn, such as cash transfers and school kits, to young parents (first-time parents) or scholarships for firstborn children; is expected to reduce within household inequality in children’s education that is due to birth order disadvantage.
-However, these two policies need to be combined for an effective reduction in disadvantaged-based inequality.
-This is due to the possibility of displacement of disadvantage from one group to another.
-In particular, if the policy only targets firstborn children, the disadvantage against daughters might increase, and vice versa.
-9 Tables and Figures
+Conversely, approximately 31% of parents without schooling have children who likewise lack schooling, compared to only $\approx$ 10% for parents with schooling (See Figure 1).
+These statistics provide suggestive evidence of both inter-generational educational mobility  [^7] Children are more educated than their parents and inter-generational educational persistence [^8] Children’s education is correlated with their parents’ education..
+Third, regarding within-family inequality, the average within-household range in children’s education is about 7 years with a maximum of 21 years of education.
+In addition the standard deviation of the within-family range of children’s education is $\approx 4$ years of education, signaling high variability in the with-household inequality across households.
 (a) Children’s education as function of parents’ education
@@
 Figure 1: Parents and children’s education.
+3 Empirical Evidence
+Empirical Evidence 1: $2/3$ of the variation in educational attainment in the sample arises from disparities within households.
+I compute the average within-household variation in educational attainment and compare that to the overall variation in educational attainment in the sample.
+Furthermore, I examine how within and between-household variances in education are related.
+Let $q_{h}=(q_{h,1},q_{h,2},\dots,q_{h,N_{c_{h}}})$ be the vector of adult children’s educational attainment in household $h$ with $N_{c_{h}}$ adult children, and let $q=(q_{1},\dots,q_{n})$ be the educational attainment of adult children in the sample.
+$Var(q)=Var\Big{[}\mathbb{E}[q|h]\Big{]}+\mathbb{E}\Big{[}Var[q|h]\Big{]}.$
+The variance of $q$ is the sum of the average within-household variance ($\mathbb{E}\big{[}Var[q|h]\big{]}$) and between household variation $\big{(}Var\big{[}\mathbb{E}[q|h]\big{]}\big{)}$ in $q$.
+The estimates of these quantities in my sample are the following:
+$\widehat{\mathbb{E}}\Big{[}s(q|h)\Big{]}=22.63\text{ and }s(q)=33.66,$
+where $s(q|h)$ is the sample variance of adult children’s educational attainment $q_{h}$ in household $h$, $s(q)$ is the sample variance of adult children’s educational attainment $q$ in the whole sample, and $\widehat{\mathbb{E}}$ is the sample average.
+This indicates that $2/3$ of the variation in $q$ arises from variation within-households.
+Furthermore, in the absence of within-household inequality  [^9] When we consider households with disparities in their children’s educational attainment., the estimate of the between-household variance of children’s educational attainment is 20.9.
+However, in the presence of within-household inequality, the estimate of the between-household variance of children’s educational attainment is 11.2.
+These statistics suggest that, on average, households with some degree of within-household inequality exhibit lower between-household inequality compared to households with no within-household inequality.
+In conclusion, the analysis highlights on one hand the substantial contribution of within-household inequality to the overall inequality in educational attainment.
+On the other hand, no within-household variation in educational attainment of children is associated with higher between households variance.
+Empirical Evidence 2: Within-household disparities in children’s educational attainment is heterogeneous across households.
+This empirical evidence focuses on the extent of variation in within-household inequality across households.
+Understanding these differences can provide valuable insights into the factors that contribute to within-household inequality and the potential mechanisms that can be employed to reduce it.
+Figure 2 depicts the empirical distribution of the within-household range and standard deviation of the educational attainment of adult children.
+This figure reveals that the magnitude of inequality varies across households, with some household having all of their children with the same education attainment while some have at least a child with some college education and at least a child with no formal education.
 (a) Range
 (b) Standard Deviation
-Figure 2: Empirical cdf of within household range and standard deviation of education attainment of children.
+Figure 2: Empirical cdf of within-household range and standard deviation of education attainment of children.
+These findings highlight the importance of considering household-level dynamics when addressing educational attainment inequality and suggest that interventions aiming at reducing disparities in education must be tailored to the unique circumstances of each household.
+The inequality is present even within gender, although in lower magnitude.
+The average within household variance of daughters’ educational attainment (resp.
+sons) is 4.97 (resp.
+8.77).
+The within-household variance of children’s education is non-zero on average for all level of parents’ education and household wealth (see Figure 3).
+However, it appears that within-household variance of adult children’s educational attainment decreases with parents’ education level and household wealth.
+Specifically, we observe a first order stochastic dominance between the empirical cdf of within-household inequality in children’s educational attainment of college educated (resp.
+high wealth index) and non-college educated (resp.
+low wealth index) parents.
 (a) Parents Education
@@
 Figure 3: Distribution of inequality by socio-economic groups.
-Figure 4: Distribution of inequality by number of children.
-Figure 5: Within household average and standard deviation of children’s education as function of parents’ level of education.
-Figure 6: Distribution of average education attainment of children.
-Figure 7: Average years of education by gender and households gender composition ($N_{c}=2$)
-Figure 8: Average years of education by birth order ($N_{c}=2$)
-(a) Between daughters and sons
-(b) Between first and second born
-Figure 9: Histogram of within household difference in average education (Benin, 2013)
-(a) By gender
-(b) By birth order
-Figure 10: Distribution of children’s education for number of $N_{c}=2$ (Benin, 2013)
-(a) As function of within household average education of children
-(b) As function of head of household’s education
-Figure 11: Effect of gender and birth order disadvantages on within household inequality ($N_{c}=2$)
-(a) As function of within household average education of children
-(b) As function of head of household’s education
-Figure 12: Inequality decomposition ($N_{c}=2$)
-(a) Extensive Margin
-(b) Intensive margin
-Figure 13: Inequality decomposition as function of head of household’s education ($N_{c}=2$)
-(a) Non-educated head of household
-(b) College educated head of households
-Figure 14: Effect of gender and birth order disadvantages on inequality ($N_{c}=2$)
-(a) Firstborn vs. second born
-(b) Daughter vs. son
-Figure 15: Distribution of estimated ability-based parameters ($N_{c}$ = 2 and non-educated parents)
-(a) Firstborn vs. second born
-(b) Daughter vs. son
-Figure 16: Distribution of estimated ability-based parameters ($N_{c}$ = 2 and college educated parents)
-(a) Daughter v.s son
-(b) Firstborn v.s second born
-Figure 17: Distribution of the difference in children’s education ($N_{c}=2$)
-(a) Daughter v.s son
-(b) Firstborn v.s second born
-Figure 18: Distribution of the difference in children’s education ($N_{c}=2$)
-Figure 19: Distribution of difference in daughter and son’s education by parents’ education level
-Figure 20: Plot of $\hat{Q}(\theta)$
-Figure 21: Histogram of relative GPA in junior high school and histogram of random draws from Beta (28.82, 28.78).
-Tables
-Table 1: Descriptive Statistics
-Statistic N Mean St. Dev. Min Max
-Age 89,594 29.452 3.993 25 40
-Female 89,594 0.380 0.485 0 1
-Years of education 89,594 7.760 5.802 0 21
-At least one years of education 89,594 0.776 0.417 0 1
-Range of children’s education 89,594 6.821 4.873 0 21
-Standard deviation of children’s education 89,594 3.874 2.744 0.000 14.142
-Educated head of household 85,407 0.390 0.488 0 1
-Number of children between 25 and 40 89,594 3.055 1.557 2 16
-Number of children 89,594 6.340 4.354 2 79
-Educated with educated head of household 85,407 0.353 0.478 0 1
-Non-educated with non-educated head of household 89,594 0.180 0.384 0 1
-$\bar{q}$ 89,594 7.760 4.444 0.143 20.250
-$q_{T}$ 89,594 22.354 15.660 1 148
-Table 2: Regression of within household maximum years of education on within household inequality and of within household standard deviation of children’s education on households’ characteristics
+Empirical Evidence 3: At the household level, a negative association emerges between the maximum educational attainment among adult children within the household and the proportion of adult children within that household who have achieved that maximum education level.
+Consider the OLS regression of the within-household maximum years of educational attainment of adult children on the proportion of children with educational attainment equal to that maximum.
+$q^{\max}_{h}=\beta_{0}+\beta_{1}\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}1\{q_{i}=q^{\max}_{h}\}+\gamma^{\prime}X_{h}+\varepsilon_{h},$ (1)
+where $X_{h}$ include number of children, HWI, area of residence, religion, gender composition of children, and head of household’ s education.
+$q_{h}^{\max}$ is the maximum educational attainment of children in household $h$, $\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}1\{q_{i}=q^{\max}_{h}\}$, is the proportion of children with that maximum educational attainment within the household.
+The estimation results in column (3)- (4) of Table 2 indicate that, on average, households with a 0.5 higher proportion of children attaining the maximum years of education within the household tend to have around 1.3 years lower maximum educational attainment for the children within the household.
+Table 2: Regression of within- household maximum years of education on within-household inequality and of within-household standard deviation of children’s education on households’ characteristics
 Maximum years of education Standard deviation
@@
 Null hypothesis value outside the confidence interval.
-Table 3: Regression of children’s education on their gender and birth order with household fixed effect $(N_{c}=2)$
+Additionally, an OLS regression of the within-household maximum years of children’s educational attainment on the within-household standard deviation of children’s educational attainment indicates that households characterized by higher levels of educational inequality demonstrate, on average, higher within-household maximum educational attainment (see column (1)- (2) of Table 2).
+These findings suggest a trade-off involved in households’ educational decision.
+The same argument as Becker and Lewis (1973) applies here, i.e.
+on one hand an increase in quality  [^10] Here quality refer to the within-household maximum years of education of children is more expensive if there are more children with that quality.
+On the other hand, an increase in quantity [^11] Quantity refers to the number of children with the within-household maximum years of education of children is more expensive if children are of high quality.
+This trade-off is a direct effect of the limited education resources available to households.
+In conclusion, due to financial constraints within the household, parents are facing a trade-off between reducing inequality within the household or reducing inequality between them and other households.
+Empirical Evidence 4: The relationship between household-level mean and standard deviation of children’s education is inverted U-shaped.
+The level of education attained by the head of a household has been found to be a significant factor associated with the level of inequality in children’s educational attainment within that household.
+In particular, an increase in the head of household’s education level is associated with a decrease in inequality.
+However, it remains unclear whether this is a direct result of high educated parents’ aversion for inequality or an indirect result of their preference for education.
+To shed light on this issue, this section will investigate the factors that contribute to the observed negative correlation between parents’ educational attainment and within-household inequality.
+In addition to having lower level of inequality, households with more educated head of household also tend to have higher average years of education for their children (See Panel (a) of Figure 4).
+This observation is particularly interesting given the hump-shaped relationship between inequality and average educational attainment of children (See Panel (b) of Figure 4).
+This inverted U-shaped relationship between average and standard deviation of children’s education is consistent with the educational Kuznets curve theory (Thomas et al.
+(2003)).
+According to the Kuznets curve theory with education distribution, as we move from zero to maximum level of education, the variance first increases and then decreases.
+This is empirically shown for a set of developing countries in Londoño (1990) and Ram (1990).
+To investigate this relationship further, I estimate an OLS regression model of within-household inequality on average education of children, and parents’ level of education, with a quadratic interaction between between this two variables.
+$\text{Inequality}_{h}=\alpha+\beta_{1}\bar{q}_{h}+\beta_{2}\bar{q}^{2}_{h}+\beta_{3}\text{hh\_Educ}_{h}+\beta_{4}\bar{q}_{h}\text{hh\_Educ}_{h}+\beta_{5}\bar{q}^{2}_{h}\text{hh\_Educ}_{h}+\gamma^{\prime}X_{h}+\varepsilon_{h},$ (2)
+[…]
+Figure 4: Distribution of average education attainment of children.
+Empirical Evidence 5: Daughters with brothers and firstborn children receive on average less education compared to other children.
+In the preceding sections, I have presented evidence at the household level, revealing that various factors contribute to the heterogeneity observed in the level of educational inequality across households.
+Notably, factors such as budget constraints, total investment in education, and parents’ education play significant roles.
+In this section, the focus is on exploring the observed characteristics of children who received less education compared to their siblings.
+The examination of these characteristics is essential for developing effective strategies to address inequality and promote equality of opportunity for all children.
+Figure 5 graphs the average years of education based on the gender of children and the gender composition of households.
+To ensure accurate comparisons, the graph holds the within-household average educational attainment of children constant.
+In the first panel, the analysis centers around households that are only able to finance primary school education for all their children.
+In the second panel, households that can only afford to provide education up to junior high school level are considered.
+The figure reveals that, girls from only-daughter households, on average, have the same level of education as the household average, while boys from only-son households have similar education levels as well.
+However, in both-gender households, girls’ average education is lower than the household average, whereas boys’ average education is higher.
+These findings suggest that there is discrimination against daughters when it comes to the allocation of education resources, when the alternative of giving more to a son is available.
+Figure 5: Average years of education by gender and households gender composition (for $N_{c}=2$ and total within household educational attainment equal 12 and 20)
+Figure 6 allows similar analysis in terms of children’s birth order after holding fix the number of children, and the within-household average years of educational attainment of children.
+Figure 6: Average years of education by birth order (for $N_{c}=2$ and total within household educational attainment equal 12 and 20)
+In panel (a) of Figure 6, the plot is for households that can afford to educate all their children up to primary education, and for households that can afford to educate all their children up to junior HS education is in panel (b).
+The figure demonstrates that the average years of education for the firstborn children is below the household average for both type of households, whereas the average years of education of the second-born children is above the household average.
+This monotonic increase in educational attainment by birth order applies to any family size (See Appendix B).
+The findings of Figure 6 suggest that there is disadvantage in birth order regarding the allocation of educational resources.
+In summary, a child’s gender, the gender of their siblings, and their birth order are key determinants of the years of education they receive.
+Despite taking into account observed household and children characteristics, a significant amount of variation in inequality across households remains unexplained, as evidenced by the $R^{2}$ value obtained from the regression of within-household standard deviation of children’s education on those observed household and children characteristics (See columns (5)-(7) of Table 2).
+In addition, despite the presence of gender disadvantage against daughters in average educational attainment, it appears that in some households, daughters receive higher education than their brothers (see Figure 7).
+This suggest an heterogeneity in gender discrimination across households.
+A modeling assumption which I adopt in this paper stipulates that the unexplained difference in inequality can be attributed to the variance in children’s innate abilities, which differs across households.
+In other words, the fact that some daughters receive higher education compared to their brothers despite gender disadvantages can be attributed to high ability draws by these girls.
+This is a significant aspect of the household’s education distribution model, which I present in the next section.
+(a) Between daughters and sons ($N_{c}=2$)
+(b) Between 1st and 2nd born ($N_{c}=2$)
+Figure 7: Histogram of within-household difference in average education (Benin, 2013)
+Empirical Evidence 6-1: Average within-household inequality in children’s education is negatively related to parents’ education.
+Among households with non-educated heads and one child of each gender, over two-thirds of the average inequality is due to gender and birth order, while among college-educated parents, only one-third is due to these factors.
+In the previous section, I have presented some empirical evidence about the observed characteristics of children which explain the within-household inequality in their educational attainment.
+In this section, I will provide a decomposition of the average within-households inequality, categorizing it into components associated with gender disparity, birth order effects, and variations in children’s unobserved abilities (or any unobserved factors affecting educational resources distribution).
+The decomposition is conducted across various within-household average educational levels on one hand and parents’ education level on the other hand.
+I used a household fixed-effect regression approach to achieve this breakdown.
+Regression with Household Fixed Effects
+To decompose the average within-household inequality into components categorized as gender and birth order effects and unobserved differences, I consider the following regressions:
+$\text{Educ}_{i,h}=\beta_{1}\text{Female}_{i,h}+\beta_{2}\text{Firstborn}_{i,h}+\beta_{3}\text{Female}_{i,h}\times\text{Firstborn}_{i,h}+\nu_{h}+\varepsilon_{i,h}$ (3)
+$\text{Educ}_{i,h}=\beta_{1}\text{Firstborn}_{i,h}+\nu_{h}+\varepsilon_{i,h}$ (4)
+where $\text{Educ}_{i,h}$ is the years of education of child $i$ in household $h$, $\text{Female}_{i,h}$ is a gender indicator variable equal to 1 if child $i$ in household $h$ is a daughter, $\text{Firstborn}_{i,h}$ is a birth order indicator variable equal to 1 if child $i$ in household $h$ is a firstborn, and $\nu_{h}$ is the household fixed effect.
+Equation 3 is for households with both sons and daughters, while equation 4 is for households with either only sons or only daughters.
+The estimates from equation 3 and 4 are presented in Table 3 by average education of children and in Table 4 by parents’ education.
+The results suggest, on one hand, that about $63\%$ of the observed within-household inequality in children’s education is due to gender and birth order effects for households with both son and daughter.
+On the other hand, for households with only daughters or only sons, about 33% of the observed inequality is due to birth order effects.
+This change suggests that part of the unobserved sources of inequality is muted by gender disadvantage.
+For the primary analysis, which focuses on households with just two adult children living at home, the reliability of the estimates shown in Tables 3 and 4 may be compromised.
+This unreliability stems from the incidental parameter problem, a consequence of having only two data points per household for the fixed effect regressions.
+To validate the initial findings, I use the following alternative regression for a more robust examination.
+$\Delta_{\text{daughter-son}}\text{Educ}_{h}=\beta_{0}+\beta_{1}\text{Firstborn\_daughter}_{h}+\varepsilon_{h},$ (5)
+[…]
+Table 3: Regression of children’s education on their gender and birth order with household fixed effect by within-household total years of children’s education $(N_{c}=2)$
 (1) (2) (3) (4) (5) (6)
@@
 Note: Columns (2), (4), and (6) are for households with only daughters. For households with only sons the decomposition is 19% birth order + 81% ability.
-Table 4: Regression of children’s education on their gender and birth order with household fixed effect $(N_{c}=2)$
+Table 4: Regression of children’s education on their gender and birth order with household fixed effect by parents’ education $(N_{c}=2)$
 (1) (2) (3) (4) (5) (6)
@@
 For the whole sample it is 18% birth order + 82% ability.
+The estimates are summarized in Figure 9.
+Figure 9 illustrates the mean disparity in educational attainment between daughters and sons, for households with a firstborn son and a firstborn daughter separately.
+These measurements are provided across various average educational levels of the children in the panel a) and across education of the head of household in panel b), and are use to decompose the average absolute difference in children’s educational attainment  [^12] Let $q_{h}=(q_{1,h},q_{2,h})$, and $\text{Range}_{h}=\max(q_{h})-\min(q_{h})=|q_{1,h}-q_{2,h}|$.
+by household’s observable characteristics as follows:
+[…]
+[…]
+I use equations 6 and 7 to get the average effect of gender and the average effect of birth order on within-family disparities in educational attainment.
+Note that in households with firstborn daughters, the same child is affected by both gender and birth order disadvantages.
+Given that $|\beta_{0}+\beta_{1}|$ is the average effects of gender and birth order on educational attainment disparities within families, $E[\text{Range}|\text{household has a firstborn daughter}]-|\beta_{0}+\beta_{1}|$ is the unexplained residual131313 $E[\text{Range}|\text{household has a firstborn daughter}]\geq|\beta_{0}+\beta_{1}|$.
+I use these calculations to break down the average inequality found within-households into the three factors illustrated in Figure 8.
+Figure 8 displays how the average inequality is divided among gender effect, birth order effect, and differences in unobserved factors.
+It reveals that gender disadvantage is the predominant factor contributing to inequality.
+As the average educational level of children increases, the influence of unobserved ability differences becomes more significant, while the impact of birth order diminishes.
+Similarly, as parents’ education level increases, total inequality is smaller on average, and the shares of gender and birth order disadvantages reduce.
+This indicates not only that there is variability in the degree of average inequality in children’s educational attainment across different levels of children’s average education and parents’ education but also in the way it is broken down.
+(a) As function of within-household average education of children
+(b) As function of head of household’s education
+Figure 8: Inequality decomposition ($N_{c}=2$)
+(a) As function of within-household average education of children
+(b) As function of head of household’s education
+Figure 9: Effect of gender and birth order disadvantages on within-household inequality ($N_{c}=2$)
+Empirical Evidence 6-2: Intra-household educational inequality is present both at extensive and intensive margin.
+Compared to the extensive margins, the unexplained component has higher share in the average inequality for the intensive margin.
+The decrease in inequality by parents’ education is mostly present in the extensive margin.
+It is relevant to analysis how within-household inequality in education is decomposed for the extensive margin compared to the intensive margin.
+To analyze that, I run the previous fixed effect regression in equations 3 and 4 for households with only educated children— for the intensive margin analysis—, and the following regression for households with at least one non-educated child— for the extensive margin analysis.
+$1\{Educ_{i,h}>0\}=\beta_{1}Female_{i,h}+\beta_{2}Firstborn_{i,h}+\beta_{3}Female_{i,h}\times Firstborn_{i,h}+\nu_{h}+\varepsilon_{i,h}$ (8)
+The estimates are presented in Table 5, and the decomposition of inequality at extensive and intensive margins is presented in Figure 10.
+The numbers indicate that parents’ education is negatively related to inequality in children’s education mostly at the extensive margin.
+In particular, panel a) of Figure 10 shows the proportion of households with a non-educated child by parents’ education.
+That number is the highest among non-educated parents ($\approx 50\%$) and close to 00 ($\approx 3\%$) among college educated parents.
+There is also a substantial heterogeneity in the decomposition of inequality at the extensive margin.
+Specifically, for most of households with a non-educated head of household and a non-educated child, the non-educated child is either a daughter or a firstborn.
+This is not true among college educated parents.
 Table 5: Regression of children’s education on their gender and birth order with household fixed effect (Extensive vs Intensive margin) $(N_{c}=2)$
@@
 Null hypothesis value outside the confidence interval.
-Note:
-Table 6: Estimates of $\hat{\theta}$
+(a) Extensive Margin
+(b) Intensive margin
+Figure 10: Inequality decomposition as function of head of household’s education ($N_{c}=2$)
+4 Structural Model of Educational Attainment Choice
+4.1 Setup
+The model I propose considers children as investment goods rather than simple consumption goods.
+In other words, the number of children does not enter parents’ utility function directly like in Becker and Lewis (1973).
+Parents’ choices consist of 2 distinct stages.
+In the first stage, households make decisions regarding the number of children, denoted as $N_{c}$, and observe their abilities and other characteristics (which are unobserved to the econometricien), represented by the vector $\omega=\big{(}\omega_{1},\dots,\omega_{N_{c}}\big{)}$.
+They then choose the aggregate total years of educational attainment, denoted $q_{T}$, for these $N_{c}$ children.
+This leads to a within-household average years of education of children, denoted $\bar{q}=\frac{1}{N_{c}}q_{T}$.
+This initial stage can be viewed as choices derived from solving a fertility choice model, resembling the one described in Becker and Tomes (1976), with the distinction that each child is not assumed to receive $\bar{q}$ years of education.
+In other words, the decisions made in the first stage are based on the quantity-quality trade-off theory.
+This leads to different choices on average for parents with different level of education.
+Parents unobserved financial, social and cultural constraints
+In addition to allowing an influence of parents’ education on the quantity of children and resources devoted to their schooling, I account for an unobserved heterogeneity that reflects parents’ financial, social and cultural constraints, leading to limited educational opportunities for some of their children.
+In particular, there is an unobserved type for parents which creates barriers to school entry for some of their children.
+These barriers include limited household resources that might not be sufficient to educate all children and labor demands in agricultural families that makes them keep certain children out of school to contribute to farming activities.
+They also include religious beliefs about formal schooling.
+The unobserved type dictates the percentage of uneducated children observed within the family.
+For households with two children, the unobserved type can take 2 possible value {High (1), or Low (0)}.
+The type “High” means high financial, social, and cultural constraints and is associated with the presence of an uneducated child, whereas the type “Low” means low constraints and is associated with the absence of an uneducated child.
+I model parents’ unobserved types using the following binary choice model:
+For each household $h$ (Parents’ education and number of children hold fixed), let $\nu_{h}^{\star}$ be the unobserved financial, social, and cultural constraints, and $\nu_{h}$ be the observed discrete outcome of number of children with no formal education.
+$\nu_{h}^{\star}=X_{h}\beta+\varepsilon_{h}$, where $X_{h}$ includes a constant, an indicator for rural area of residence, an indicator for agricultural households, HWI, religion, gender composition, and average education of children.
+$\nu_{h}=1$ if $\nu_{h}^{\star}>0$, and 00 otherwise.
+I assume $\varepsilon$ follows a standard normal distribution, and estimate $\beta$ using a probit regression.
+Given $\hat{\beta}$, for each household $h$, $\nu_{h}$ is modeled as follows:
+$\nu_{h}\sim Bernoulli(\Phi[X_{h}\hat{\beta}]).$
+For number of children $>2$, I use a multinomial model.  [^14] The estimates of $\hat{\beta}$ are presented in Figure 20
+The maximization problem
+In the second stage of their decision, households decide on the distribution of $q_{T}$.
+This decision is function of their unobserved type combined with children’s observed and unobserved characteristics.
+Specifically, each household is characterized by a type $\nu_{h}$ (their level of financial hardship, social and cultural constraints).
+Given $\nu_{h}$, the household distributes $q_{T}$ among the children taking into consideration their gender, birth order and innate ability/other unobservables.
+The decision of parents is to choose the distribution $(q_{1},\dots q_{N_{c}})$ of $q_{T}$, which maximizes the household’s utility function.
+$\max_{q_{i}}U(q,\theta)$ (9)
+$\text{subject to }\sum q_{i}\leq q_{T},\text{ }q_{i}>0,\text{ }q_{i}\leq q_{\max}$
+I assume that $U(.)$ is increasing and concave, and $q_{\max}$ is the maximum years of education a child can receive.
+$\theta$ is the vector of parameters described in the next section.
+The model analyzes decisions in the second stage, taking choices in the first stage as given.
+Functional form of households’ utility [for households with 2 adult children]
+I use a generalized utilitarian social welfare function  [^15] The use of a generalized utilitarian social welfare function allows for flexibility in how children’s utilities are aggregated within the household.
+Individual utilities can be weighted or transformed, allowing for a range of preferences about inequality, risk, or the relative importance of individuals’ well-being.
+to represent parents’ utility function.
+This function incorporates a concave utility function derived from the education levels of each child.
+I chose this functional form because it better aligns with the observed mean-standard deviation curve, which differs from the one predicted by a linear utility function and more closely resembles a concave utility function from children’s educational outcomes (see Figure 17).
+Let $q_{h}=(q_{1,h},\dots,q_{N_{c},h})$ be the distribution of $q_{T,h}$ in household $h$.
+The utility function for households with 2 children has the following expression [^16] The utility function for households with more than 2 children is presented in Appendix B.:
+$U(q_{h})=(1-\nu_{h})\Big{[}\sum_{i=1}^{2}a_{i,h}.(q_{i,h})^{\delta_{i,h}^{\text{low}}}-\alpha_{i}^{\text{low}}q_{i,h}\Big{]}+\nu_{h}\Big{\{}\sum_{i=1}^{2}\Big{[}e_{i,h}.\big{(}a_{i,h}.(q_{i})^{\delta_{i,h}^{\text{high}}}-\alpha_{i}^{\text{high}}q_{i,h}\big{)}\Big{]}\Big{\}},$ (10)
+where,
+  • $a_{i,h}=\frac{\omega_{i,h}}{\sum_{j=1}^{N_{c_{h}}}\omega_{j,h}}\sim G(.)$ captures the unobserved ability-based weight on child $i$’s education relative to other children in household $h$.
+By definition $\sum_{i=1}^{2}a_{i,h}=1$.
+  • $\nu_{h}$, is parents’ level of financial, social, and cultural constraints.
+  • $e_{i,h}=1\{a_{i,h}.(q_{T})^{\delta_{i,h}^{\text{{high}}}}-\alpha_{i}^{\text{high}}q_{T}>a_{j,h}.(q_{T})^{\delta_{j,h}^{\text{high}}}-\alpha_{j}^{\text{high}}q_{T}\}$, $e_{j,h}=1-e_{i,h}$, $e_{i,h}$ and $e_{j,h}$ are indicator of whether not children $i$ and $j$ in household $h$ have some education.
+  • $\delta_{i,h}^{\text{low}}=\delta(\text{gender}_{i,h},\text{gender\_comp}_{h})=\gamma-\theta_{1}^{low}\text{Female}_{i,h}(1-\text{Female}_{j,h})$.
+  • $\delta_{i,h}^{\text{high}}=\gamma-\theta_{1}^{\text{high}}\text{Female}_{i,h}(1-\text{Female}_{j,h})$. [$\gamma$ is normalized to 0.5].
+  • $\delta_{i,h}$ is the marginal benefit from giving an additional year of education to child $i$ in household $h$.
+  • $\alpha_{i}^{\text{low}}$, and $\alpha_{i}^{\text{high}}$ are the costs (financial and opportunity costs) of giving a year of education to $i^{th}$ child at the extensive and intensive margin respectively.
+$\alpha_{2}^{\text{high}}$, and $\alpha_{2}^{\text{low}}$ are normalized to 0.
+  • $q_{i,h}$ is the years of education of child $i$ in household $h$.
+The vector of parameters of interest is
+$\theta=\Big{(}\theta_{1}^{\text{low}},\theta_{1}^{\text{ds, high}},\theta_{1}^{\text{sd, high}},(\alpha_{1}^{\text{high}}-\alpha_{2}^{\text{high}}),(\alpha_{1}^{\text{low}}-\alpha_{2}^{\text{low}})\Big{)}.$
+The utility from providing a $q_{i,h}$ level of education for each child [$u_{i,h}=a_{i,h}.(q_{i,h})^{\delta_{i,h}^{\text{low}}}-\alpha_{i}^{\text{low}}q_{i,h}$ or $u_{i,h}=a_{i,h}.(q_{i})^{\delta_{i,h}^{\text{high}}}-\alpha_{i}^{\text{high}}q_{i,h}$] in the parents utility has two parts: the benefit and the cost parts.
+Note that $a_{i,h}$ and $q_{i,h}$ are complementary in the benefit part of the utility from providing a $q_{i,h}$ level of education to child $i$ in household $h$.
+In other words, holding everything else fix, parents get higher utility by providing higher $q_{i,h}$ to child $i$ compared to child $j$ if $a_{i,h}>a_{j,h}$.
+Assumption 1: $a_{i,h}$ of a child $i$ in household $h$ is drawn from a distribution $G(.)$, with the constraint that $\sum_{i=1}^{N_{c_{h}}}a_{i,h}=1$.
+I assume that $G(.)$ is independent of gender and birth order.
+The incorporation of differences in $\delta$ across the children’s genders and the household’s gender composition within the model allows for the consideration of disadvantage that females face at the extensive and intensive margin in terms of human capital investment when they have a brother.
+This parameter models the difference in educational attainment by gender, reflecting potential gender disadvantage that may exist within the household— as evidenced in Figure 5.
+The assumed functional form is designed to capture the idea that girls with brothers receive a penalty in the distributional decision of the education resources made by parents.
+Additionally that penalty is an increasing function of the proportion of boys among the siblings.
+Similarly, the model allows for differences in $\alpha$ across children’s birth order to capture the monotonic increase in educational attainment as birth order advances as observed in Figure 6.
+To estimate the vector of parameter $\theta$, I used simulated method of moments.
+The procedure is outlined in the next section.
+The parameters are estimated for each level of education of parents holding number of children fixed.
+4.2 Estimation and Inference Strategy
+In this section, I provide an overview of the data moments used to estimate the key parameters in the model.
+I use two sets of moments for the parameters’ estimation.
+First, the difference in average educational attainment between daughters and sons in households with one child of each gender and no uneducated children, while holding fixed parents’ education and number of children; and the average educational attainment by birth order in households with children of the same gender only and no uneducated children, holding fixed the head of household’s education and number of children.
+This moments provide data variations to estimate $\delta^{\text{low}}$ and $\alpha_{(i)}^{\text{low}}$ respectively.
+The second set of moments includes the proportion of educated daughters and firstborn children in households with an uneducated child, while holding fixed parents’ education and number of children.
+These moments help estimate parameters in $\delta^{\text{high}}$ and $\alpha_{(i)}^{\text{high}}$.
+Specifically, the proportion of educated firstborn children in households with only children of the same gender and one uneducated child allow for the estimation of $\alpha_{(i)}^{\text{high}}$.
+The proportion of educated daughters from households with both genders and one uneducated child is used to estimate $\theta_{1}^{\text{high}}$.
+For the rest of this section, let’s define the variables $Y_{h}^{d}$ as daughters’ education in household $h$ and $Y_{h}^{s}$ as sons’ education in the same household $h$.
+And let $Y_{h}^{1}$, $Y_{h}^{2}$, be the education of firstborn and second born children respectively.
+Additionally, let $Z$ be a vector of observables, such as the education of the head of the household, the number of children ($N_{c}$), the aggregate education of children ($q_{T}$), and the gender composition of children.
+Note that for households with the same observed (Z) and unobserved ($\nu$) types, any differences observed in the variables $Y_{h}^{d}$ and $Y_{h}^{s}$, or in $Y_{h}^{1}$ and $Y_{h}^{2}$ between these households stem from disparities in the unobservable difference in children’s characteristics (eg.
+innate ability).
+[…]
+  1. 1.
+    $S_{d}^{d}=s\times H^{d}$ predictions of the educational attainment of daughters in households with only daughters, where $H^{d}$ is the number of simulated households with only daughters.
+  2. 2.
+    $S_{d}^{b}=s\times H^{b}$ predictions of the educational attainment of daughters in households with both genders, where $H^{b}$ is the number of simulated households with both genders.
+  3. 3.
+    $S_{d}^{i}=s\times H^{d}$ predictions of the educational attainment of the $i^{th}$ born daughter in households with only daughters.
+  4. 4.
+    $S_{s}^{i}=s\times H^{s}$ predictions of the education of the $i^{th}$ born son in households with only sons, where $H^{s}$ is the number of simulated households with only sons.
+These predicted educational attainments represent the educational outcomes based on the given parameter values.
+I then take the average of the $S_{l}^{m}$ predictions for each moment, where $l,m\in$only sons (s), only daughters (d), both gender (b) }.
+To do inference on the parameters in $\delta^{\text{low}}$, the model and data moments are matched across various gender compositions.
+This process involves normalizing the parameter $\gamma$ to 0.5 and estimating $\theta_{1}^{\text{low}}$ by matching the model’s predictions with the observed data in terms of the difference in educational attainment for daughters and sons from households with both genders.
+For the inference on the parameters $(\alpha_{(t)}^{\text{low}})_{t=1}^{N_{c}}$ associated with birth order, the model and data moments are matched across different birth orders.
+This process entails normalizing $\alpha_{(N_{c})^{\text{low}}}$ to 0 and estimating $(\alpha_{(t)}^{\text{low}})_{t=1}^{N_{c}-1}$ by comparing the model’s predictions to the observed data regarding the difference in educational attainment between $t^{th}$ and $(t+1)^{th}$ born children.
+For the parameters $\delta^{high}$, and $(\alpha_{(t)}^{high})_{t=1}^{N_{c}}$, the data moments and the model moments on the proportion of educated firstborn children from one gender households, firstborn daughter, and second born daughters from mix gender households are matched with the model moments.
+Let $\hat{\mu}_{l}^{d}(\theta,Z)$ represent the predicted average education attainment of daughters in different household types, where $l\in{d,s,b}$ denotes households with only daughters, only sons, and both genders, respectively.
+Similarly, let the vector $\hat{\mu}(\theta,Z)=(\hat{\mu}_{1}(\theta,Z),\dots,\hat{\mu}_{N_{c}}(\theta,Z))$, be the predicted average education attainment by birth order.
+Finally, let $\hat{\pi}=(\hat{\pi}_{1},\hat{\pi}_{d})$, be the model prediction of the proportions of firstborn children, and daughters for households with an uneducated child.
+These simulations provide estimates of the model’s predictions for various household compositions, gender and birth orders, allowing for the comparison of the model’s outcomes with the observed data.
+The data moments are defined as follows: Let $T_{\text{Educ}}$ be the total number of educated children.
+  • $m_{1}=E[Y^{s}|\text{Gender\_Comp}=b,T_{\text{Educ}}=2]-E[Y^{d}|\text{Gender\_Comp}=b,T_{\text{Educ}}=2],$
+  • $m_{2}=E[1\{Y^{1}>0\}|\text{Gender\_Comp}=s\text{ or }d,T_{\text{Educ}}=1],$
+  • $m_{3}=E[1\{Y^{d}>0\}|\text{Gender\_Comp}=b,T_{\text{Educ}}=1],$
+  • $m_{t+3}=E[Y_{t+1}|\text{Gender\_Comp}=s\text{ or }d,T_{\text{Educ}}=2\text{ }\&\text{ birth\_order }=t+1]-E[Y_{t}|\text{Gender\_Comp}=s\text{ or }d,T_{\text{Educ}}=2\text{ }\&\text{ }\text{birth\_order}=t],\text{ }t\in\{1,\dots,N_{c}\}.$
+I matched the following data and model moments to estimate $\theta$.
+$m_{1}=\hat{\mu}^{d}_{d}-\hat{\mu}^{d}_{b},\text{ }m_{2}=\hat{\pi}_{1},\text{ }m_{3}=\hat{\pi}_{d},\text{ and }m_{t+3}=\hat{\mu}_{t+1}-\hat{\mu}_{t};\text{ }t\in\{1,\dots N_{c}-1\}.$
+The corresponding sample objective function is the following expression:
+$\hat{Q}(\theta)=({\bar{Y}_{d,z}^{d}-\bar{Y}_{b,z}^{d}}-({\hat{\mu}_{d,z}^{d}-\hat{\mu}_{b,z}^{d}}))^{2}+({\hat{m}_{2}}-{\hat{\pi}_{1}})^{2}+({\hat{m}_{3}}-{\hat{\pi}_{d}})^{2}+\sum_{l\in\{d,s\}}({\bar{Y}_{2,z}^{l}-\bar{Y}_{1,z}^{l}}-({\hat{\mu}_{2,z}^{l}-\hat{\mu}_{1,z}^{l}}))^{2}.$ (11)
+$\hat{\theta}=\text{ argmin}_{\theta\in\Theta}\hat{Q}(\theta).$
+The sample objective function possess a unique local optimizer (See Figure 22). SMM standard errors are computed.
+4.2.1 Estimation of G(.)
+I use auxiliary data to estimate the parameters of the distribution $G(.)$ of the ability-based weight on children’s education outside of the model.
+In particular, I assume that the ability-based weights on children’s education are i.i.d across household and from a Dirichlet distribution.
+$a_{h}\sim^{i.i.d}\text{Dirichlet}(\eta_{1},\dots,\eta_{N_{c_{h}}})\text{ where, }a_{h}=(a_{1,h},\dots,a_{N_{c_{h}},h})$
+$(\eta_{1},\dots,\eta_{N_{c_{h}}})$ are estimated using auxiliary data.
+Specifically, I used data on average GPA in junior high school for a sample of student in Benin in 2018 to estimate $(\eta_{1},\dots,\eta_{N_{c_{h}}})$ using maximum likelihood method.
+The Dirichlet distribution seems to be a good fit for the distribution of relative ability (See Figure 25).
+5 Estimation Results and Counterfactual Analysis
+5.1 Estimation Results
+The estimates of $\theta$ are provided in Table 6 for households with $N_{c}=2$ children, for non-educated and college educated parents.
+Table 6: Estimates of $\hat{\theta}$, ($N_{c}=2$)
 Non-educated parents College educated parents
-$\hat{\theta}_{1}$ $\hat{\theta}_{2}$ $\hat{p}_{1}$ $\hat{p}_{fb,d}$ $\hat{p}_{sb,d}$ $\hat{\theta}_{1}$ $\hat{\theta}_{2}$
-Estimates 0.0218 0.0018 0.3663 0.1124 0.3217 0.0115 0.0016
-Standard errors 0.0017 0.005 0.0145 0.0119 0.0136
-Number of observations 11,270 942
-Table 7: Data and Model Moments: Targeted Moments
-Data Moments Model moments Difference
-$\bar{Y}_{d}^{d}-\bar{Y}_{d}^{b}$ 1.7028 1.7169 -0.0141
-Non-educated parents $\bar{Y}_{sb}^{d}-\bar{Y}_{fb}^{d}$ 0.2754 0.4502 -0.1748
-$\bar{Y}_{sb}^{s}-\bar{Y}_{fb}^{s}$ 0.3682 0.1616 0.2066
-Proportion of educated firstborn 0.3642 0.3678 0.0036
-Proportion of educated firstborn daughters 0.1116 0.1156 0.004
-Proportion of educated second born daughters 0.3177 0.3177 0
-$\bar{Y}_{d}^{d}-\bar{Y}_{d}^{b}$ 1.56 1.55 0.01
-College educated parents $\bar{Y}_{sb}^{d}-\bar{Y}_{fb}^{d}$ 2.61 2.76 -0.15
-$\bar{Y}_{sb}^{s}-\bar{Y}_{fb}^{s}$ 2.47 2.88 -0.41
-Proportion of educated firstborn
-Proportion of educated firstborn daughters
-Proportion of educated second born daughters
-Table 8: Data and Model Moments: Non targeted Moments
-Data Moments Model moments Difference
-Average inequality 5.87 5.97 -0.1
-Average inequality (only daughters households) 5.29 5.46 -0.17
-Average inequality (only sons households) 5.71 5.87 -0.16
-Average inequality (both gender households) 6.29 6.29 0
-Average inequality ($\bar{q}<6$) 5.05 5.04 0.01
-Non-educated parents Average inequality ($106\leq\bar{q}<$ 10) 8.46 7.56 0.9
-Average inequality ($1310\leq\bar{q}<$ 13) 5.27 5.34 -0.07
-Average inequality ($\bar{q}\geq 13$) 2.1 5.62 -3.52
-Average inequality (intensive margin) 3.59 3.77 -0.18
-Standard deviation of firstborn’s education 5.55 5.5
-Standard deviation of second born education 5.19 5.29
-Average inequality
-Average inequality (only daughters households)
-Average inequality (only sons households)
-Average inequality (both gender households)
-Average inequality ($\bar{q}<6$)
-College educated parents Average inequality ($106\leq\bar{q}<$ 10)
-Average inequality ($1310\leq\bar{q}<$ 13)
-Average inequality ($\bar{q}\geq 13$)
-Average inequality (intensive margin)
-Standard deviation of firstborn’s education
-Standard deviation of second born education
-Table 9: Difference in Average education of men and women
-Observed Unobserved
-Average years of education
-Male 5.35 4.66
-Female 4.73 2.03
-Difference 0.62 2.63
-Proportion of married
-Male 0.37 0.84
-Female 0.38 0.88
-Difference 0.01 -0.04
-Quantity-quality trade-off in Benin
-Table 10: Regression of education attainment dummy on number of children and other covariates
-1998-2013 1970-1998 1960-1970
-OLS IV OLS IV OLS IV
-(Intercept) $0.74^{*}$ $0.71^{***}$ $0.44^{*}$ $0.46^{***}$ $0.16^{*}$ $0.62^{*}$
-$[0.74;0.75]$ $(0.01)$ $[0.43;0.44]$ $(0.02)$ $[0.12;0.19]$ $(0.26)$
-Family size:Urban $-0.01^{*}$ $-0.01^{***}$ $-0.00^{*}$ $-0.00$ $0.00$ $0.01$
-$[-0.01;-0.01]$ $(0.00)$ $[-0.00;-0.00]$ $(0.01)$ $[-0.01;0.01]$ $(0.06)$
-Family size:Rural $-0.02^{*}$ $-0.02^{***}$ $-0.00^{*}$ $-0.01^{*}$ $0.00$ $-0.16$
-$[-0.03;-0.02]$ $(0.00)$ $[-0.00;-0.00]$ $(0.00)$ $[-0.00;0.01]$ $(0.10)$
-Education of hh $0.02^{*}$ $0.02^{***}$ $0.03^{*}$ $0.03^{***}$ $0.05^{*}$ $0.05^{***}$
-$[0.02;0.02]$ $(0.00)$ $[0.03;0.03]$ $(0.00)$ $[0.05;0.06]$ $(0.00)$
-Female hh $0.09^{*}$ $0.10^{***}$ $0.11^{*}$ $0.11^{***}$ $0.15^{*}$ $0.03$
-$[0.09;0.09]$ $(0.00)$ $[0.11;0.12]$ $(0.01)$ $[0.13;0.17]$ $(0.09)$
-HWI $0.04^{*}$ $0.04^{***}$ $0.06^{*}$ $0.06^{***}$ $0.07^{*}$ $0.07^{***}$
-$[0.04;0.04]$ $(0.00)$ $[0.06;0.06]$ $(0.00)$ $[0.06;0.08]$ $(0.01)$
-Urban $0.02^{*}$ $0.04^{***}$ $0.13^{*}$ $0.11^{***}$ $0.16^{*}$ $-0.36$
-$[0.02;0.02]$ $(0.01)$ $[0.12;0.13]$ $(0.03)$ $[0.12;0.19]$ $(0.30)$
-Birth order $0.02^{*}$ $0.01^{***}$ $0.01^{*}$ $0.02^{***}$ $0.01$ $0.12$
-$[0.01;0.02]$ $(0.00)$ $[0.01;0.01]$ $(0.00)$ $[-0.01;0.03]$ $(0.08)$
-Female $-0.04^{*}$ $-0.04^{***}$ $-0.08^{*}$ $-0.08^{***}$ $-0.12^{*}$ $-0.14^{***}$
-$[-0.04;-0.04]$ $(0.00)$ $[-0.09;-0.08]$ $(0.00)$ $[-0.14;-0.10]$ $(0.02)$
-R2 $0.14$ $0.14$ $0.21$ $0.21$ $0.27$ $-0.05$
-Adj. R2 $0.14$ $0.14$ $0.21$ $0.21$ $0.27$ $-0.05$
-Num. obs. $2069182$ $2069182$ $403850$ $403850$ $6461$ $6461$
-RMSE $0.40$ $0.43$ $0.40$
-${}^{***}p<0.001$; ${}^{**}p<0.01$; ${}^{*}p<0.05$ (or Null hypothesis value outside the confidence interval).
-10 Appendix A: Additional Figures
+$\hat{\theta}_{1}^{\text{low}}$ $\hat{\alpha}_{1}^{\text{low}}$ $\hat{\theta}_{1}^{\text{ high}}$ $\hat{\alpha}_{1}^{\text{high}}$ $\hat{\theta}_{1}^{\text{low}}$ $\hat{\alpha}_{1}^{\text{low}}$
+Estimates 0.0238∗∗ 0.0014 0.13∗∗ 0.017∗∗ 0.00759∗∗ 0.00045
+Standard errors 0.0017 0.0004 0.0019 0.001 0.0006 0.0012
+Number of observations 6187 5083 942
+**∗ ∗ significant at 5% level of significance.
+Result 1: For parents without formal education and with high financial, social and cultural constraints, parents’ perceived average utility at high school level of education is $\approx$ 11.5% higher for a second born child compared to firstborn child of the same gender.
+The estimate of the marginal educational cost difference between firstborn and second born children suggests that, on average, for parents without formal education and with high financial, social and cultural constraints, the likelihood of the firstborn child being educated compared to a second-born child of the same gender is approximately 0.3663, which correspond to an average cost difference of $0.02$.
+Holding everything else equal, this cost difference translate into a utility gap of $\approx$ 11.5% for high school level of education.
+Result 2: For parents without formal education and with high financial, social and cultural constraints, perceived average utility at high school level of education is $\approx$ 40% higher for the second born son compared to the firstborn daughter.
+Among parents without formal education and high financial, social and cultural constraints, the marginal utility from an additional year of education is 33% high if given to a son compared to given to a daughter.
+After factoring in birth order, these parents’ perceived average utility at high school level of education is $\approx$ 40% higher for the second born son compared to the firstborn daughter.
+Their perceived average utility of graduating high school is $\approx$ 18% higher for the firstborn son compared to the second born daughter .
+Note that estimates for the extensive margin parameters ($\hat{\theta}_{1}^{\text{high}}$ and $\hat{\alpha}_{1}^{high}$) are not provided for college-educated parents, as nearly all of them— approximately 98%— have only educated children.
+Result 3: Among parents with low financial, social and cultural constraints, the ones without formal education perceive a 6.6% higher utility on high school of education for sons compared to daughters, while for college-educated parents, the utility gap is approximately 2%.
+The parameter estimates for parents with low financial constraints indicate those without formal education perceive a 6.6% higher utility on graduating high school for sons compared to daughters, while for those with college education the difference is approximately 2.2%.
+The cost difference between providing a given level of education to firstborn children compared to second born is very small.
+For both non-educated and college-educated parents the estimates of the marginal costs different are not statistically different from 0 at 5% significance level.
+Parents’ perceived average utility at high school level of education is $\approx$ .9% (resp.
+0.3%) higher for a second born child compared to firstborn child of the same gender among non-educated parents (resp.
+college educated parents).
+5.2 The Model’s Fit
+In this section, I access the fit of the model.
+Using the estimated parameters $\hat{\theta}$, I solve the household maximization problem, to obtain the optimal distribution $q^{\star}_{h}$ of $q_{T_{h}}$ among children within each household $h$ across a simulated sample of $H$ households.
+I then compare $q^{\star}$ with the observed data to evaluate the model’s fit.
+First, I compute the intra-household educational difference between daughters and sons’ educational attainment, as well as between firstborn and second-born children using $q^{\star}$ for the simulated households.
+The empirical distributions of these within-household differences are compared with the ones from actual educational attainment observed in the data.
+Figure 11 provides a visual comparison of the model’s distributions with the observed data, showing a clear match.
+Second, I use $q^{\star}$ to derive targeted moments for estimation, alongside selected non-targeted moments.
+These analytical outcomes are summarized in Figure 23.
+It shows that there is no significant difference between the model and the data for all the targeted moments and most of the non-targeted moments as well.
+(a) Firstborn vs. second born
+(b) Daughter vs. son
+Figure 11: Empirical distribution of key moments: Data vs. Model ((For non-educated parents))
+5.3 Counterfactual Analysis
+Counterfactual 1: Without gender and birth order effects, the distribution of the average educational attainment difference between daughters and sons exhibits first-order stochastic dominance over the distribution in cases with such effects.
+In the absence of any gender and birth order effects, differences in education within a household between daughters and sons primarily result from variations in their individual unobserved characteristics/abilities.
+If children’s innate abilities are assumed to be distributed independently of gender and birth order, the average educational difference between daughters and sons, without considering gender or birth order effects, follows a symmetric distribution centered around 0.
+However, this distribution shifts towards the negative side in instances when gender and birth order disadvantages are present.
+In other words, without gender and birth order effects, the distribution of the average education difference between daughters and sons exhibits first-order stochastic dominance over the distribution in cases with such effects (see Figure 12 for non-educated parents).
+The distribution with no gender and birth order effects can be achieved through the substitution effect of a targeted educational cost reduction policy.
+Note however that such scenario did not change significantly the overall average within-family disparities in children’s educational attainment among non-educated parents.
+(a) Daughter v.s son
+(b) Firstborn v.s second born
+Figure 12: Distribution of the difference in children’s education ($N_{c}=2$)
+Counterfactual 2: At the intensive margin, the average educational attainment difference between firstborn daughters and second born sons is equal to 0 in the presence of gender and birth order effects when the firstborn daughter’s ability draw is $\approx 13\%$ (resp.
+8%) higher than the ability draw of the second born son for households with non-educated head of household, (resp.
+households with college educated head of household).
+In this counterfactual analysis, my primary goal is to assess how the unobserved source of inequality interact with the observed sources in the educational attainment choice within-family.
+I do this by quantifying the additional ability needed to counterbalance educational inequality due to gender and birth order effects.
+First, I compute the extra ability needed by daughters and older siblings to offset the effect of gender and birth order on their educational attainment.
+In order to do that I solve the household maximization problem in equation 9 with (using $\hat{\theta}$) and without (setting $\theta=0$) gender and birth order effects for a grid of relative ability of children for two-child families with a firstborn daughter and a second-born son, and compute the following quantities:
+  1. 1.
+    Ability of the firstborn daughter relative to the second born son at which the average difference between daughter’s and son’s education is equal to zero even in the presence of gender and birth order effects.
+  2. 2.
+    The change in inequality due to gender and birth order effects, by level of relative ability of the firstborn daughter.
+Figure 13 presents the results for non-educated and college educated parents with low financial, social and cultural constraints.
+It suggests two main conclusions.
+First, for the same ability draws, gender and birth order effects reduce the education attainment of the first born daughter by $\approx 2.2$ years and 1.2 years for non-educated parents and college educated parents respectively.
+Second, the average education difference between firstborn daughters and second born sons is equal to 0 in the presence of gender and birth order effects when the firstborn daughter’s ability draw is $\approx 13\%$ (resp.
+8%) higher than the ability draw of the second born son for households with non-educated head of household, (resp.
+households with college educated head of household).
+(a) Non-educated head of household
+(b) College educated head of households
+Figure 13: Effect of gender and birth order on inequality ($N_{c}=2$)
+5.3.1 Counterfactual 3
+This counterfactual is analyzed only among non-educated parents only, and the outcomes are compared to college-educated parents’ observed outcomes.
+Barriers to School Entry for all Children: In Theory
+The objective of this section is to examine the effectiveness of a counterfactual focusing on removing barriers to school entry, primarily addressing obstacles arising from parental decisions, to ensure that every child is enrolled in the school system.
+Theoretically this means setting $\nu_{h}$ to 0 for all households.
+Figure 14 displays the distribution of the difference in education between daughters and sons in panel (a), and between firstborn and second-born children of the same gender in panel (b), across three distinct situations.
+The elimination of barriers to school entry reduces gender and birth order effects and overall average within-family disparities in children’s educational attainment.
+In particular, the distribution of the difference between daughters’ and sons’ (resp.
+firstborn and second born children’s) educational attainments, after removing barriers to school entry for all children, second order stochastically dominates both the distributions with and without gender (and birth order) effects.
+This means that, compared to the situations there is no gender and birth order effects, removing barriers to school entry leads to more favorable and equitable educational outcomes.
+The overall distribution shifts in a way that is consistently better, resulting in a notable reduction in average inequality in the sample.
+(a) Daughter v.s son
+(b) Firstborn v.s second born
+Figure 14: Distribution of the difference in children’s education for non-educated parents ($N_{c}=2$) [Observed vs. with no barriers to school entry for all children]
+In summary removing barriers to school entry for all children reduces part of the gender and birth order effects on the within-household educational inequality.
+In addition, overall average within-household inequality is reduced by 50%.
+However, a significant share ($\approx 40\%$) of the average inequality is attributable to gender effect.
+In particular, it reduces the gender effect by 60% and birth order effect by 78%.
+That is consistent with the fact that gender effect is equally present in both the extensive and intensive margin where as the birth order effect is mostly present in the extensive margin.
+Barriers to School Entry for all Children: In Practice
+In the previous section, I set the financial, social and cultural constraint parameter to low for all households and analyze how within-family disparities in children’s educational attainment is affected.
+In this section, I analysis how this theoretical situation can be achieved with practical policy tools.
+Recall that financial, social and cultural constraints are determined by parents’ observable characteristics such as their area of residence (rural vs.
+urban), their sector of occupation (agricultural vs.
+non-agricultural), their religion (Christian vs.
+non-Christian), their level of wealth (low vs.
+high), and the average educational attainment of their children (a proxy for educational resources available).
+Some of these characteristics can be changed through a policy intervention.
+In particular, the financial constraints—wealth level and education resources.
+Observed data evidence show that children of college-educated parents with two adult children have an average education level of 14.5 years, compared to 9.2 for children of non-educated parents.
+In addition, average HWI [Household Wealth Index] is 0.03 among non-educated parents compared to 1.06 among college educated parents.
+In this section, I increase these two quantities for non-educated parents to the same level as college educated parents.
+The results show a decrease by 10% in the average disparities in children’s educational attainment within-family— not nearly as low as college-educated parents’ level— with 51% of the remaining inequality due to gender effects (see Figure 15).
+What happened is that the increase in income and educational resources was successful at removing barriers to school entry for children with non-educated parents [we observed a significant increase in the proportion of non-educated children], however, it also created more room for gender bias to create inequality in the intensive margin.
+And that is why it was not as effective as expected.
+In summary, some combination of awareness campaigns or laws promoting gender equality in education and an increase in households’ educational resources would be effective.
+Figure 15: Effects of increase in income and educational resources
+The following Table summarize the effectiveness of these different education policies counterfactual.
+Table 7: Summary of Education Policies Counterfactual
+Eliminate gender & Reduce average
+birth order effects inequality
+Remove Gender and birth order disadvantages (1)
+Remove barriers to school entry (2)
+Increase in income and educational resources (3)
+6 Conclusion
+In this paper, I examine the interaction between the three empirically known sources of disparities in children’s educational attainment within-households.
+I propose and estimate a structural model of households’ distribution of education resources among children, allowing for the influence of factors such as gender, birth order and ability of children.
+The model not only allows me to decompose, for each relative ability draw, the total observed inequality into parts due to gender, birth order, and ability differences; it also gives a platform for analysing how different education policies affect within-household inequalities.
+The construction of the model is motivated by contexts similar to the one of Benin; a setting marked by notable disparities in children’s education within-households, coupled with evidence of gender and birth order disadvantages.
+To ensure tractability, certain aspects of the parental decision-making process regarding education resources distribution are omitted.
+Notably, the model adopts a static approach, although the education decision of children is inherently dynamic.
+The primary objective of the paper being to rationalize the observed differences in children’s education, attributing them to gender effect, birth order effect, or variations in innate ability draws; despite its static nature, the model proves relevant, as it effectively incorporates and analyzes the interactions among these three factors.
+Additionally, the paper attributes any unexplained differences in children’s education, not accounted for by gender and birth order, to differential draws of innate ability.
+However, it acknowledges the potential influence of other unobserved factors, such as varying favoritism for wives in polygamous households, which could lead to increased parental investment in the education of specific children.
+In recognizing this, the interpretation of unexplained inequality within-households is acknowledged as an upper bound of the effect of differential ability.
+In light of the findings in this paper, we can expect a reduction in the opportunity cost of girls education such as education support in the form of cash transfers, scholarships, and school kits for girls; to reduce within-household inequality in children’s education that is due to gender disadvantage.
+Additionally, a reduction in the opportunity cost of education for firstborn, such as cash transfers and school kits, to young parents (first-time parents) or scholarships for firstborn children; is expected to reduce within-household inequality in children’s education that is due to birth order disadvantage.
+However, these two policies need to be combined for an effective reduction in disadvantaged-based inequality.
+This is due to the possibility of displacement of disadvantage from one group to another.
+In particular, if the policy only targets firstborn children, the disadvantage against daughters might increase, and vice versa.
+Finally, a compulsory education policy is the most effective in reducing average inequality in the sample.
+However, as long as there is budget constraint, as we move toward maximum education for everybody, there will always be a positive within-household inequality.
+7 Tables and Figures
+Figure 16: Distribution of inequality by number of children.
+Figure 17: Within-household standard deviation of children’s education as function of average education of children (Min vs. Max vs. Observed for $N_{c}=2$).
+(a) By gender
+(b) By birth order
+Figure 18: Distribution of children’s education for number of $N_{c}=2$ (Benin, 2013)
+(a) Gender composition of households as function of number of uneducated children
+(b) Number of uneducated children as function of gender composition of households
+Figure 19: Number of uneducated children by gender composition ($N_{c}=2$)
+(a) Estimate
+(b) Predicted probability as function of $\bar{q}$
+Figure 20: Estimates of $\hat{\beta}$
+Figure 22: Plots of the objective function
+(a) Average range
+(b) Standard deviation
+Figure 23: Data vs Model moments (For non-educated parents)
+Figure 24: Average Inequality after subsequent shot down of sources of inequality
+Figure 25: Histogram of relative GPA in junior high school and histogram of random draws from Beta (28.82, 28.78).
+(a) Mean
+(b) Standard Deviation
+Figure 26: Mean and standard deviation of adult between 25 and 40 years old
+(a) Women
+(b) Men
+Figure 27: Empirical cdf of the years of education of adults between 25 and 40 years old
+Tables
+Table 8: Educational inequality on average education
+(1) (2) (3) (4) (5) (6)
+Full sample Non-educated hh College educated hh
+Average education ($\bar{q}$) $1.25^{*}$ $1.04^{*}$ $1.44^{*}$ $1.26^{*}$ $0.87^{*}$ $0.64^{*}$
+$[1.24;1.27]$ $[1.02;1.07]$ $[1.42;1.46]$ $[1.23;1.29]$ $[0.81;0.94]$ $[0.54;0.75]$
+$\bar{q}^{2}$ $-0.08^{*}$ $-0.06^{*}$ $-0.09^{*}$ $-0.08^{*}$ $-0.05^{*}$ $-0.04^{*}$
+$[-0.08;-0.08]$ $[-0.06;-0.06]$ $[-0.09;-0.09]$ $[-0.08;-0.08]$ $[-0.05;-0.04]$ $[-0.04;-0.03]$
+R2 $0.69$ $0.71$ $0.73$ $0.74$ $0.61$ $0.62$
+Covariates
+Num. obs. $32729$ $32729$ $19558$ $19558$ $1438$ $1438$
+Null hypothesis value outside the confidence interval. Covariates include parents’ education, area of residence, religion, number of children,
+and gender composition of the household.
+8 Appendix A: Robustness of Estimates to Missing Siblings
+The main sample used for this analysis comprises adult children who were living in the same household as their parents during the census period.
+This sample represents a specific subgroup within the larger population of adult children.
+Importantly, the decision for children to leave the parental home is often influenced by factors such as their occupation and educational accomplishments, making it an endogenous process.
+Moreover, the motives for leaving home frequently differ between daughters, commonly associated with marriage, and sons.
+Given these dynamics, there’s a potential for bias in our estimates.
+This would be particularly concerning if, firstly, the children who remained at home are more similar to each other, and secondly, if they significantly differ from those who moved out.
+The wide range in both educational attainment and gender among children residing in the same household as their parents suggests that the first concern may not be significant.
+The second concern could lead to either overestimation— if women who moved out are more educated compared to ones who stayed and men who moved out are less educated compared to the ones who stayed— or underestimation— if women who moved out are less educated and men who moved out are more educated, compared to those who remained at home.
+In this section, I delve into the potential bias in estimating the effect of gender disadvantage on within-household inequality.
+To investigate this, I compare the educational attainment of adult women and men living in the same households as their parents to those who have moved out.
+The mean comparison between these two groups is presented in Figure 26.
+This comparison suggests that the difference in average education between men and women is more pronounced in the sub-sample that is not included in my analysis.
+In addition we observe a clear first order stochastic dominance between the empirical distribution of the education of adult female living in the same households as their parents and those who do not (see Figure 27).
+Such first order stochastic dominance is not as pronounced among men.
+As a result, it implies that, if anything, I may be underestimating the effect of gender disadvantage.
+Consequently, my estimate of gender disadvantage can be interpreted as an estimate of the lower bound of the true parameter.
+A similar argument to the one presented in the previous section also applies to the birth order disadvantage parameter.
+The decision for children to move out is closely linked to their age, with older children being more inclined to leave their parents’ household.
+Consequently, we may have a selected sample of younger children in some households.
+In specific cases, children referred to as firstborns in certain households might actually be of a higher birth order.
+Additionally, more accomplished younger siblings may have already moved out.
+It’s important to note that both of these situations would potentially bias our estimate of the birth order disadvantage parameter downward.
+In particular, if we maintain the assumption that firstborn children receive less education than other children, the older firstborn children with less education— who already moved out of the family house— are not included in our analysis.
+This leads to an underestimation of the birth order effect.
+In addition if high educated children are more likely to move out— because they have better and stable socio-economic status— we observe uniformly less educated children in our sample.
+In summary, we are likely to have in our sample, less educated children.
+On one hand, if the age effect dominates the education effect, we will have less firstborn children in our sample, which biases our estimate downward.
+On the other hand, if the education effect dominates the age effect, we have less second born children in our sample, which also biases our estimate downward.
+Appendix B
+8.1 Appendix B1: Summary Statistics
+Table 9: Summary Statistics for all the individuals between 25 and 40 years old residing in the same household as their parents
+Mean Standard Deviation
+All 5.11 5.94
+Rural 3.42 5.18
+Urban 7.05 6.14
+Farmer 2.91 4.83
+Education Non-farmer 6.75 6.15
+Non-Christian 3.43 5.23
+Christian 7.13 6.1
+Non-educated parents 3.45 5.16
+College-educated parents 13.8 4.42
+No education 0.48
+Female 0.38
+Rural 0.53
+Farmer 0.43
+Christian 0.45
+Non-educated parents 0.72
+College-educated parents 0.03
+HWI 0.03 0.98
+Number of children 5.35 4.1
+Number of children btw 25 and 40 2.1 1.51
+Average educ in hh is 0 0.4
+hh with only one child btw 25 and 40 0.45
+Number of observation 253,734
+8.2 Appendix B2: Model for households with Number of Children equal 3
+For $N_{c}=3$, let $ne\in\{0,1,2\}$ be number of children with no formal and education, and $\nu_{h}^{L}=1\{ne_{h}=0\}$, $\nu_{h}^{M}=1\{ne_{h}=1\}$, and $\nu_{h}^{H}=1\{ne_{h}=2\}$
+$U(q_{h},\theta)=\nu_{h}^{L}\Big{[}\sum_{i=1}^{N_{c}}a_{i}.(q_{i})^{\delta_{i,h}^{L}}-\alpha_{i}^{L}q_{i}\Big{]}+\nu_{h}^{M}\Big{\{}\sum_{i=1}^{N_{c}}\Big{[}e_{i}^{M}.\big{(}a_{i}.(q_{i})^{\delta_{i,h}^{M}}-\alpha_{i}^{M}q_{i}\big{)}\Big{]}\Big{\}}+\nu_{h}^{H}\Big{\{}\sum_{i=1}^{N_{c}}\Big{[}e_{i}^{H}.\Big{[}\big{(}a_{i}.(q_{i})^{\delta_{i,h}^{H}}-\alpha_{i}^{H}q_{i}\big{)}\Big{]}\Big{\}}$ (12)
+where,
+$\delta_{i,h}^{type}=\gamma-\theta_{1}^{type}Female_{i}\frac{1}{N_{c}-1}\sum_{\{i,j\in h\},j\neq i}(1-Female_{j}),\text{ with }type\in\{low(L),medium(M),high(H)\}$
+$e_{i}^{M}=1\{\exists j:\big{(}a_{i}.(q_{i})^{\delta_{i,h}^{M}}-\alpha_{i}^{M}q_{i}\big{)}>\big{(}a_{j}.(q_{i})^{\delta_{j,h}^{M}}-\alpha_{j}^{M}q_{j}\big{)}\},$
+$e_{i}^{H}=1\{\big{(}a_{i}.(q_{i})^{\delta_{i,h}^{H}}-\alpha_{i}^{H}q_{i}\big{)}>\big{(}a_{j}.(q_{i})^{\delta_{j,h}^{H}}-\alpha_{j}^{H}q_{j}\big{)},\forall j\neq i\},$
+The vector of parameters of interest is
+$\theta=\Big{(}\theta_{1}^{L},\theta_{1}^{M},\theta_{1}^{H},\alpha_{(1)}^{L},\alpha_{(2)}^{L},\alpha_{(1)}^{M},\alpha_{(2)}^{M},\alpha_{(1)}^{H},\alpha_{(2)}^{H}\Big{)}$
+$\alpha_{(3)}^{L}$, $\alpha_{(3)}^{M}$, $\alpha_{(3)}^{H}$ are normalized to 0.The dimension of $\theta$ is $1\times 9$.
+For each household $h$, [Parents’ educ and number of children hold fixed], let
+  • $\nu_{hne}^{\star}$ be the utility [due to unobserved financial, social & cultural constraints] from having $ne$ children with no formal education.
+  • $ne\in\{0,1,\dots,N_{c}-1\}$ be the observed discrete outcome of number of children with no formal educ.
+$\nu_{hne}^{\star}=X_{h}^{\prime}\beta_{ne}+\varepsilon_{hne}$
+[$X_{h}$ includes a cst, rural, farmer, HWI, religion, gender composition, and average education of children.]
+The choice of children with no formal education is
+$ne\text{ if }\nu_{hne}^{\star}>\nu_{hne^{\prime}}^{\star}\forall ne^{\prime}\neq ne$
+$ne_{h}=argmax_{ne}X_{h}^{\prime}\hat{\beta}_{ne}+\varepsilon_{hne}^{Sim}\text{, with }\varepsilon_{hne}^{Sim}\sim\mathcal{N}(0,\hat{\Sigma})$
+Table 10: Estimates of $\hat{\theta}$, ($N_{c}=3$)
+$\hat{\theta}_{1}^{L}$ $\hat{\alpha}_{1}^{L}-\hat{\alpha}_{2}^{L}$ $\hat{\alpha}_{2}^{L}-\hat{\alpha}_{3}^{L}$ $\hat{\theta}_{1}^{M}$ $\hat{\alpha}_{1}^{M}-\hat{\alpha}_{2}^{M}$ $\hat{\alpha}_{2}^{M}-\hat{\alpha}_{3}^{M}$ $\hat{\theta}_{1}^{H}$ $\hat{\alpha}_{1}^{H}-\hat{\alpha}_{2}^{H}$ $\hat{\alpha}_{2}^{H}-\hat{\alpha}_{3}^{H}$
+Estimates 0.033∗∗ 0.0026∗∗ 0.0010 0.0761∗∗ 0.0094 0.0001∗∗ 0.1596∗∗ 0.0120∗∗ 0.0105∗∗
+Standard errors 0.0052 0.0006 0.0006 0.0058 0.0016 0.0019 0.0073 0.0019 0.0015
+Number of observations 3644
+**∗ ∗ significant at 5% level of significance, * significant at 10% level of significance.
+8.3 Appendix B3: Generalized Households’ utility function for any Number of Children
+Let $type\in\{0,1,\dots N_{c}-1\}$, where
+  • 00 corresponds to the least constrained household,
+  • and $N_{c}-1$ corresponds to the most constrained household.
+$U(q_{h},\theta)=\sum_{c=2}^{N}\Big{\{}1\{N_{c_{h}}=c\}.U^{c}(q_{h},\theta)\Big{\}}\text{ where,}$, (13)
+  • $\delta_{i,h}^{type}=\gamma-\theta_{1}^{type}Female_{i}\frac{1}{N_{c_{h}}-1}\sum_{\{i,j\in h\},j\neq i}(1-Female_{j})$,
+  • $q_{h}=(q_{1,h},\dots,q_{N_{c},h})$
+9 Appendix C: Additional Figures
 (a) All Children.
@@
 Figure 29: Histogram of years of education.
-Figure 24: Distribution of within household range of children’s education.
+Figure 30: Distribution of within-household range of children’s education.
+Figure 31: Average years of education by gender and households gender composition and parents’ education ($N_{c}=2$)
 Figure 32: Average education by birth order for households with $N_{c}=2$ (Benin, 2013)
@@
 Figure 35: Average education by birth order for households with $N_{c}=5$ (Benin, 2013)
-Figure 29: Average education by birth order for households with $N_{c}=6$ (Benin, 2013)
-(a) 2 daughters, 1 son
-(b) 2 sons, 1 daughter
-Figure 30: Effect of gender and birth order disadvantages on within household inequality as function of head of household’s education ($N_{c}=3$)
-(a) 2 daughters, 1 son
-(b) 2 sons, 1 daughter
-Figure 31: Effect of gender and birth order disadvantages on within household inequality as function of head of household’s education ($N_{c}=3$)
-11 Appendix B: Households’ utility function for Number of Children equals 3
-The utility function maximized by households with 3 children has the following expression:
-$U(q_{h},\theta)=\nu_{h}\Big{[}\sum_{i=1}^{N_{c}}a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i}\Big{]}+\nu_{m}\Big{\{}\sum_{i=1}^{N_{c}}s_{h}\Big{[}\nu_{m,i}^{s}.(a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i})\Big{]}+$ (16)
-$\sum_{gender\_comp}\sum_{i=1}^{N_{c}}b_{h}^{gender\_comp}\Big{[}\nu_{m,i}^{{gender\_comp}}.(a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i})\Big{]}\Big{\}}+\nu_{l}\Big{\{}\sum_{i=1}^{N_{c}}s_{h}\Big{[}\nu_{l,i}^{s}.(a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i})\Big{]}+$
-$\sum_{gender\_comp}\sum_{i=1}^{N_{c}}b_{h}^{gender\_comp}\Big{[}\nu_{l,i}^{{gender\_comp}}.(a_{i}.(q_{i})^{\delta_{i,h}}-\alpha_{i}q_{i})\Big{]}\Big{\}}$
-where
-$gender\_comp\in\{(d,s,s);(d,d,s);(d,s,d);(s,d,d);(s,s,d);(s,d,s)\}$
-$s_{h}=Female_{i}.Female_{j}=(1-Female_{i}).(1-Female_{j})$
-$b_{h}^{d,s}=Female_{1}.(1-Female_{2}),\text{ }b_{h}^{s,d}=(1-Female_{1}).Female_{2}$
-For households with 3 children, aversion for having uneducated children takes three possible values (High: H/ Medium:M/ Low:L).
-Households with high aversion for having an uneducated child choose to educate both children.
-Households with medium aversion choose to educate 2 of their children, whereas households with low aversion choose to educate only child out of the three children.
-For those who choose to educate two children and only one child, choose the educated children based on a draw $\nu^{gender\_comp}_{m,i}$ and $\nu^{gender\_comp}_{l,i}$.
-$\delta_{i,h}=\delta(gender_{i},gender\_comp_{h})=\gamma-\theta_{1}Female_{i}\frac{1}{N_{c}-1}\sum_{\{i,j\in h\},j\neq i}(1-Female_{j})$
-$\nu_{h}\sim Bernoulli(p_{h}),\text{ }\nu_{m}\sim Bernoulli(p_{m}),\text{ }\nu_{l}\sim Bernoulli(p_{l}),\text{ with the constraint that }\nu_{h}+\nu_{m}+\nu_{l}=1$
-In practice, for each simulated household, the algorithm is as follows:
-  1. 1.
-    Take draw for $\nu_{h}\sim Bernoulli(p_{h})$,
-  2. 2.
-    If $\nu_{h}=1$, $\nu_{m}=\nu_{l}=0$ for that household,
-  3. 3.
-    If $\nu_{h}=0$, take draw for $\nu_{m}\sim Bernoulli(p_{m})$,
-  4. 4.
-    If $\nu_{m}=1$, $\nu_{l}=0$ for that household,
-  5. 5.
-    If not $\nu_{l}=1$. Note that this work because by construction in the data $\hat{p}_{h}+\hat{p}_{m}+\hat{p}_{l}=1$.
-$\nu_{m,i}^{s}\sim Bernoulli(p_{m,i}^{s}),\text{ }\nu_{l,i}^{s}\sim Bernoulli(p_{l,i}^{s})\text{ with }i\in\{1,2,3\}\text{ and }\sum_{i=1}^{3}\nu_{m,i}^{s}=2\text{ and }\sum_{i=1}^{3}\nu_{l,i}^{s}=1$
-Note that $p_{m,i}$ and $p_{l,i}$ are parameters of interest to be estimated. In practice for households with medium aversion, for each simulated household, the algorithm is as follows:
-  1. 1.
-    Take a draw for $\nu_{m,1}^{s}\sim Bernoulli(p_{m,1})$,
-  2. 2.
-    If $\nu_{m,1}^{s}=1$, take a draw for $\nu_{m,2}^{s}\sim Bernoulli(p_{m,2})$,
-  3. 3.
-    If $\nu_{m,2}=1$, $\nu_{m,3}=0$. If not $\nu_{m,3}=1$.
-For households with low aversion, the algorithm is similar to the one for $\nu_{h}$, $\nu_{m}$ and $\nu_{l}$.
-$\nu_{m,1}^{b_{gender\_comp}}\sim Bernoulli(p_{m}^{gender\_comp}),\text{ }\nu_{m,2}^{b_{gender\_comp}}\sim Bernoulli(p_{m}^{gender\_comp}),$
-$\nu_{m,3}^{b_{gender\_comp}}\sim Bernoulli(p_{m}^{gender\_comp})\nu_{l,1}^{b_{gender\_comp}}\sim Bernoulli(p_{l}^{gender\_comp}),\text{ }$
-$\nu_{l,2}^{b_{gender\_comp}}\sim Bernoulli(p_{l}^{gender\_comp}),\nu_{l,3}^{b_{gender\_comp}}\sim Bernoulli(p_{l}^{gender\_comp})$
-with the following conditions on $\nu_{m,i}^{gender\_comp}$s, and $\nu_{l,i}^{gender\_comp}$s.
-$\sum_{i=1}^{3}\nu_{m,i}^{gender\_comp}=2\text{ and }\sum_{i=1}^{3}\nu_{l,i}^{gender\_comp}=1$
-The vector of parameters of interest is
-$\theta=(\theta_{1},\alpha_{(1)}-\alpha_{(2)},\alpha_{(2)}-\alpha_{(3)},p_{m,1}^{s},p_{m,2}^{s},p_{l,1}^{s},p_{l,2}^{s},p_{m,1}^{gender\_comp},p_{m,2}^{gender\_comp},p_{l,1}^{gender\_comp},p_{l,2}^{gender\_comp})$
-and
-$p_{m,3}^{s}=1-(p_{m,1}^{s}+p_{m,2}^{s}),\text{ }p_{l,3}^{s}=1-(p_{l,1}^{s}+p_{l,2}^{s})$
-$p_{m,3}^{gender\_comp}=1-(p_{m,1}^{gender\_comp}+p_{m,2}^{gender\_comp}),\text{ }p_{l,3}^{gender\_comp}=1-(p_{l,1}^{gender\_comp}+p_{l,2}^{gender\_comp})$
-The dimension of $\theta$ is $1\times 31$.
-12 Appendix C: Model Performance
-For a fix $\theta$ define the function
-$g:a_{1},a_{2}\in(0,1)^{2}\longrightarrow[0,21]^{2}$
-$(q_{1},q_{2})=g(\theta;a_{1},a_{2})$
-Proposition 1: $g(\theta;.)$ is injective, i.e. $g(\theta;a_{1},a_{2})=g(\theta,a_{1}^{\prime},a_{2}^{\prime})$ implies that $(a_{1},a_{2})=(a_{1}^{\prime},a_{2}^{\prime})$.
-Proof of Proposition 1
-Let $g(\theta;a_{1},a_{2})=g(\theta;a_{1}^{\prime},a_{2}^{\prime})=(x,q_{T}-x)$. From the household’s maximization we must have $\forall x^{\prime}\neq x$
-$a_{1}x^{\delta_{1}}+a_{2}(q_{T}-x)^{\delta_{2}}-\alpha_{1}x-\alpha_{2}(q_{T}-x)\geq a_{1}x^{\prime\delta_{1}}+a_{2}(q_{T}-x^{\prime})^{\delta_{2}}-\alpha_{1}x^{\prime}-\alpha_{2}(q_{T}-x^{\prime})$
-and
-$a_{1}^{\prime}x^{\delta_{1}}+a_{2}^{\prime}(q_{T}-x)^{\delta_{2}}-\alpha_{1}x-\alpha_{2}(q_{T}-x)\geq a_{1}^{\prime}x^{\prime\delta_{1}}+a_{2}^{\prime}(q_{T}-x^{\prime})^{\delta_{2}}-\alpha_{1}x^{\prime}-\alpha_{2}(q_{T}-x^{\prime})$
-This implies that
-$(a_{1}+a_{1}^{\prime})x^{\delta_{1}}+(a_{2}+a_{2}^{\prime})(q_{T}-x)^{\delta_{2}}\geq(a_{1}+a_{1}^{\prime})x^{\prime\delta_{1}}+(a_{2}+a_{2}^{\prime})(q_{T}-x^{\prime})^{\delta_{2}}$ (17)
-which is not necessarily true if $(a_{1},a_{1})\neq(a_{1}^{\prime},a_{2}^{\prime})$. Hence $g(.)$ is injective.
-To see that equation 17 only holds if $(a_{1},a_{1})\neq(a_{1}^{\prime},a_{2}^{\prime})$, consider the following:
-  • $N_{c}=2$, $q_{T}=20$, $Female_{1}=0$ and $Female_{2}=1$.
-  • Initial ability vector correspond to $(0.57,0.43)$ with optimal distribution of $q_{T}$ equal to $q_{1}=15$, $q_{2}=5$
-  • The blue curve correspond to utility with change in ability but no change in education distribution.
-  • The red curve correspond to utility with change in ability and the corresponding optimal distribution of education.
+Figure 36: $\hat{\beta}$.
+Figure 37: Effect of gender and birth order on within-household inequality in education and Inequality Decomposition ($N_{c}=3$, Benin, 2013).